A Selected Bibliography of Publications by, and about, Wolfgang Pauli

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

22 November 2022 Version 1.119

Title word cross-reference

1/2 [SG78]. $27.95 [vB10]. $28 [Ano45a]. $29.95 [Sta07]. 3V [Raj77]. $6.00 [Dys60]. $80.00 [Sta07]. ^{+} [GP31]. a + A → C^{∗} → b_{1} + b_{2} + b_{3} [Kra65]. A_{n−1} [PZ88]. α [SKS85]. β [Fer34, Gau14]. f [Pau47c]. g [Kij72]. H [Pau28, PF37]. λ [Pau43a]. ∫↕(n, C) [Han11]. SL(3, C) [dGPS00, HNPT06]. n [PZ88]. npα [NMS84]. P [Pau38b]. R [ST82, ST84]. sl(n, C) [HPPT02]. sl(n^{k}, C) [Han10]. sl(p^{2}, C) [PST06]. V θ [Nel72]. X − Y [Suz73].

-Limiting [Pau43a]. -matrix [Kij72, ST84, ST82]. -Theorem [Pau28, PF37, Pau28, PF37]. -values [Pau47c]. -Werten [Pau47c].

/Friedrich [MKD^{+}91]. /Fritzsch [MKD^{+}91]. /Max [BT58, DT58]. /Max-Planck-Medaille [BT58, DT58]. /Rath [GRE^{+}01].

0 [Dys02, Str04a]. 0-19-856479-1 [Coo03, Dys02, Str04a].

1 [Dys02, Str04a]. 1.5 [GRE^{+}01]. 10 [Pau32a]. 100th [Enz00]. 11 [Pau33e]. 112 [FW07]. 137 [MM11]. 15.12.1958 [Hei59]. 175 [Ein79]. 19 [Ein79]. 19-175 [Ein79]. 1901-1950 [dVNHdVH53]. 1911 [Meh75]. 1929 [HEB^{+}80, HvMW79]. 1931 [Ehr31]. 1933 [CCJ^{+}34]. 1939 [OSWR86, vMHW85, HW87]. 1945 [Pau45b, Pau47d]. 1949 [vM93, Bro95]. 1952 [vM96]. 1954 [vM99]. 1954/Mladjenovic [ARK^{+}99]. 1956 [vM01]. 1958 [Mei92, Pei60b, vM05b, Enz05, Pes05, Sch02]. 1960s [ARK^{+}99]. 1960s/Frenkel [ARK^{+}99]. 1972 [Meh73]. 1983 [GHMP87]. 1984 [Wes85]. 1985 [LM85]. 1998 [Fis99, GR00]. 1st [GHMP87].

2002 [JG04]. 2006 [Ber06]. 2009 [Pet11]. 20th [Meh73]. 22 [CCJ^{+}34]. 23.50 [Che85]. 298.00 [Hen86].

3-540-13609-6 [Hen86]. 31 [Ehr31].

4/Pauli [OSWR86]. 49/Atmanspacher/Primas [MGG^{+}95].

60th [Pau45a, Pau94i].

7 [Pau29b]. 70th [KP55, PRW55].

8 [Fis99, Pau30a]. 89 [Sch01].

90-277-1648-X [Che85]. 9O [Bro86]. 9O-277-1648-X [Bro86].

= [Uns00].

A. [Pau26a, Pau31b, Pau51d, Pau53c]. A./Kraus [GRE^{+}01]. a.o [Hen86, HvMW79, vMHW85, vM93, vM96, vM99, vM01, vM05b]. abendländische [Pau56f, Pau59]. Abhandlungen [Bor63, Cha68]. Abschlusses [Pau25c]. Absorption [Dir27, Pau25d]. abstract [PD42b]. Abt [BDR84]. Abteilung [BDR93]. Acceleration [HEB^{+}80]. Acceleration/Brosche [HEB^{+}80]. Accelerator [JBE^{+}00, NFG^{+}91]. according [Fin49, PM26]. Account [Str04b, Str07, BSI^{+}82, ZSJY82]. action [Łop91]. actually [Bel86]. actuel [Pau53d]. Adam [DT18]. Adams [Pau25a]. Addendum [Hei79b, Hei79a]. Addition [BT58, DT58]. Additions [BT58, DT58]. adjoint [BP03]. Adventures [Wei91a]. AE [Ein79]. after [Kae48]. against [Mas04]. Age [Gil09]. ago [Enz82]. Aharonov [BP03, BBMS08, GS04, Hen81b]. Ähnlichkeitstransformationen [Pau40b]. akausaler [JP52]. Albert [Ano05, Pau61c, Kle05, Pau58a, Pau61c, Pau94a, Pau94f, Sch49a, Sch49b]. Alchemical [vExx]. Alchemistische [vExx]. algebra [Bay96, BPP64]. Algebraic [Nel72]. algebras [CP93, PZ88]. algorithm [FF87]. Allan [Sta07]. Allgemeine [Pau29a]. allgemeinen [Hei84, Pau33g, Pau58b]. allowed [FRKK95]. alpha [ACD90]. älteren [Pau57c, Pau61d]. am [Ehr31]. among [Sto24]. Analyses [dB86]. analysis [Ara90, GR00]. analysts [Kev00]. Analytical [QS98, BPP64]. Analytische [BPP64]. Anaximander [Sti77]. ancestors [Smu90]. Angular [Kra65, ZSJY82, HGF91, ZF00]. Anima [vE98]. Anima-Figure [vE98]. Anlage [Hei79b, Hei79a]. Anniversary [Pau01]. annotated [Sch01]. Announcement [Pau56b]. anomalen [Pau23b]. anomalies [Mas04]. Anomalous [BU26, Pau23b]. anomaly [Gau14]. Anordnung [ET60]. Ansprache [Ehr31, Pau56c]. Answers [Sta07]. anti [Dit15]. anti-Dirac [Dit15]. Anton [Fis10, Fis12, Fis10, Fis12]. Antwort [Pau31a]. Anwachsen [Pau28]. Anwendung [Pau74, Fre28]. Any [Tch72]. appearing [Pau25e]. Application [SZ97, Ara90, PJ44, Pau74, SS75, Fre28]. Applications [HEB^{+}80, Pau43a, PW35]. Applications/Lehmann [HEB^{+}80]. Applied [BT58, DT58]. Approximate [KM89]. approximation [RL77, SP77]. approximations [BP03]. April [JG04]. Äquivalenzverbot [JW28]. Arbeiten [ARK^{+}99, Pau24d, Pau77b]. Arbitrary [Fie39, DC86, FP39b, FP39a, Hal91]. Archetypal [Pau94g, Krö54, Pau52b]. Archetype [Pes05, Sch02, MEF01]. Archetypes [VNB08]. archetypischer [JP52, Krö54, Pau52b]. Argyris [MGG^{+}95]. Argyris/Faust [MGG^{+}95]. Armin [HW87]. Arnold [Ano38, BC51, Pau51a, Pau51b, Pau51c, Pau51e, Pau94b]. arrangement [ET60]. Arthur [vB10]. article [Dys60]. Ascona [AvMMH^{+}07]. aspects [Meh75, Pau54b]. Aspekte [Pau54b]. assignment [Pau23c]. assistance [All56]. Assistant [Pei85a, Enz00]. associated [Han10, Han11, HPPT02]. astounding [Haw11]. Asymptotic [Pau38a, QS98, Ara90]. Atom [Gie90, MEF01, MKD^{+}91, Pau25c, Ser77, vM90, Lau88, Nau13, Pes05, Sch02]. Atombaues [Pau29a]. Atomic [ARK^{+}99, Boh21, BBSR69, FR13b, Pau94u, RS74, SM08, Sto24, CCJ^{+}34, NFG^{+}91, Pau58c, ST82, ST84]. atomiques [CCJ^{+}34]. Atomkernes [Pau32b]. Atommechanik [Pau25b]. Atomphysik [MKD^{+}91, Pau58c]. Atomphysik/Ignatovich [MKD^{+}91]. Atoms [Ano45a, QS98, Str04b, Str07, Lau97, LT91, LT97, LT01, LT05, NM91]. attempts [Nav06]. Aufbau [Boy14]. Aufbruch [MGG^{+}95]. Aufgabe [MKD^{+}91]. Aufsätze [Pau61b, Pau61a]. Ausbreitung [Pau20a]. ausgeübten [PB27]. Ausgewählte [Bor63]. auspices [CCJ^{+}34]. Ausschliessungsprinzip [Pau77b]. äusseren [Pau58g, Pau58h, Pau23c]. Austrian [Ano93]. Automorphisms [HPPT02]. autostructure [Bad11]. autres [Kle05, Ano05]. Award [Ano45a, Ehr31].

B [BDR84, vM05b]. Back [GB30, GB29, GB30]. Back-Effekt [GB30]. Background [VNB08, CV02]. Band [ARK^{+}99, BPP64, Eis00, HvMW79, MGG^{+}95, Pau29b, Pau30a, Pau32a, Pau33e, Pau51e, Pau51d, vMHW85, vM93, vM96, vM99, vM01, vM05b, KP23]. Bandenspektren [KP23]. Bandwidth [CP84]. Bargmann [FRKK95]. barrier [SSF82]. Based [SZ97, ZSJY82]. Basis [GR31, FRKK95, dGPS00].

Bass [OSWR86]. Bau [Pau32b]. Bd [Eis02, HEB^{+}80]. Be [Ano45a, BT58, Coo03, DT58, Hei79a, Kap13, Tel03, Dys02, Enz02, Gou76, Str04a, Sul62]. beam [Pau19a]. Becker [DT18]. Bedeutung [APWB95, MGG^{+}95, Pau50b]. Beeinflussung [Pau24a]. Began [Kae48]. Begriff [Pau52a, vdWPR52]. Behandlung [Pau26a]. behavior [PB40, UG25]. Beitrag [Pau55b, Pau01]. beiträge [Pau48b, Pau35a]. beliebigem [FP39b, Fie39]. belief [vM87]. Believe [MKD^{+}91]. Believed [Ano45a]. Bell [BH09]. Belyaev [MKD^{+}91]. Bemerkungen [Pau20b, Pau24d]. Bericht [Hei84]. Berlin [Eis00, Hen86, Sch01]. Berlin/Heidelberg/New [Eis00]. Bern [Pau01]. Berner [Pau01]. Bestimmung [Pau47c]. Beta [Jen00]. betreffenden [Pau33a]. Between [Bur58, Pau40a, CP92, Del61, Ehr27b, Ehr27a, Fis99, Fis04, Fis11, LZ58, Pau23a, Pau24d, Pau50a]. bewegten [Pau20a]. bewegter [Pau33h]. Beyond [Lau88, Nat03, MKD^{+}91, vM90, Gie90]. bezüglich [UG25]. Bicentennial [Pet11]. Bifurkation [MGG^{+}95]. Big [Hür22, Scu07c]. Bildung [JP52, Krö54, Pau52b]. binding [BLV05, HSS11]. Biografie [Enz05]. Biographical [Enz94b]. Biographie [RP79]. biographies [Ano64, SS98]. Biography [Coo03, Enz05, Tel03, Dys02, Enz02, RP79, Str04a]. Biot [Sza97]. Bird [Pei85b]. Birth [BH83a, BH83b, Gar63a, BH82]. Birthday [Pau45a, Pau94i, Enz00, KP55, PRW55, Mor56a]. bis [Fis99, Fis10, Fis12, Sti77]. Bisherige [Hei79b, Hei79a]. bisherighen [Hei84]. bismuth [BG28, GB27, Tho26b, ZBG30]. black [Pau29d]. black-body [Pau29d]. Blockade [Pec93]. Body [MKD^{+}91, NMS84, Pau29d, SOCg82]. Bohm [BP03, BBMS08, GS04, Hen81b]. Bohmsche [GRE^{+}01]. Bohr [All56, ARK^{+}99, Bel86, Bro86, Bro95, Che85, Eis00, GRE^{+}01, HEB^{+}80, Hen86, HvMW79, Lak96, MGG^{+}95, Mor56a, OSWR86, Wes85, dB86, vMHW85, vM93, vM96, vM99, vM01, vM05b, Deg89, CG14, De 14, Hen84, KP55, Mas04, Nau13, OBxx, Pau45a, PRW55, PB58, Pau94i, Pau01, SM08]. Bohr-Festschrift [Pau01]. Boltzmann [Eis91]. Book [Bal65, Bel86, Ber06, Bor47, Bro86, Bro95, Che85, DT18, Dys60, Dys02, Gie90, HW87, Hen86, Kra09, Mor56a, Pau24b, Pau25b, Pau25a, Pau26a, Pau29b, Pau29c, Pau30a, Pau30b, Pau30c, Pau31b, Pau31c, Pau32a, Pau32b, Pau33e, Pau33f, Pau35b, Pau47a, Pau51e, Pau51d, Pau56d, Pes05, Sch02, Sta07, Sti77, Str04a, Tel03, Wes85, YK73, vB10, vM90, BPP64, BSRD^{+}55, LPB^{+}56]. Books [All56, BDR84]. Bor [PB58]. Born [Pau25b, Pau30b]. Boru [PB58]. Bose [Pau36a]. boson [Ara90]. Boston [Che85, Wes85]. bound [Sch01]. Brandmüller [MGG^{+}95]. Break [Pau72a]. Brecht [Sch99a, Sch99b]. Breit [Bad11, BG88, FWS^{+}85, FM68b, FM68a, HGF91, RL77, ST82, ST84, ZF00]. Breit-type [RL77]. bridge [BS12, Fis04, Fis11]. Brief [Coo03, Tel03, Dys02, Enz02, Str04a]. Briefe [Hei79b, vM68]. Briefwechsel [ARK^{+}99, Eis00, GRE^{+}01, HEB^{+}80, Hen86, HvMW79, MGG^{+}95, Mei92, OSWR86, vMHW85, vM93, vM96, vM99, vM01, vM05b, Eis02]. Brod [MGG^{+}95]. Brosche [HEB^{+}80]. Brücke [BS12]. Brücken [Fis04, Fis11]. Bruno [Bon05]. Brussels [CCJ^{+}34]. Bruxelles [CCJ^{+}34, Far01].

Buchbesprechungen [BSRD^{+}55, LPB^{+}56, MGG^{+}95, OSWR86]. Bücher [ARK^{+}99, BDR84, GRE^{+}01, HEB^{+}80]. Buchwald [Pau24b]. bulk [Eck70, Eck73, OBxx]. Bunde [MGG^{+}95]. Burger [Pau24d].

C [Coo03, Enz00, Fis99, JP52, MM11, Pau24d, YK73]. C. [Fis99, Gie05, Pes05, Sch02]. cachés [Pau53b]. calculate [ST82, ST84]. Calculation [SZ97, ZSJY82]. calculations [BG88, Bay96]. Calculus [Eck26, BPP64]. Cambridge [Sta07]. can [Mas04, Hei79a, Kap13, MKD^{+}91]. Cape [Ber06]. Carl [GRE^{+}01, vB10, Eck70, Eck73, Mil09]. Case [PH45, dR99, Gau14, Pau40b]. Casimir [Łop91]. Cassidy [MGG^{+}95]. Catalogue [Eur69]. Catalonia [GHMP87]. Causality [Pau94r, Pau36c]. Cavendish [Nav06]. centennial [vM00]. Center [FM68b, FM68a]. Centre [Meh73]. century [Bra09, Cas83, Cas92, CS10, Meh73, FW60]. Ceremony [Pau31a]. CERN [Ano60a]. Ch [BT58, DT58]. Challenge [Ber06, Sta09]. champs [Pau53d]. change [Gau14]. Chaos [MGG^{+}95, BGTJ99]. Chaos/Plaschko/Brod [MGG^{+}95]. Chaotische [MGG^{+}95]. Charge [Hri71, Tch72, CP92, JP28, Pau55a, Pau57a]. charge-free [JP28]. Charles [Dys02, Str04a, Tel03]. che [Fer24]. chemical [Fin49]. Chemie [MGG^{+}95]. Chemie/Dictionary [MGG^{+}95]. chemischen [Fin49]. Chemistry [PW35, Hal86, Ano45a, MGG^{+}95]. Chile [Ano45a]. Chinese [vE98]. Chiral [Xue91]. Chiral-gauge [Xue91]. CII [RC25]. circuits [Pau74]. class [CC83, CCS95]. classic [EFK90]. Classical [Pau33c, SP77, GG05]. Closed [Ell48, Goe48, Goe49, KK73]. Closing [Pau56c]. cloth [vB10]. Cloud [Wei88]. cluster [FRKK95, SOCg82]. cohesive [Fre28]. coincidence [Cas92, Fis99]. collaboration [Sch01, vM05a]. Collected [BDR84, BDR93, Cha68, KW64, Pau64a, Bal65]. collection [Eur69]. Collective [HEB^{+}80]. collision [Pau24d]. combination [Pau25e]. combined [SKS85]. comet [PB27]. commemoration [FW59]. Comment [Cav00, CNT00, PT59]. Commentary [Got83]. Comments [AP80, Pau24d]. Commission [BT58, DT58]. Company [Bro86, Che85, vB10, Coo03]. compared [Pau40b]. Complementarity [Pau94m, De 14, Pau48a, Pau50b, Scu07c]. complete [FRKK95]. Completed [Hei79a]. completion [Pau25c]. complex [Pau23c, Pau25c, Pau58d, vO69]. Complexes [SZ97]. component [Pau19b]. Components [QS98]. Composite [GH99, KS81]. computing [HGF91, ZF00]. concentration [CP84]. Concept [Gar63a, Pau48a, Pau50b, Pau52a, SSF82, vdWPR52]. ConceptDraw [GRE^{+}01]. conception [Lau85b, Meh73]. concepts [Enz85a]. Concerning [VNB08]. Condensed [ARK^{+}99]. condition [Kij72]. Conference [Fis99, Pet11, Pau30d, Far01, Pau56c, Pau01]. conferences [Meh75]. Confirmation [CRH^{+}56]. conflict [vM85]. Conformal [Hri71]. Congruence [Nel72]. conjecture [AF14]. Conjugation [Hri71, CP92]. connected [Pau56a]. Connection [Bur58, CP92, LZ58, Pau40a, Del61, Pau50a]. Conrad [MGG^{+}95]. conscience [Dys02, EM88]. conseil [CCJ^{+}34]. Consensus [Jen00]. Conservation [GLR14, Pau57a]. conserved [CV02]. considerations [Pau38b, Fer24]. Considerazioni [Fer24]. Consistency [GH75]. Constant [VNB08, Pau21a, PM26, Sch99a, Sch99b]. constrained [BP22, BP63]. constraint [MS94]. contain [Fer24]. Contemporary [AvMMH^{+}07, AP06]. contengono [Fer24]. context [AP06, Pet11]. Continuing [Ber06]. Continuous [Pau56e, Pau65a]. continuum [ST82, ST84]. contractions [HNPT06]. Contribution [Pau94c, Ano85, Pau55b, Pau65b, Pau01]. Contributions [Pau36b, Pau49, Pau94q, Pau48b, Str00, Pau35a]. Control [VNB08]. Controversy [Jen00, Scu07c]. Conversation [BM04]. Cooper [PCC10]. coordinates [Pau33c, Pau33d]. Copenhagen [Lau85a]. Correct [Fre30, KS81]. Corrected [BT58, DT58]. Correction [Ehr27a]. corrections [BCP89, Kal89]. Correlation [QS98, Suz73]. Correlation-Kinetic [QS98]. correlations [KK73]. Correspondence [Bro95, Eis00, Hen86, HvMW79, Mei92, vMHW85, vM93, vM96, vM99, vM01, vM05b, Eis02]. Cosmic [MGG^{+}95, Mil09, vB10, MM11]. Cosmological [ARK^{+}99]. Cosmology [Rei72]. Cosmos [Fis04, Fis11]. Coulomb [QS98]. Coulombic [KM89]. Coupling [PD42a, PH45, Tch72, PD42b, Pau43c]. crack [BS12]. Creation [Bel86, Bro86, Che85, Hen84, Wes85, dB86]. Creativity [MKD^{+}91]. creators [Str11]. criterion [Pau39]. Critical [Sta09]. Critique [Ano05]. crystals [Pau25d]. Curies [Kae48]. curious [Gau14]. current [Hei84, Pau53d].

D [Bro86, Che85, KP55]. D. [Pau56d]. dans [Pau53b]. David [OBxx]. Debye [Pau35b]. Decay [Jen00, Gau14, Pau38b]. Déchirure [Pau72b]. Deciphering [Mil09, vB10]. Decomposition [Pec93]. Dedicated [Mor56a, KP55, PRW55]. Defect [HEB^{+}80]. Defects [MGG^{+}95]. Defects/Weiss [MGG^{+}95]. define [Boy14]. Defining [ARK^{+}99, De 14]. Definition [SG78]. deflection [Pau19a]. Deformation [Pec93]. deformed [JMS80]. degeneracy [Pau27a]. Degeneration [Pau90b]. degli [Fer24]. deinen [MKD^{+}91]. d’Einstein [PS32a, PS32b]. del [Maj37]. dell [Maj37]. della [Pau58e]. dello [Gul54]. demon [Scu07a]. Demonstration [Ein41]. demonstrative [Mas04]. Denken [Pau56f, Pau59]. Denkens [Sti77]. Denkstoffe [Fis04, Fis11]. densities [BCP89, Kal89, ZY88]. Density [SZ97, KM89, KK73]. density-matrix [KK73]. dependence [Pau25f]. Dependent [PH45, BBMS08]. depth [Gie05]. detect [Nav06]. Detection [CRH^{+}56, RC53]. Determination [GR31]. determining [Pau47c]. Deterministically [Hei79a]. deterministische [Hei79a]. deuteron [BSI^{+}82]. Deutsch [GRE^{+}01]. Deutsch-Englisch [GRE^{+}01]. Deutsch-Englisch/ConceptDraw [GRE^{+}01]. Deutschland [MGG^{+}95]. Deutung [Pau24a]. Development [All56, Meh73, Mor56a, Pau94a, Meh75, PRW55, Pau58a, PB58]. developments [OS00]. DF1 [Che85]. Dialog [APWB95, Fis99, MGG^{+}95]. Dialogue [Che85, dB86, Bel99, Fis99, Gie05, Hen84, Bro86, Wes85, Bel86]. diamagnetism [Pau20b]. Diamagnetismus [Pau20b]. diatomic [Pau21a].

Dictionary [MGG^{+}95]. Did [Fin92]. dielectric [Pau21a, PM26]. Dielektrizitätskonstante [Pau21a, PM26]. Dienstakten [EGO97]. differential [Nag10]. Difficulties [GM89a, GM89b, Pau47b]. Diffraction [Pau38a]. Diffusion [ARK^{+}99]. Dimension [Pau24d]. dimensional [Bes00]. dimensions [ASG05, Hal91, PZ88]. dipole [Pau21a, PM26]. Dipolgase [Pau21a]. Dipolgasen [PM26]. Dirac [Lak96, Pau30c, ASG05, BDQR11, Bel94, CV02, Dit15, GH75, Hor11, Pau32c, PS32a, PS32b, Pau33d, Pau35a, Pau36b, Pau40b, Pau43b, PJ44, Pau46a]. Diracs [Pau32c, Pau46a]. Dirac’schen [Pau35a, Pau40b, Pau33d]. direct [Han10]. Discours [Pau47d]. discoveries [Bra09]. discovery [Gou71, Kae48]. Discussion [Pau34, Pau53c]. discussions [CCJ^{+}34]. disk [BGTJ99]. disregarding [Cav00, CNT99, CNT00]. dissertation [vM00]. Dissidents [Fre15]. dissociated [Pau58h]. dissoziierter [Pau58g, Pau58h]. distorted [Bad11]. Distribution [Sto24, ZSJY82]. DM [Hen86, Sch01]. do [Mas04]. Domenico [Sch01]. Dordrecht [Che85]. Dordrecht/Boston/Lancaster [Che85]. Doublets [BU26]. doubly [JMS80]. doubly-even [JMS80]. d’ouvrages [dB86]. Draft [Fer34]. dreams [Dys02, Fis04, Fis11, Haw11]. due [ATY92]. durch [Pau24a, Pau56c, Pau70, Pau90a, UG25]. during [vM05a]. Dürr [GRE^{+}01]. Dutch [Cas92, GU26, vE95]. Dynamics [Eck26, BNP05, HP29, HP30, Pau24d]. Dynamik [MGG^{+}95, Pau24d].

E. [Pau24b, Pau25a]. each [UG25]. Earlier [Pau94j, Pau91]. Early [Nav06]. Earth [HEB^{+}80]. Easy [Sta07]. Eckart [Eck70, Eck73]. Eddington [Pau26a]. Edited [Eis00, All56, Mor56a, Sch01]. Editor [DGB^{+}01, Pes05]. Editorial [Pau48a]. editors [YK73]. Edmund [Nau13]. eeuw [Cas92]. Effect [BU26, GB29, Hel67, JMS80, BG28, BBMS08, GS04, GB30, Pau23b, Pau25f, Pau47c, ZY88]. effective [SKS85]. Effects [Aoi71, Aoi73, PR36]. Effekt [GB30]. ego [PB58]. Ehrenfest [Pau33b, Pau94k]. Eidgenössischen [EGO97]. eigenfunctions [Pau39]. Eigenfunktionen [Pau39]. Eigenschaften [Hei84]. eigentlich [Bel86]. Eigenvalues [Bes00]. Eighteen [Got83]. Ein- [Pau39]. einatomigen [Fer26]. einatomiger [Pau20b]. einem [Pau24d]. Einfluss [JP52, Krö54, Pau52b, Pau25f]. Einige [Pau33a]. einiger [Pau24a]. Einstein [Ano05, ARK^{+}99, Bro95, Eis00, GRE^{+}01, HEB^{+}80, Hen86, HvMW79, LM85, MGG^{+}95, OSWR86, Pau24d, Pau58a, SH99, vMHW85, vM93, vM96, vM99, vM01, vM05b, Deg89, CG14, FW07, Kle05, PS32a, PS32b, Pau36a, Pau49, Pau55b, Pau58a, Pau61c, Pau65b, Pau94a, Pau94c, Pau94f, Sch49a, Sch49b]. Einsteina [Pau65b]. Einsteins [Pau55b]. Einsteinschen [Pau24d]. einzelnen [UG25]. Elaboration [KMS80]. elastic [SHS81]. Elected [BC51]. electric [Pau25e, Pau33f]. electricity [Pau19c]. Electrodynamics [Hei79b, Hei79a, Pau73a, PE00a, Sch58, Sch03, Esp07, JP28]. electromagnetic [FP39b, FP39a, SG78]. Electron [BU26, PE57, Pau75b, SB53, Tho26a, Fre28, GU26, Gou71, Maj37, Pau24d, Pau25f, Pau27b, Pau32c, Pau32d, RL77, UG25, SB33]. electronic [Pau74]. Electrons [BBB^{+}10, PBB^{+}11, OSWR86, RS74, Sto24, UG26, ATY92, Boh26, NFG^{+}91, Pau23a, Pau25c, Pau33h, Pau73b, PE00b, SB53]. elektrischen [Pau25e]. Elektrizität [Pau19c]. elektromagnetischen [FP39b]. Elektron [Pau24d]. Elektronen [Pau23a, Pau33h]. Elektronengastheorie [Fre28]. Elektronengruppen [Pau25c]. Elektronenmasse [Pau25f]. Elektronentheorie [PE57, SB33]. elektronische [Pau74]. Elektrons [Pau32c, Pau27b, UG25]. element [Pau40b, Rig02]. Elementare [Pau30b]. Elementarteilchen [Hei84]. Elementary [BH83a, HEB^{+}80, Hei84, HP93, Pau41, Pau93, BH82, PB40, PT59]. elementary-particle [BH82]. Elemente [Fin49, Pau52c]. elementi [Fer24]. Elements [Pau94o, Fer24, Fin49, Pau55c, dGPS00, Pau52c, Pau56d]. elettrone [Maj37]. Elliott [LT91, LT97, LT01, LT05]. Elster [MKD^{+}91]. Emission [Dir27, FP38, PF38, Pau46a, PJ44]. end [FW07]. Energiekomponenten [Pau19b]. Energies [Pec93]. Energy [GLR14, ET60, Pau19b]. engineering [BPP64]. Englisch/ConceptDraw [GRE^{+}01]. enhanced [BLV05]. Enrico [Mil07, Ore01]. entanglement [CMM^{+}07, Gil09]. Entropie [Pau28]. entropy [Pau28]. Entwicklung [Hei60, Pau58a]. envision [Gau14]. Enz [Coo03, Dys02, Str04a, Tel03, YK73, Enz00]. episodes [Bra09]. epistemological [Pau54b]. Epistemology [Pau94e, Pau61b, Pau61a]. Equation [ARK^{+}99, Pag65, ASG05, BBMS08, FF87, Mie81, NN99, Pau32c, PW34, RL77, Raj76, SS75]. Equation/Gaponenko [ARK^{+}99]. Equations [Eck26, DC86, EP43, FP39b, FP39a, MS94, PS32a, PS32b, Pau33d, Pau40b, Sza97, Zha02, PS32a, PS32b]. equidistant [BCP89]. equilibrium [Pau23a, Pau58h]. equivalence [JW28]. Era [MGG^{+}95]. Erforschung [Fis10, Fis12, MGG^{+}95]. Ergänzung [Hei79a]. Ergebnisse [Pau29b, Pau30a, Pau32a, Pau33e]. Erinnerungen [Bet00, Hei60, vW59]. erkenntnistheoretische [Pau54b]. Erkenntnistheorie [Pau61b, Pau61a]. Erklärung [GRE^{+}01]. Erkundigungsfragen [Pau33a]. Erlebtes [Pau90a]. Ernst [Pai00a, Pei60b, FS06, Smu90]. erscheinenden [Pau25e]. Ersetzung [UG25]. erzählt [FS06]. Erzähltes [Pau90a]. Essay [Coo03]. Essays [Mor56a, Lau97, PRW55, SH99, Pau61b, Pau61a]. essential [Rig02]. estimates [BLV05]. était [Ano05, Kle05]. Etat [Pau53d]. etc [Sch01]. ETH [EGO97]. Ettore [EAW15]. Even [Ber06, JMS80, Woi06]. evidential [Fra04]. evolution [Fle07, Hei60]. Exact [BCP89, Mie81, NM91, vM87]. exakten [Pau29b, Pau30a, Pau32a, Pau33e]. Exchange [QS98]. excited [JMS80]. exciton [ZSJY82]. Exclusion [Bar10, BBB^{+}10, CM30, PBB^{+}11, Ehr27b, Ehr27a, Fre30, Hal86, JBE^{+}00, Kap13, KLL90, Ken30, Kra09, LL79, Mar44, Pau55a, Pau64b, Pau94d, PCC10, RS74, Sta07, YS93, Fle07, Mas05, NFG^{+}91, ZY88, Pau46c, Pau47d, Pau77b, Rus24, Str04b, Str07]. exclusive [Kap02]. exerted [PB27]. exhibit [YS93]. Exist [Gar63a]. Existence [Cha32b, Cha32a, Ein41, EP43]. existing [Gau14]. Expanding [BSI^{+}82]. Expansions [FM68b, FM68a]. experienced [Pau90a]. Experiment [PBB^{+}11, Pau94s, Pau52d, Pau52d]. Experimental [Bar10, BBB^{+}10, PBB^{+}11, JBE^{+}00]. Experts [Ano45a]. Explanation [BU26]. exploration [Pau33a]. Extended [NN99]. extension [BDQR11, vM00]. extensions [BP03]. external [KM89, Pau23c, Pau58h, SG78]. Extraordinary [Gib19].

F [Bal65, HW87]. F. [PT59]. faces [CM09]. Fachschulen [BPP64]. Fachwörterbuch [GRE^{+}01]. Factors [Tch72, ET60]. Fail [GLR14]. Failure [Ber06, Woi06]. Faktoren [ET60]. Falk [MKD^{+}91]. Fall [MGG^{+}95, Pau40b]. Faust [MGG^{+}95]. Feinstrukturen [GB27]. Feld [FP39b]. Felde [Pau25e]. Felder [JP28, Pau24a]. Feldern [Pau23c]. Feldgleichungen [FP39b]. Feldquantisierung [Pau46a]. Feliu [GHMP87]. Fermi [Bel94, Fre28, Mil07, Ore01, ZSJY82]. fermion [MS94]. Fermischen [Fre28]. Festschrift [Pau01]. Few [MKD^{+}91]. Few-Body [MKD^{+}91]. Feynman [Cav00, CNT00, CNT99]. fiction [Cas03]. Field [Dys60, GM87, GM88, GM89a, GM89b, Pau41, PD42a, Pau43a, Pau43b, PK43a, PK43b, EP43, Enz85a, FP39b, FP39a, KP55, Łop91, Pau19b, Pau43c, PJ44, Pau46a, Pau47b, Pau53d, Pau53a, PK56, Pau73c, PE00c, SG78, SB53]. field-quantization [PJ44]. fields [Ein41, HP29, HP30, JP28, Pau23c, Pau24a, Pau25e, Pau56a, Pau58h, SS75]. Fierz [Ara90, BLV05, BNP05, DC86, HSS11]. Fifty [Gou76, Kae48, Uhl76]. Figure [vE98]. Filme [HEB^{+}80]. filosofia [Lau82]. Fine [PST06, RC25, VNB08, Han10, Han11, HPPT02, GB27]. finest [PZ88]. Finland [LM85]. Finn [Ano45a]. Finnish [Lau80, Lau82]. fireballs [BSI^{+}82]. First [MGG^{+}95]. Fischer [FS06]. five [Pau33c, Pau33d, Pau77b]. Fixer [MGG^{+}95]. Fixer/Wenske [MGG^{+}95]. fiziki [PB58]. flash [Gar63b]. flip [Hel67]. Fluctuations [Aoi73, Aoi71]. Fluid [GR31]. Fock [FRKK95]. Fois [Ano05, Kle05]. Fölsing [MGG^{+}95]. For.Mem.R.S. [Ano53, Kem59]. forbidden [SHS81]. Force [Fie39, Pau58h, UG25]. Force-Free [Fie39]. Forces [Pau46b, Fre28, Bor47]. Forderung [UG25]. Foreword [Boh60]. Form [Rus24, Tch72, MS94, NM91]. formation [Gau14, Krö54, Pau52b]. Formulation [Fre30, BCP89, GH75, KK73, Pau33c, Pau33d]. Formulierung [Pau33c, Pau33d]. Found [Sul62]. Foundations [Fre15, LM85, Pau47a, Ble91a, Ble91b, GR00, Hal86]. Four [Ada72]. Fractals [MGG^{+}95]. Fractional [Hal91]. Frage [Pau23c, Pau24a]. Franklin [Sta07]. Frederick [JG04]. Free [CRH^{+}56, Fie39, RC53, Ein41, JP28, Pau23a, Pau33h]. freien [Pau23a]. freier [Pau33h]. French [CCJ^{+}34, Kle05, Pau32d, PS32a, PS32b, Pau36b, Pau36a, Pau47d, Pau53d, Pau53b, Pau72b]. Frenkel [ARK^{+}99]. frequencies [Pau46a]. frequency [PJ44]. Frequenzen [Pau46a]. Friction [HEB^{+}80]. Friedrich [ARK^{+}99, GRE^{+}01, MKD^{+}91]. Friend [Enz94a]. Friendship [Mil09, vB10]. Fritzsch [MKD^{+}91]. Functional [SZ97]. Functions [Pau38a]. fünf [Pau33c, Pau33d, Pau77b]. fusion [Rei72]. Future [Hei79b, Hei79a, Pet11].

G [Dys60, Fis99, Gie05, HW87, JP52, MM11]. G. [Pau32b]. Gabriela [Ano45a]. Galilei [MGG^{+}95]. Gallery [Pai00b]. game [vE95]. Gamow [Pau32b, Rei72]. Gaponenko [ARK^{+}99]. gaps [Bes00]. gas [Fer26, Fre28, McC86, Pau27a, Pau90b]. Gase [Pau20b]. Gasentartung [Pau27a]. gases [Pau20b, Pau21a, PM26, Pau73e, PE00e, Fer26]. Gauge [OS00, Xue91]. gebunden [Sch01]. Gedankenexperiment [LM85]. gegenüber [Pau40b]. geht [Pau70, Pau90a]. Geist [vEA02]. Geisteswissenschaften [MKD^{+}91]. Geisteswissenschaften/Vischer [MKD^{+}91]. General [BT58, DT58, Hei84, Pau80, HGF91, Pau33g, Pau58b, ST82, ST84, ZF00]. generalization [Hal91]. Generalized [Raj76, Han11, HPPT02, PST06, Suz73, vO69]. generated [PST06]. Generation [Hür22]. generations [Ada72]. generator [dGPS00]. Genius [Pai00b, Gar63b]. geometric [Pau32c]. Geometrie [BPP64]. geometrische [Pau32c]. Geometry [DH91, BPP64]. George [Rei72]. German [Dys60, Eis00, Ano93, APWB95, BG28, BSRD^{+}55, BPP64, Ble91b, BDR84, BDR93, BP22, Bor63, BP63, Cha68, Ehr31, Ein22, EFK90, Eis91, Eis02, ET60, EM88, EGO97, Enz05, Fer26, FP38, FP39b, Fie39, Fin49, Fis99, Fis04, FS06, Fis10, Fis11, Fis12, Fre28, GB27, GB30, GP31, HP29, HP30, Hei60, Hei79b, Hei79a, Hei84, JW28, JP28, KP23, Krö54, LPB^{+}56, Mei92, MM11, Pau19b, Pau19a, Pau19c, Pau20a, Pau20c, Pau20b, Pau21c, Pau21b, Pau21a, Pau22, Pau23c, Pau23b, Pau23a, Pau24d, Pau24a, Pau24c, Pau25d, Pau25f, Pau25e, Pau25c, PM26, Pau26b, Pau26c, Pau27a, Pau27b, PB27, Pau28, Pau29d, Pau31a, Pau32c, Pau33g, Pau33a, Pau33c, Pau33d, Pau33h, PW34, Pau35a, Pau36c, PF37, PF38, Pau39, Pau40b, Pau46a, Pau47c, Pau47e, Pau48a, Pau48b, Pau50b, Pau52a, Pau52b]. German [Pau52c, Pau52d, Pau54b, Pau54c, Pau55b, Pau56c, Pau56f, Pau57c, Pau57b, PE57, Pau58a, Pau58b, Pau58h, Pau58c, Pau59, Pau61d, Pau61b, Pau61a, Pau61c, Pau70, Pau74, Pau77b, Pau85a, Pau90a, PG00, Pau01, BS12, RP79, Rös91, Sch01, SB33, Tho26b, UG25, ZBG30, vExx, vEA02, vdWPR52, vM68, vM05a, vW59]. Gershenfeld [ARK^{+}99]. gesagt“ [Enz05]. Gesammelte [BDR84, BDR93, Cha68]. Geschichte [MGG^{+}95, Pau52c, Pau57c, Pau61d]. Geschwindigkeitsabhängigkeit [Pau25f]. Gesetzen [Pau58c]. Gesetzmäßigkeiten [Pau23b]. gestörter [BP22, BP63]. Gets [Ano45a]. gewesen [Bel86]. Gewissen [EM88]. ghost [Hei56]. Giulini [Sch01]. given [Pau47d]. Gleichgewicht [Pau23a]. Gleichgewichtsgemische [Pau58g, Pau58h]. Gleichungen [Pau33d]. God [vE95]. Gödel [FW07]. Goetzberger [MGG^{+}95]. Going [Sta09]. Gordon [CV02]. Göttingen [Got83]. Göttinger [GRE^{+}01]. Goudsmit [Lak96]. Graded [HNPT06]. grading [Han11, HPPT02, PST06]. gradings [Han10, PZ88]. Graßmann [Pet11, Hor11]. Gravitation [Pau19c, Pau19a]. gravitational [Ein41, Pau19b, Sch99a, Sch99b]. Gravitationsfeldes [Pau19b]. Gravitationstheorie [Pau19a]. gravity [GG05]. Great [Ano58, Kae48, Lin04]. Greatest [Hür22]. Grenzen [Hei84]. Gröbner [Ano93]. Grosse [ARK^{+}99]. Großforschung [MGG^{+}95]. ground [KK73, NM91]. Group [Far01, HP93, PT59, Pau93, Han11, Pau55a]. groups [Han10, Pau25c, Pau56e, Pau65a]. Grundkurs [MGG^{+}95]. Grundlage [GRE^{+}01]. Grundlagen [Ble91b, Pau29a, Pau31b]. guide [Boy14]. Guíxols [GHMP87]. Gürsey [PT59]. Gustav [Mei92].

H [LT91, LT97, LT01, LT05, Pau24d, Pau33f]. H. [Pau29c, Pau47a]. Haar [Pau56d]. Haas [HEB^{+}80]. Hahn [GRE^{+}01, MGG^{+}95, Ano45a]. Hahn-Meitner-Instituts [MGG^{+}95]. Hahn-Meitner-Instituts/Fölsing [MGG^{+}95]. Halbleiterbauelemente [OSWR86]. Halbleiterbauelemente/Stitch [OSWR86]. half [Cas83, Cas92, CS10, IiS03]. half-century [Cas92]. halve [Cas92]. Hamiltonian [BG88, BP03, ET60, FWS^{+}85, FM68b, FM68a, HGF91, Nou06, Pau53a, SZ97, ST82, ST84, ZF00]. Hamiltonians [CC83, CCS95]. Hamiltonoperator [ET60]. Handbook [OSWR86]. hänen [Lau80]. Haphazard [Cas83, CS10]. hardback [Coo03, Sta07]. harmony [Fis04, Fis11]. harvest [Bra09]. Havlin [MGG^{+}95]. HC [Pau24d]. heart [Pau70, Pau90a]. Heat [Den27]. heaven [Cas03]. Heidelberg [Eis00, Hen86]. Heisenberg [ARK^{+}99, BDR84, BDR93, Bro95, Eis00, GRE^{+}01, HEB^{+}80, Hen86, HvMW79, Lak96, MGG^{+}95, OSWR86, Pau31c, vMHW85, vM93, vM96, vM99, vM01, vM05b, BC09, CP84, CG14, Hei79b, Mas04, Mie81, Pau34, Ser77, dR99, vM05a]. Heisenbergs [Hei79b, vM05a]. Heitjans [ARK^{+}99]. Held [Meh73, CCJ^{+}34, GHMP87]. helical [SS75]. Hendry [Che85, Wes85, Bro86, dB86]. Henrik [OBxx]. Herausgegeben [Eis00, Sch01]. Here [Ano45a]. heretics [DT18]. Hermann [HEB^{+}80, HW87, Pau19c, Cha68, Pau19c]. Hertha [BS12, BS12]. Herz [Pau90a, Pau70]. Heuristic [dR99]. Hi [MKD^{+}91]. Hi-Tech [MKD^{+}91]. Hi-Tech/Scribal [MKD^{+}91]. hidden [BC09, Pau53b]. Hierarchy [CV02]. Hilbert [Hei56]. Hintertreppe [Fis10, Fis12]. Historical [OS00, Str04b, Str07, GG05, GR00, GP00]. History [Enz94a, GHMP87, HW87, Pau46c, Pau85b, Pau94j, Wei88, dB86, Bag11, Fra04, KHFA67, Pau52c, Pau57c, Pau61d, Pau91]. Hochschule [EGO97]. hole [BCP89, Kal89, ZY88]. homogenen [Pau33c, Pau33d]. homogeneous [Pau33c, Pau33d]. honor [SH99]. Honorary [BC51]. Honored [Ano45a]. Houtermans [ARK^{+}99]. Humanökologie [MKD^{+}91]. Hund [Boy14, Bon81]. Hydrogen [Den27, Lak96, Rig02, Pau21c, Pau22, Pau26c]. Hyperfeinstruktur [GP31, ZBG30, GB30]. Hyperfine [GB29, Bel82, GB30, GP31, ZBG30]. Hypothese [UG25]. Hypothesis [Pau85a, UG25, Gau14].

Idea [Pau94m, Bro78]. ideal [Fer26]. idealen [Fer26]. Ideas [AvMMH^{+}07, GHMP87, Pau94g, Pau94e, AP06, Ble91a, Ble91b, Krö54, Pau52b, Pau54b, vM68]. Idee [Pau48a, Pau50b]. Ideen [Ble91b, Pau54b, vM68]. identical [Fer24, Mil07]. identici [Fer24].

Ignatovich [MKD^{+}91]. ihre [Pau74]. ihrer [Pau24a]. II [ARK^{+}99, BT58, DT58, Eis00, HW87, Hei84, OSWR86, Pau33d, vMHW85, vM99, Goe49, GP31, HP30, Hei56, Hei84, LD68, Pau33d, SB53]. III [BDR93, BT58, Bro95, DT58, Eis02, Hei84, vM93, vM01, Hei84]. im [ET60, FP39b, Fis99, Pau25e, Pau25c, Pau40b]. impact [AF14]. Impenetrability [Ehr27b, Ehr27a]. implementation [Bad11]. implications [RC88]. Importance [Mar44, SOCg82]. important [Str00]. Impressionen [Pau61c]. Impressions [Pau94f, Pau61c]. incarnation [vE95]. including [Ano64, SS98]. incorporating [ST82, ST84]. increase [Pau28]. indefinite [Pau58d]. independent [Kap02]. induction [Mas04]. inequalities [BH09, Suz73]. inequality [CP84]. Infinitesimalrechnung [BPP64]. Influence [Pau94g, Krö54, Pau24a, Pau25f, Pau52b, SKS85, ZSJY82]. Information [Bon81]. Ingenieur [BPP64]. Ingenieur- [BPP64]. inhibited [CP93]. inner [UG25]. Innere [MGG^{+}95]. inneren [UG25]. innermost [Gie05]. innocence [FW07]. Inquiry [Gib19]. ins [MGG^{+}95]. Insight [Wei91b]. institut [CCJ^{+}34]. Institute [CCJ^{+}34, Ano45a]. Instituts/Fölsing [MGG^{+}95]. integrals [HGF91, ZF00]. intelligenten [GRE^{+}01]. Intensität [Pau33h]. Intensitäten [Pau25e]. intensities [Pau25e]. intensity [Pau33h]. Interaction [DET74a, DET74b, HEB^{+}80, Pec93, Hei84, KM89]. Interactions [PH45, KS81, KMS80, PD42b, SKS85]. interferometers [YS93]. interior [Pla89]. intermediate [Cav00, CNT99, CNT00]. International [BT58, DT58, GHMP87, Meh73, CCJ^{+}34]. Interpretation [CM30, Hug89, JP55, Pec93, Bon81, Pau24a, vM85]. Interpretations [Deg89]. Introduction [Enz94b, PW35]. invariance [Pau40b]. Invariant [PV49]. Invarianz [Pau40b]. invented [Enz82]. inventory [KHFA67]. Ion [HEB^{+}80, Pau21c, Pau22]. ions [NM91]. ISBN [Bro86, Che85, Coo03, Dys02, Hen86, Str04a]. Isospin [HP93, Pau93]. Isotopic [Pau53c]. Isotropico [Gul54]. Issues [Wes85, Pau33a]. Italian [Fer24, Fer34, Maj37, Pau58e]. Italy [GR00, Meh73]. iuber [Pau51e, Pau51d]. IV [ARK^{+}99, BT58, DT58, Eis00, Eis02, Pau51e, Pau51d, vM96, vM99, vM01, vM05b, Far01]. IV/Teil [ARK^{+}99].

J [HW87]. J. [OBxx, Pau33f]. ja [Lau80, Lau82]. Jahr [Pau85a, Pau01]. Jahre [RP79, Sti77, vM05a]. jedes [UG25]. Joensuu [LM85]. John [Bro86, Wes85, dB86, Che85]. Jonathan [Ber06]. Jordan [Pau30b, Deg89, Sch99a, Sch99b]. journalist [BS12]. Journalistin [BS12]. Journey [FR13b]. Joy [Wei91b]. June [LM85]. Jung [APWB95, AF14, Fis99, MGG^{+}95, Mei92, MM11, Pes05, Sch02, APWB95, Fis99, GP00, Gie05, Lin04, Mei92, MEF01, Mil09, Mil10, MM11, vE95, vB10]. Jung-Dialog [APWB95]. just [EFK90].

K. [vM90]. kabbalists [Kev00]. Källén [Hei56, Nel72, Pag65, Raj76]. Kärger [ARK^{+}99]. Karl [Eis00, HW87, Sch01, Bro95]. Kastler [Ano85].

Kausalitat [Pau36c]. Kay [MKD^{+}91]. Kepler [JP52, Krö54, Pau52b, Pau94g, Krö54]. kernel [Gie05]. Kernmoment [BG28]. Kernphysik [Fis04, Fis11, Pau35b]. Kinetic [QS98, Pau73e, PE00e]. Kirche [MGG^{+}95]. kirjeensä [Lau80]. Kirschner [OSWR86]. Klassiker [EFK90]. Klassische [Pau33c]. Klein [CV02]. kleiner [Pau46a]. kleinsten [Fis10, Fis12]. Knobloch [MGG^{+}95]. Knowledge [Sta07, Cas03]. known [PB40]. Koch [HEB^{+}80]. Kohäsionskräfte [Fre28]. Kohn [QS98]. Kombinationslinien [Pau25e]. Kometenschweifen [PB27]. kommentiert [Sch01]. Komplementaritat [Pau50b, Pau48a]. Komplexstruktur [Pau25c]. Komplexstrukturterme [Pau23c]. Konferenz [Pau56c]. Koordinaten [Pau33c, Pau33d]. Korrespondenzprinzip [Pau24b]. kosmischen [MM11]. Kosmos [Fis04, Fis11]. kräftefreier [Fie39]. Kraftfeldern [Pau58g, Pau58h]. Kraus [GRE^{+}01]. Kristallen [Pau25d]. Kriterium [Pau39]. Kronig [Bal65]. Kunz [HEB^{+}80]. kvantovoi [Pau75a, Pau77a]. kvantovuyu [Pau65b].

L [All56, Mor56a, Pau24d]. Laboratory [Nav06]. ladungsfreier [JP28]. Lancaster [Che85]. Lanczos [FF87]. Langenscheidts [GRE^{+}01]. langwelliger [FP38, PF38]. Laser [OSWR86]. last [Enz00]. laureate [Ano64, SS98]. Laurikainen [MKD^{+}91, vM90, Gie90]. law [Gau14, Sza97, Woi06]. Laws [Ber06, Pau94u, Pau23b, Pau33c, Pau33d, Pau58c]. leap [Fis10, Fis12]. Leben [ARK^{+}99, MGG^{+}95, BS12]. Lecce [GR00]. Lecture [Pau47d]. Lectures [YK73, Ano64, LM74, Pau61b, Pau61a, SS98, MKD^{+}91, PE57]. Lee [KP55, Raj77]. legend [Sta80]. Lehmann [HEB^{+}80]. Lehr [BPP64]. Lehr- [BPP64]. Lehrbuch [HEB^{+}80]. Leipzig [vM05a]. Leipziger [vM05a]. L’électron [Pau32d]. lengths [Nou06]. lepton [Pau57a]. Letter [Ein79, Pau30d, Hei79b]. Letters [DGB^{+}01, Pes05, Sch02, Lau80, MEF01, vM68]. Level [Coo03, NM91]. Levels [Sto24]. Li [GP31, PK56, GP31]. Lichtes [Pau20a]. Lichtquant [Pau24d]. Lichtquanten [Pau24d, FP38]. Lichtquanten. [PF38]. Lie [PZ88]. Lieb [LT91, LT97, LT01, LT05]. Life [MGG^{+}95, Dys02, BS12]. Light [Pau38a, Sta80, FP38, Pau20a, Pau24d, PF38]. Limit [BBB^{+}10, PBB^{+}11, GH99, Ara90, BNP05, KLL90]. Limiting [Pau43a]. Limits [BMT99, Cas03, Hei84]. Lindner [MGG^{+}95]. line [Pau40b]. lines [Pau24a, Pau25e, RC25]. Linienelementes [Pau40b]. Literature [Ano45a]. lives [Coo03]. livre [Ano05]. Local [GM87, GM88, GM89a, GM89b, Pau53a, YS93]. local-realism [YS93]. logic [Boy14]. London [Ber06, Dys60]. long [FP38, PF38]. long-wave [PF38]. long-wavelength [FP38]. Lorentz [Pau29c, Ehr31, Pau31a, Pau55a]. Lorentzmedaille [Ehr31, Pau31a]. Lösung [MGG^{+}95]. Lösung/Meÿenn [MGG^{+}95]. love [CG14]. Low [McC86, PJ44]. Low-temperature [McC86]. lower [Pau46a]. LS [Pau24d]. Lubanski [Łop91]. Ludwig [Eis91, Eis91].

M [Pau30c]. M. [Pau25b, Pau30b, Pau34]. Mach [Smu90]. Made [Pei92, Cli87, Haw11]. Magnetic [BU26, Pau75b, Pau24a, Pau27b, Pau32d, Pau33f, SS75]. magnetique [Pau32d]. magnetische [Pau24a]. magnetischen [Pau27b]. Magnetism [Pau75b, Pau32d]. magnetisme [Pau32d]. magnetized [BGTJ99]. magneton [Pau20c, Pau20c]. Mai [Fis99]. Majorana [Esp07, EAW15]. Make [Fin92]. making [Bel99]. man [Ore01]. Managing [MKD^{+}91]. MAPLE [Bay96]. March [Pau31b]. Martin [ARK^{+}99]. Mass [JBE^{+}00, Ein41, FWS^{+}85, NFG^{+}91, Pau25f, Pau40b, Pau53c]. mass-velocity [FWS^{+}85]. Massimi [Kra09, Sta07]. massless [DC86]. master [FF87]. matematicheskoi [PK56]. Materie [MGG^{+}95, vEA02]. Materie/Brandmüller [MGG^{+}95]. Materiewellen [Pau33d]. Mathematical [ARK^{+}99, BT58, DT58, EAW15, KP55, Pau35a, PK56, Pau36b]. Mathematics [HW87, Pau85b, Scu07b, Ano93, BPP64, Rös91]. Mathematik [Ano93, BPP64, Rös91]. mathématiques [Pau36b]. mathematischen [Pau35a]. mathematischer [Pau26a]. matrices [BH09, Han11, HPPT02, PZ88, Pau35a, Pau36b, PST06]. Matrix [ST84, Kij72, KK73, ST82, dGPS00]. Matrizen [Pau35a]. Matter [ARK^{+}99, DL67, Ehr27b, Ehr27a, LD68, Pau54a, Pau94h, AP06, LT91, LT97, LT01, LT05, Pau33d, Wei70, vEA02]. Matter/Friedrich [ARK^{+}99]. matters [Rei72]. Max [Fis10, Fis12, Fis10, Fis12, MKD^{+}91]. Max-Planck-Medaille [BT58, DT58]. Maximal [BH09]. Maxwell [Scu07a, Sza97]. May [AvMMH^{+}07, Fis99]. Mayer [PS32a, PS32b]. Meaning [APWB95, DT18]. Means [BU26]. mécanique [Pau53b]. Mechanical [Ser77, BP22, BP63, Kap02]. Mechanics [Bel86, Bro86, Che85, Fre15, Hei79a, Hug89, Pau47d, Pau47e, Pau94d, PW35, Scu07c, dB86, vdW67, ACD90, GR00, Hei60, Hen84, PM26, Pau26c, Pau27b, Pau28, Pau33g, Pau33a, PF37, Pau39, Pau53b, Pau56e, Pau58b, Pau64b, Pau65a, Pau73d, Pau73f, Pau80, PE00d, PE00f, Set09, Wes85, dlPA71, vdW68, vM85, Pau47a, Pau30c, Pau56d]. Mechanik [GRE^{+}01, Pau47e]. mechanischer [BP22, BP63]. Medaille [BT58, DT58]. Medal [Pau31a, Ehr31]. media [Pau20a]. Medien [Pau20a]. Mediziner [HEB^{+}80]. Mediziner/Hermann [HEB^{+}80]. Meeting [CCJ^{+}34, GHMP87, Lin04]. mehr [EFK90]. Meier [Pes05, Sch02]. mein [Pau70, Pau90a]. Meitner [MGG^{+}95]. Members [BC51]. Memorial [Ano60a, FW60, Rei72]. Memories [Gle86, Wei85, Hei60, vW59]. Men [Cli87]. menswording [vE95]. Mercer [FW07]. Mercury [Pau19a]. Merkurperihelbewegung [Pau19a]. Meson [PD42a, Pau43a, PK43a, PK43b, Pau46b, Pau43c, Bor47]. mesotron [PD42b]. message [Lau97]. Metal [SZ97]. Metalle [SB33]. metals [SB33, SB53]. Method [Pau43b, Pec93, BDQR11, PJ44, Pau47c, ST82, ST84]. Methode [Pau47c]. methodological [Set09]. metric [Pau58d]. Meyenn [Bro95, Eis00, GRE^{+}01, HEB^{+}80, HW87, Sch01, ARK^{+}99, MGG^{+}95].

Michela [Sta07, Kra09]. might [Gou76]. Miller [vB10]. Millikan [BC51]. Mind [Hür22, AP06]. minds [Lin04]. minimal [Nou06]. Minnesord [Kle59]. Miramare [Meh73]. Mises [Ano93]. Misprints [BT58, DT58]. Mistral [Ano45a]. Mitwirkung [Sch01]. Mixed [PK43a, PK43b]. mixtures [Pau58h]. Mladjenovic [ARK^{+}99]. Model [Joh33, Ara90, BCP89, BNP05, HSS11, KP55, Pau21c, Pau22, Raj77, ST82, ST84, Suz73, ZSJY82]. Modelers [Lak96]. Modeling [ARK^{+}99]. Modeling/Origin [ARK^{+}99]. Modell [Pau21c, Pau22]. Models [Lak96]. Modern [APWB95, Gib19, LM85, Pau94r, Str08, dB86, Bra09, EFK90, OS00, Pau36c, Uns00]. moderne [APWB95, MGG^{+}95, Uns00]. modernen [EFK90, Pau36c]. Modified [BG88, Pec93]. möglich [Hei79a]. molecular [Pau21c, Pau22]. Molecule [Den27]. Molekülions [Pau21c]. moment [BG28, CG14]. Momenta [Kra65]. moments [Bag11]. monatomic [Fer26, Pau20b]. Monte [AvMMH^{+}07]. most [Haw11, Str00]. Motion [Tho26a, ZSJY82]. Mott [MKD^{+}91]. Moving [GR31, Pau20a, Pau33h]. multiparticle [ZY88]. Multispinor [DC86]. Mushrooming [Wei88]. Must [Gar63a]. my [Pau70, Pau90a]. Mystery [FR13b, Scu07a]. mystical [Deg89]. Mysticism [Scu07b]. mystics [Scu07a].

Några [Kle59]. nach [Fin49, MM11, PM26]. Nachtträume [Fis04, Fis11]. Nanocrystals [ARK^{+}99]. Nanocrystals/Gershenfeld [ARK^{+}99]. narrated [Pau90a]. National [RW94]. Natur [MKD^{+}91]. Natur- [MKD^{+}91]. Natural [Pau94e, Boy14, Pau52a, vdWPR52, Pau54b]. Nature [ARK^{+}99, JP55, Meh73, Nat03, Pau33c, Pau33d]. Naturerklärung [JP52]. Naturgesetze [Pau33c, Pau33d]. Naturwissenschaften [Pau29b, Pau30a, Pau32a, Pau33e, Pau52a, vdWPR52]. Naturwissenschaftliche [Pau54b]. naturwissenschaftlicher [JP52, Krö54, Pau52b]. natuurkunde [Cas92]. Neu [Sch01]. neuen [Pau26c, Pau28]. neueren [Pau57c, Pau61d]. Neutrino [Ano56, CRH^{+}56, GLR14, Mor56b, Pau85a, Pau94j, RC53, Sul62, Win91, Bro78, Enz82, Nav06, Pau31d, Pau57c, Pau61d, Pau77b, Pau91, Pau77b]. Neutrinohypothese [Pau85a]. Neutrinos [Gar63a, Nat03, Bon05, Fra04, Pau57c, Pau61d]. Neutron [Cha32b, Cha32a]. neutrons [Bon05, SS75, MKD^{+}91]. Neutrons/Belyaev [MKD^{+}91]. nicht [ET60]. Nichtlineare [MGG^{+}95]. Niels [KP55, Mor56a, OBxx, Pau45a, PRW55, PB58, Pau94i, SM08, All56, Mor56a]. Nil [PB58]. Nil’su [PB58]. NMR [SZ97]. No [Enz02, Sta07, HSS11, Tel03, Coo03, Dys02, Str04a]. no-binding [HSS11]. Nobel [Enz05, Pau47d, Ano45a, Ano45b, Ano64, dVNHdVH53, dVNHdVH71, Pau45b, Pau47d, SS98]. Nobelpreisträgers [Enz05]. Nomenclature [BT58, DT58]. Nomenclature/Additions [BT58, DT58]. Non [GP00, Ein41, EP43, Pau53a, ZF00, Coo03]. non-existence [Ein41, EP43]. non-local [Pau53a]. non-orthogonal [ZF00]. Non-separability [GP00]. non-specialist [Coo03]. non-vanishing [Ein41]. noncommuting [ET60]. nonlinear [Mie81]. nonrelativistic [Ara90]. Norton [vB10]. Note [Den27, BDQR11, Bel82, GG05]. notes [Dys60]. novel [Pau72b]. Noyau [Pau34]. noyaux [CCJ^{+}34]. Nuclear [ARK^{+}99, Jen00, Kra65, MGG^{+}95, Pau46b, Ada72, Bor47, Fis04, Fis11, Gou61, Kae48, PD42b, vO69, BG28]. Nuclei [Ell48, Goe48, Goe49, CCJ^{+}34, JMS80, KK73, vO69]. Nucleon [GRE^{+}01, ACD90]. nucleon-alpha [ACD90]. Nucleon/Sube [GRE^{+}01]. Nucleons [GH99, BMT99, LL79, NFG^{+}91]. Nucleus [DET74a, DET74b, FR13b, DL76, SSF82, SHS81]. nucleus-nucleus [SSF82, SHS81]. Nullpunktsenergie [ET60]. Number [Mil09, VNB08, vB10, MM11]. nur [EFK90]. NUT [CV02].

o [Bro95, PK56]. obéissant [Pau36a]. obeying [Pau36a]. Obituary [Pau33b, Ano59b, Lan59b, Lan59a, Pau51b, Pau51c]. Object [GR31]. Objection [LM17]. Objekten [GRE^{+}01]. Obsession [Mil10]. Occasion [Mor56a, KP55, PRW55]. October [CCJ^{+}34, Ehr31, Far01, GR00]. octobre [CCJ^{+}34]. Oktober [Ehr31]. old [Pau57c, Pau61d]. Olson [HEB^{+}80]. onde [Pau53b]. One [FM68b, FM68a, Gib19, Sta09, Cas03, Pau39]. One- [FM68b, FM68a, Pau39]. only [PD42b]. open [GR00]. Opening [Pau56c]. Operations [BT58, DT58]. Operator [Eck26, BLV05, Bes00, GS04, Hen81b, Łop91]. operators [CV02, IiS03]. Oppenheimer [OBxx]. Optical [ARK^{+}99, Rus24]. Optics [PE57, Pau73b, PE00b, Pau32c]. Optik [PE57, Pau32c, Pau51e, Pau51d]. Optoelektronische [OSWR86]. Oral [Wei88]. orbitals [Boy14, ZF00]. order [GB27]. Origin [ARK^{+}99, Kra09, Sta07, Mas05]. Original [BDR93]. Originalarbeiten [BDR93]. origins [OS00, Set09]. Ornstein [Pau24d]. orthogonal [ZF00]. oscillating [Bon05]. Oscillations [Fra05]. oscillator [Pau47c]. Österreichische [Ano93]. Oszillatorenstärken [Pau47c]. other [MGG^{+}95, Rei72]. Otto [Ano45a, GRE^{+}01]. overlapping [YS93]. Oxford [Coo03, Dys02, Str04a].

P [Coo03, Dys02, Enz00, Pau25a, Pau30b, Str04a, Tel03, YK73]. P. [Pau30c, Pau35b]. Paarläufe [FS06]. pages [Eis00, Sch01, vB10]. Pair [PH45, FS06, PCC10]. pairing [BCP89]. Pais [Pau53c]. paper [Dit15, Pau34]. Papers [KW64, Pau64a, Sch58, Sch03, BDR84, Eur69, Eck70, Eck73, Haw11, OBxx, BDR93, Bal65]. parafermionic [CP93]. Paramagnetism [Aoi73, Aoi71, Pau27a, Pau90b]. Paramagnetismus [Pau27a]. paramètres [Pau53b]. paranormal [Deg89]. Part [BDR93, Eis02, vM96, vM99, vM01, vM05b, Eis00, GP31, Pau33c, Pau33d]. participants [Pau30d]. Particle [ARK^{+}99, BH83a, BH83b, BCP89, BH82, GG05, Gau14, Kal89, KS81, SG78]. particle-hole [BCP89, Kal89]. Particles [Fie39, Hei84, HP93, Pau41, Pau93, FP39b, FP39a, Fis10, Fis12, Mil07, PB27, Pau36a, PB40, PT59, Raj77, SP77, dlPA71]. particules [Pau36a]. partly [SHS81]. Paschen [GB30, GB29, GB30]. Passage [Pei85b]. Passions [Wei91b]. past [Pet11]. path [SH99]. Paul [OSWR86, Enz94a, Pau33b, Pau94k]. Pauli [All56, Ano53, APWB95, ARK^{+}99, AvMMH^{+}07, AF14, Bal65, Bel86, Ble91b, Bro86, Bro95, Cav00, Che85, CNT00, CP84, Dys60, Ehr31, Ein22, Ein79, EFK90, Eis91, Eis00, EM88, EGO97, Enz05, FWS^{+}85, FR13c, Fie59, Fis99, Fis04, Fis11, FM68b, Fre28, GRE^{+}01, HEB^{+}80, HW87, Hei59, Hen86, HvMW79, Kem59, Krö54, MGG^{+}95, Mei92, Mie81, MM11, Mor56a, OSWR86, BS12, Pei60b, Pes05, Sch01, Sch02, ST84, Sta07, Uns00, Wes85, YK73, dB86, vEA02, vB10, vMHW85, vM93, vM96, vM99, vM01, vM05a, vM05b, ASG05, ATY92, AP80, ACD90, Ano45a, Ano45b, Ano58, Ano59a, Ano59b, Ano60a, Ano60b, Ano85, Ano93, Deg89, Aoi71, Aoi73, Ara90, APWB95, AP06, BC09, Bad11, BGTJ99, BCP89, Bar10, BLV05, BG88]. Pauli [Bar79, BMT99, BBB^{+}10, BSI^{+}82, Bay96, Bel82, BNP05, Bes00, Bet72, Bet00, Ble91a, Ble91b, BS73, BT58, BH09, Bon81, BP03, Bor47, BBMS08, Boy14, CC83, CMM^{+}07, CCS95, CM30, Coo03, CV02, CP92, CP93, CNT99, CM09, PBB^{+}11, DT58, Dit15, DC86, DET74a, DET74b, DL76, DPS97, DGB^{+}01, Dys02, Eur69, Ehr27b, Ehr27a, Ehr31, Ein22, EFK90, Eis91, Eis02, Enz73, Enz82, Enz85a, Enz85b, EM88, Enz94a, Enz94b, EvM94, EGO97, Enz00, Enz02, Enz05, Esp07, FW07, FR13a, FW59, FW60, FRKK95, Fin92, Fin49, Fis99, Fis04, Fis11, Fle07, FM68a, Fra05, Fre28, Fre30, FF87, Gar63a, GS04, GP00, Gie05, Gla01, Gle86, Gou61, GM87, GM88, GM89a, GM89b, GH99, GH75, Hal91, Hal86, Han10, Han11, HPPT02, Hei56, Hel67, Hen81a]. Pauli [Hen84, HGF91, HSS11, Hor11, HNPT06, HP70, IiS03, JBE^{+}00, Joh33, JMS80, JW28, Kac74, Kal89, Kap02, Kap13, KLL90, Ken30, Kij72, KS81, Kle59, KM89, Kra09, Kra65, Krö54, KW64, KMS80, Lan59b, Lan59a, Lau80, Lau82, Lau85b, Lau85a, Lau88, Lau97, LM17, LM74, Lin04, LL79, Łop91, Mas04, Mas05, McC86, MEF01, MKD^{+}91, Mey04, Mil09, Mil10, MM11, MS94, NM91, Nag10, NMS84, Nel72, NN99, NFG^{+}91, Nou06, Pag65, Pai00a, PZ88, BS12, Pec93, Pei85a, Pei92, PST06, Pla89, PCC10, QS98, RC88, RL77, Raj76, RS74, RP79, SSF82, SOCg82, SZ97, Sch99a, Sch99b, SS75, ST82, Scu07b, Scu07c, Ser77, Set09, Smo87, Smu90, SHS81, SKS85, Sta80, Str00, Str08, Sza97, Ter65, Uns00, Wei85, Wei91a, Xue91, YK73]. Pauli [YS93, ZF00, Zha02, ZY88, ZSJY82, dGPS00, dR99, vE95, vE98, vExx, vEA02, vM00, vM68, vM85, vM87, vMS01, vM05a, vO69, vW59, Coo03, Dys02, Gie90, Sti77, Str04a, Tel03, vM90]. Pauli-algebra [Bay96]. Pauli-allowed [FRKK95]. Pauli-correct [KS81]. Pauli-equation [SS75]. Pauli-Fermischen [Fre28]. Pauli-forbidden [SHS81]. Pauli-Jung-Dialog [MGG^{+}95]. Pauli-Prinzip [Fin49]. Pauli-Type [SZ97]. Pauli/Falk [MKD^{+}91]. Pauli/Jung [Pes05, MEF01]. Paulis [vExx, vM68]. Paulische [JW28]. Peacock [ARK^{+}99]. Pergamon [Dys60]. perihelion [Pau19a]. Periodensystem [Fin49]. Periodic [Pau94o, Fin49, Pau55c, SB53, Pau52c]. periodischen [Pau52c]. Personal [Uhl76, Wei85, Enz00]. perspective [GP00]. Perturbation [Pau24c]. Peter [Ber06]. Phänomen [Pau57b]. Pharmazeuten [HEB^{+}80]. phase [Suz73]. Phenomena [Pau57b]. phenomenological [EAW15]. Phenomenon [Pau94l]. Philosopher [Hen81a, Sch49a, Sch49b, Enz85b]. Philosopher-Scientist [Sch49a, Sch49b]. Philosophic [Pau47a]. Philosophical [AvMMH^{+}07, Gie90, Mar44, MKD^{+}91, Pau94m, vM90, GP00, Lau88, Pau50b]. philosophische [Pau50b]. Philosophy [dR99, EvM94, Lau82, Lau85a, Sce16]. Photograph [GR31]. photon [Pau24d]. Photonen [Pau46a]. photons [PJ44, Pau46a]. Photovoltaik [MGG^{+}95]. Photovoltaik/Lindner [MGG^{+}95]. Physical [BT58, DT58, HW87, Pau85b, Pau94l, Pau57b, Woi06]. Physicist [Ano58, Pei85b, Wei91b, Enz85b, Meh73]. Physicists [Eis91, Hür22, Ada72, Cli65, Cli87, Kev00]. Physics [Ano45b, ARK^{+}99, BM04, BS73, BT58, BBSR69, BH83a, BH83b, DT58, DH91, Gam85, Gil09, LM85, Meh73, MKD^{+}91, Mor56a, Pau45b, Pau94a, Pau94n, Pau94r, Pau94u, Str08, Wes85, Ano93, Deg89, Ble91a, Ble91b, Bra09, BH82, Cas92, Cli87, Dys02, EFK90, EM88, EvM94, EAW15, FW60, Fis04, Fis11, Gam66a, Gam66b, Gam72, GHMP87, GP00, Gie05, Haw11, dVNHdVH53, dVNHdVH71, Kae48, KHFA67, McC86, Meh75, Pau36c, Pau54c, PRW55, Pau58a, PB58, Pau58c, Pau61b, Pau61a, RG76, SS98, Str00, Str11, Uns00, Win91, Ano64, Ber06, CCJ^{+}34, Pau47d, All56, DT18, YK73]. Physics/Grosse [ARK^{+}99]. Physics/Peacock [ARK^{+}99]. Physik [Ano93, GRE^{+}01, HEB^{+}80, MGG^{+}95, Pau51e, Pau51d, Pau61b, Pau61a, Sti77, Ble91b, EFK90, EM88, MKD^{+}91, Pau36c, Pau54c, Pau58a, Uns00, Pau29c]. Physik-Lehrbuch [HEB^{+}80]. Physik/Vilenkin/Shellard [MGG^{+}95]. Physikalische [Pau57b]. physikalischen [Pau31c, Sti77]. Physiker [Eis91, MKD^{+}91]. Physiker/Kay [MKD^{+}91]. physique [CCJ^{+}34, Pau47d]. pierces [Pau70, Pau90a]. pilot [Pau53b]. pilot-wave [Pau53b]. pilote [Pau53b]. Pion [Bet72, DET74a, DET74b, DL76]. Pion-Nucleus [DET74b]. pioneering [vM68]. Pittsburgh [Sta07]. Planck [BT58, DT58, Fis10, Fis12, MKD^{+}91, CG14, Fis10, Fis12]. plane [IiS03, MKD^{+}91]. Plaschko [MGG^{+}95]. Plato [Scu07b]. platonic [Cas03]. Podolsky [LM85]. Poincaré [IiS03]. point [ET60]. Polarization [PR36, SG78]. Polarized [OSWR86]. Poles [HP70]. politics [Sch99a, Sch99b]. polya [PK56]. Pontecorvo [Bon05]. Portrait [Pai00b, Smo87]. portraits [Far01]. Positron [PR36, Maj37]. positrone [Maj37]. Possible [BU26, Cha32b, GM87, GM88, GM89a, GM89b, ATY92, CP92]. Postscript [Pau93]. postulate [Kap02, Pau31d]. posvjascennyj [PB58]. Potential [QS98, KM89, NM91, Nag10, NMS84, SOCg82, SB53]. potentials [KS81, ST82, ST84]. Pp [Che85, Coo03, Hen86, Sta07, Bro86, Dys60, Dys02, Str04a]. Practice [Sta09, BPP64]. Prague [Smu90]. Präsident [Pau56c]. Prediction [GLR14]. Preface [Enz94a]. Prejudice [RG76]. prescription [Cav00, CNT99, CNT00]. presence [BBMS08, Nou06]. Present [Hei79b, Hei79a]. presentation [Ano64, SS98]. President [Pau56c]. Press [Coo03, Dys60, Dys02, Sta07, Str04a]. pressure [PB27]. price [Dys02, Str04a]. Primas [MGG^{+}95]. Princeton [Ano45a]. Principle [Cav00, CNT00, Kra09, Kra65, Mar44, Pau94d, Sta07, BGTJ99, Bon81, CP93, CM09, Fle07, JMS80, Kal89, Kap02, Mas05, NFG^{+}91, Pau38b, Pau55a, Pau64b, Pla89, SKS85, ZSJY82, ATY92, AP80, ACD90, Bar10, BMT99, BBB^{+}10, Bet72, CMM^{+}07, CM30, CP92, CNT99, PBB^{+}11, DET74a, DET74b, DL76, Ehr27b, Ehr27a, FR13a, Fin49, Fre30, GM87, GM88, GM89a, GM89b, GH99, Hal91, Hal86, HP70, JBE^{+}00, Joh33, Kap13, KLL90, Ken30, KMS80, LL79, McC86, Pau46c, Pau47d, Pau77b, PCC10, RC88, RS74, Rus24, Str04b, Str07, YS93, vO69, Kra09]. Principles [BT58, DT58, Pau30c, Pau33g, Pau58b, Pau80]. Prinzip [JP52, Fin49]. Prinzipien [Pau31c, Pau33g, Pau58b]. prix [Pau47d]. Prize [Ano45b, Enz05, Pau45b, Pau47d, Ano45a, dVNHdVH53, dVNHdVH71]. Prizes [Ano45a]. Probability [Pau94n, Pau52a, vdWPR52, Pau54c]. Problem [Fre28, GG05, PJ44, Pau53b, PCC10]. problème [Pau53b]. problems [Pau56a]. proceedings [GHMP87]. Process [Pau43a]. processes [Hel67, ST82, ST84]. product [Han10, Han11, PST06]. Production [HEB^{+}80, BSI^{+}82]. Production/Olson [HEB^{+}80]. Prof [Ano45b]. Prof. [Ano53, Ano58, Enz00, Kem59]. Professor [ARK^{+}99, Ehr31, Ano38, Ehr31, KP55]. program [HGF91, ST82, ST84, ZF00]. prohibition [JW28]. prononcé [Pau47d]. proof [Kac74]. propagation [Pau20a]. Properties [ARK^{+}99, Hei84, CCJ^{+}34, IiS03]. proposed [KMS80]. propriétés [CCJ^{+}34]. Proved [Kap13]. Pseudoscalar [PD42a, PK43a, PK43b]. Psyche [JP52, JP55]. psychology [Gie05]. Publication [Pau01]. Publisher [Enz94a]. Publishing [Bro86, Che85]. Pure [BT58, DT58]. Pursuit [Mil10, Sta07]. Pythagoras [Scu07b]. Pythagorean [Scu07a].

Quadrapole [Pau74]. quanta [FP38, Pau24d, PF38]. Quantelung [BP22, BP63, Fer26]. Quantendynamik [HP29]. Quantenelektrodynamik [Hei79b, Hei79a, JP28]. Quantenmechanik [GRE^{+}01, Hei79a, Pau27b, Pau33a, Hei60, PM26, Pau26c, Pau28, PF37, Pau30b, Pau31b]. Quantenmechanik/Thomas/Weise [GRE^{+}01]. Quantenphysik [GRE^{+}01]. Quantensprung [Fis10, Fis12]. Quantentheorie [HP30, Pau29a, Pau20c, Pau26b, Pau48b, Pau55b, Pau31c]. quantique [Pau36a, Pau53d, Pau53b]. quantiques [Pau32d]. Quantisierung [PW34]. Quantitative [BLV05]. Quantities [BT58, DT58]. Quantities/Addition [BT58, DT58]. Quantization [Pau43b, BP22, BP63, Fer24, Fer26, PW34, PJ44, Pau46a, Pau47b, Pau73c, PE00c, Pau53c]. quantized [Pau56a]. quantizzazione [Fer24]. Quantum [Bel86, Bel99, Bro86, Che85, Dir27, Eck26, Fre15, Gam85, Gil09, GM87, GM88, GM89a, GM89b, Hei79b, Hei79a, Hug89, Hür22, Pau25a, Pau47a, Pau47d, Pau49, PV49, Pau75b, Pau94c, Pau94d, Pau94q, PW35, Pei60a, Sch58, Sch03, Scu07c, Wes85, dB86, vdW67, ACD90, Deg89, Bag11, Cli65, Cli87, CG14, DT18, Fis10, Fis12, Gam66a, Gam66b, Gam72, GR00, GP00, Gie05, Haw11, HP29, HP30, Hei60, Hen84, JP28, Kap02, KHFA67, Łop91, PM26, Pau26c, Pau27b, Pau28, Pau30c, Pau33a, Pau36a, PF37, Pau48b, Pau53d, Pau53b, Pau55b, Pau56e, Pau64b, Pau65a, Pau65b, Pau80, Scu07a, Set09, Str11, dlPA71, vdW68, vM85, Pau20c, Pau26b, Pau32d]. Quarks [GH99]. quenching [BSI^{+}82]. Quest [DT18, vExx]. Queste [vExx]. question [Pau23c, Pau24a]. questioners [Cli65]. Questions [Scu07c, GR00]. Quon [GH99].

R [Bal65]. R. [BC51]. Radiation [Dir27, HEB^{+}80, BNP05, Pau23a, PB27, Pau29d, Pau33h]. Radioaktivität [Pau32b]. Radium [Kae48]. Raggi [Fer34]. Random [Web73]. Rapports [CCJ^{+}34]. Rarita [DC86]. Rath [GRE^{+}01]. Raum [Pau36c]. rays [Fer34, Pau25d]. razvitie [PB58]. Reactions [Kra65, vO69]. Real [Ano56, DT18]. realism [YS93]. Realität [Pau57b, MKD^{+}91]. Reality [Pau94l, Rös91, Cas83, Cas92, CS10, Eis91, Lau85b, Pau57b]. really [Fra04]. realm [Kae48]. Reason [Ken30]. Reborn [Gil09]. Rebuilding [Fre15]. receiving [Pau47d]. recently [KMS80]. réception [Pau47d]. Reciprocal [Ehr27b, Ehr27a]. Recollections [Pei85b]. Recommendations [BT58, DT58, BT58, DT58]. Recommendations/Additions [BT58, DT58]. Reflection [Pau94u, Pau55a, Pau58c, SS75]. Regge [HP70]. regime [HSS11]. regular [EP43]. Regularization [PV49, Xue91]. Reichenbach [Pau47a]. Reidel [Bro86, Che85, Wes85]. Rejection [Ser77]. Relation [Ehr27b, Ehr27a, Boy14, KM89, Pau25c]. relativiste [Pau36a]. Relativistic [Fie39, Pau41, PV49, SZ97, EP43, FP39b, FP39a, Łop91, MS94, PW34, Pau36a]. relativistische [FP39b, Fie39]. relativistischen [PW34]. Relatività [Pau58e]. Relativitätskongreß [Pau01]. Relativitätstheorie [Pau56c, Sch01, Ein22, Pau21b, Pau26a, PG00]. Relativity [BU26, Pau01, Sch01, Hor11, Pau21b, Pau56c, Pau58e, Pau58f, Pau81, Pau94t, PG00, Dys60, Ein22]. relevant [Pau33a]. Remark [Fin92, Pei92, Pau53c]. Remarks [PR36, Pau46c, Pau56a, Pau20b, Pau56c, Pau53b]. Remarques [Pau53b]. remembered [Bar79]. Remembering [Bet00]. Remembrance [Gla01]. Reminiscence [Mey04]. reminiscences [Kle59, Uhl76]. renormalizable [KP55, PK56]. renormalization [Pau56a, Pau53d]. renormiruyemoi [PK56]. Replacement [UG25]. Reply [CNT00]. Report [PT59, KHFA67, Hei84]. Reports [CCJ^{+}34]. representation [Tho26b]. reprints [Eur69]. requirement [UG25]. residual [Pau25d]. resolution [Gau14]. Response [Pau31a]. rest [Pau40b]. restriction [Kij72]. Reststrahlen [Pau25d]. Review [Bal65, Bel86, Ber06, Bor47, Bro86, Bro95, Che85, Coo03, DT18, Dys60, Dys02, Gie90, HW87, Hen86, Kra09, Mor56a, Pau24b, Pau25b, Pau25a, Pau26a, Pau29b, Pau29c, Pau30a, Pau30b, Pau30c, Pau31b, Pau31c, Pau32a, Pau32b, Pau33e, Pau33f, Pau35b, Pau47a, Pau51e, Pau51d, Pau56d, Pes05, Sch02, Sta07, Sti77, Str04a, Tel03, Wes85, YK73, vB10, vM90, BDR84, Ein22, Eis02]. reviews [BSRD^{+}55, LPB^{+}56]. Revolution [Kae48, Bel99, Kle05, Ano05, Kle05]. Rezension [Eis00, Eis02]. Rezensionen [MKD^{+}91]. Richard [Ano93]. Rigorous [BNP05]. Riss [BS12, Pau70, Pau90a]. Robert [Kem94]. Role [Str04b, Str07, dR99, Pau52a, SHS81, vdWPR52]. Rolle [Pau52a, vdWPR52]. roman [Pau72b]. Röntgen [MGG^{+}95]. Room [Ano60a]. roots [Pau58d]. Rosbaud [Enz94a]. Rosen [LM85]. Rosenfeld [All56, Mor56a]. rotating [GU26]. Rotation [HEB^{+}80]. Rotation/Haas [HEB^{+}80]. roteerende [GU26]. route [Bel94]. Ruhm [MKD^{+}91]. Ruhmasse [Pau40b]. rule [Bon81, Boy14]. rule-set [Boy14]. runs [FS06]. Russell [FW07]. Russian [Pau38b, PK56, PB58, Pau65b]. Rutherford [Lak96]. Rydberg [Pau55c, Pau94o].

S [Pau24d, Pau26a, PB58]. Said [Enz05]. Sallhofer [Lak96]. Sambursky [Sti77]. Sant [GHMP87]. satellite [Pau24a]. Satelliten [Pau24a]. satisfying [Kij72]. saturation [SKS85]. Savart [Sza97]. sbornik [PB58]. Scalar [PH45, SZ97, Esp07, PW34]. Scalar- [PH45]. scattered [Pau33h]. Scattering [Bet72, ACD90, DL76, Hen81b, SSF82, SHS81]. SCF [Pec93]. Schaltungen [Pau74]. scheme [Xue91]. Schicksal [ARK^{+}99]. Schicksal/Kärger [ARK^{+}99]. Schlapp [Kem94]. Schlußwort [Pau56c]. schoolroom [Hor11]. schools [BPP64]. Schriftstellerin [BS12]. Schrödinger [ARK^{+}99, CCS95, Lak96, Deg89, BC09, CC83, GH75, MS94, NN99, vM85]. Schucking [SH99]. Schumacher [HEB^{+}80]. schwachen [Pau23c]. schwarzen [Pau29d]. Schwinger [DC86]. Science [Ano56, APWB95, AvMMH^{+}07, Gib19, MGG^{+}95, MKD^{+}91, Pai00b, Pau94e, Pau94p, Pau94t, RW94, Sta07, Web73, Wei88, dB86, AP06, Cas83, CS10, FW07, FS06, Pau56c, Sce16, Pau56f, Pau59]. Science/Goetzberger/Voß/Knobloch [MGG^{+}95]. Sciences [HW87, Pau54b, Pau85b, Pau52a, vdWPR52]. Scientific [Bal65, BDR84, Bro95, Coo03, Eis00, Enz73, GHMP87, Hen86, HvMW79, Kra09, KW64, Mil10, Pau64a, Pau94g, Sta07, Tel03, vMHW85, vM93, vM96, vM99, vM01, vM05b, Ble91a, Ble91b, BDR93, Cas03, Dys02, Eur69, Enz02, Haw11, Krö54, Mas05, Pau52b, RP79, Str04a, vM68, Eis02]. Scientist [Sch49a, Sch49b]. Scientists [MKD^{+}91, Sce16, Str11]. Scope [Coo03]. Scribal [MKD^{+}91]. search [MM11, Woi06]. Sector [Nel72]. Seems [GLR14, Mor07]. sein [Ble91b, EGO97]. seine [APWB95, Ble91b, MGG^{+}95, Pau52a, vdWPR52]. Seiten [Eis00, Sch01]. Selbstbiographie [MKD^{+}91]. Selbstbiographie/Laurikainen [MKD^{+}91]. selecta [LT91, LT97, LT01, LT05]. Selected [Pau73c, PE00c, Sch58, Sch03, Bor63]. self [BP03]. self-adjoint [BP03]. Semiconductor [ARK^{+}99]. semidesjatiletiem [PB58]. sense [Fis99]. separability [GP00]. separable [KS81, Zha02]. sept [Ano05, Kle05]. September [Ein79, GHMP87, JG04, Meh73, Pet11]. septième [CCJ^{+}34].

Seriendarstellung [Tho26b]. Series [BDR84, BDR93, Pau38a, Tho26b]. Service [EGO97]. set [Boy14]. seven [Kle05, Sce16]. seven-times [Kle05]. Seventh [CCJ^{+}34, Far01]. Seventieth [Mor56a]. Sham [QS98]. Shaped [Gib19]. shell [KK73]. Shellard [MGG^{+}95]. Shells [Ell48, Goe48, Goe49]. Shielding [SZ97]. Shmuel [Sti77]. Shook [Gam85, Gam66a, Gam66b, Gam72, Haw11]. Significance [Pau94m, Pau50b]. similarity [Pau40b]. simmetrica [Maj37]. simple [PZ88]. since [Meh75]. Single [GR31, Hür22, Boy14]. singularities [Ein41]. Sinn [Fis99]. sistemi [Fer24]. Sixty [FR13b]. skalaren [PW34]. Sketches [RP79]. Skizzen [RP79]. slow [Bon05]. small [CP92]. smallest [Fis10, Fis12]. Socialism [RW94]. Society [Gib19]. Software [ARK^{+}99, GRE^{+}01]. solenoids [GS04]. Solid [BM04]. Solid-State [BM04]. Solids [HEB^{+}80, Pei60a]. Solution [Eck26, FF87, Nel72, Pag65, SS75]. solutions [EP43, Mie81]. solvable [CC83, CCS95]. Solvay [CCJ^{+}34, Far01, Meh75, CCJ^{+}34]. Some [Kle59, Pau38b, OS00, Pau24a, Pau33a]. Sommerfeld [Lak96, Pau51e, Pau51d, Ano38, BC51, Pau48b, Pau51a, Pau51b, Pau51c, Pau94b, Pau94q, Set09]. Sommerfelds [Pau48b]. Sonnenenergie [MGG^{+}95]. Sources [KHFA67, Pau85b, vdW67, vdW68, HW87]. sous [CCJ^{+}34]. Space [Pau94r, Enz85a, FRKK95, Hei56, Pau55a, Pau36c]. space-time [Pau55a]. spacing [BCP89]. Spain [GHMP87]. special [Hor11]. specialist [Coo03]. Specific [Den27]. Spectra [Rus24, UG26, Boh26, GU26, KP23, Pau25c]. Spectral [IiS03, Bes00, Pau24a]. Spectrometry [JBE^{+}00, NFG^{+}91]. spectroscopic [Mas04]. spectrum [Gau14, GB27, Pau26c, RC25, Tho26b]. Speech [Ehr31]. speeches [Ano64, SS98]. Spektrallinien [Pau24a]. Spektren [Pau25c]. spel [vE95]. Spiegelungs [Pau58c]. Spiegelungs-Symmetrien [Pau58c]. Spin [Bur58, FP39b, Fie39, Gul54, LZ58, Mor07, Pau40a, PH45, Tch72, Tom97, Ara90, BH09, Del61, DC86, DPS97, FP39a, Gou61, Gou71, Gou76, Hel67, Pau50a, Pau53c, SG78, SP77, Uhl76, dlPA71, FR13a]. spin-boson [Ara90]. Spin-Dependent [PH45]. spin-flip [Hel67]. spin-statistics [DPS97]. Spinning [Boh26, Tho26a, UG26]. spinor [Mie81]. Spirit [Gib19, vEA02]. spiritual [Scu07c]. Springer [Eis00, Hen86, Sch01]. Stability [DL67, LD68, LT91, LT97, LT01, LT05]. staircase [Fis10, Fis12]. Standpoint [Pau94e, Pau26c, Pau28]. Standpunkt [Pau26c, Pau28]. Stark [Pau47c]. Starkeffekt [Pau47c]. starken [Pau23c]. Stars [Str04b, Str07, LT91, LT97, LT01, LT05]. State [BM04, Nel72, BCP89, CNT00, Kal89, KK73, NM91, Pau53d, ZY88]. states [Cav00, CNT99, FRKK95, Hei56, JMS80, SHS81]. stationary [EP43]. Statistical [Pau73d, PE00d, PB40, Pau56d, Pau47e]. Statistics [Bur58, GH99, LZ58, Pau40a, Bel94, CP92, Del61, DPS97, Hal91, Pau36a, Pau50a]. statistique [Pau36a]. Statistische [Pau47e]. steps [Wei70]. still [Kae48]. Stitch [OSWR86]. Stochastic [dlPA71]. Stoner [Nau13]. Störungstheorie [Pau24c]. Story [Gam85, Tom97, Bag11, Esp07, Gam66a, Gam66b, Gam72]. Stoßes [Pau24d]. Strahlenablenkung [Pau19a]. Strahlung [Pau23a, Pau29d]. Strahlungsdruck [PB27]. Strange [Mil09, vB10]. Street [FW07]. strengths [Pau47c]. Streustrahlung [Pau33h]. String [Ber06, Woi06]. Strings [MGG^{+}95]. Strong [PD42a, PD42b, PH45, KMS80, Pau23c, Pau43c]. Structure [CCJ^{+}34, GRE^{+}01, GB29, Hug89, Pau34, QS98, UG26, VNB08, Bel82, Boh21, Boh26, GU26, GB27, GB30, GP31, KP55, Pau23c, Pau25c, Pau53a, PK56, PT59, Wei70, ZBG30, CCJ^{+}34]. structures [RC25, SS75]. structuur [GU26]. strukturye [PK56]. Studies [dB86]. Study [dR99, Fis10, Fis12]. stuff [Haw11]. Sube [GRE^{+}01]. subgroup [dGPS00]. substructure [ATY92]. Suche [MM11]. Sucher [MKD^{+}91]. suitable [BG88]. sulla [Fer24]. Sun [Pla89]. Sündermann [HEB^{+}80]. Superconducting [Aoi73, Aoi71]. Superficial [ATY92]. Supersymmetric [ACD90, Łop91]. supersymmetries [NN99]. supplementary [Dys60]. Surfaces [OSWR86]. Surfaces/Paul [OSWR86]. susceptibilities [Pau33f]. susceptibility [Hel67]. svjazi [PB58]. Swedish [Kle59]. Switzerland [AvMMH^{+}07]. Symbols [BT58, DT58]. Symbols/Additions [BT58, DT58]. symmetric [Fin49]. Symmetrical [Maj37]. Symmetrien [Pau58c]. Symmetries [Pau94u, ASG05, DC86, Han10, Pau58c, PST06, vM87, GHMP87]. symmetrisches [Fin49]. symmetry [Xue91]. Symposium [Meh73, LM85]. synchronicity [GP00]. Synchronizität [JP52]. Synchrotron [HEB^{+}80]. System [Pau94o, NMS84, Pau55c]. Systeme [MGG^{+}95, BP22, BP63]. Systeme/Bunde/Havlin [MGG^{+}95]. Systems [GH99, Pau52c, BP22, BP63, Fer24, FRKK95, KM89, SOCg82, MKD^{+}91]. Systems/Elster [MKD^{+}91].

T. [KP55]. Table [Pau52c, Fin49]. Tagung [Fis99]. tails [PB27]. taking [ZSJY82]. tale [Sce16]. talk [Pau53c]. talks [BDR84]. Taub [CV02]. taught [CG14]. teaching [Boy14, BPP64]. tearing [Pau70, Pau72b, Pau90a]. Tech/Scribal [MKD^{+}91]. Technical [EGO97, BPP64]. Techniques [HEB^{+}80]. Technischen [EGO97]. Technological [MGG^{+}95]. Technology [RW94]. Teil [ARK^{+}99, BDR93, Eis00, Eis02, GP31, HW87, OSWR86, Pau33c, Pau33d, vM96, vM99, vM01, vM05b, Pau25b]. Teilchen [FP39b, Fie39, Fis10, Fis12, PB27]. tells [FS06]. temperature [McC86]. temps [Pau72b]. tensor [Han11, PST06]. Tensors [SZ97]. Tentativo [Fer34]. tenu [CCJ^{+}34]. Teoria [Fer34, Gul54, Maj37, Pau58e]. teorii [PK56, Pau75a, Pau77a]. teoriyu [Pau65b]. ter [Pau56d]. term [FWS^{+}85, GB27]. Termordnung [GB27]. terms [Pau23c]. Test [JBE^{+}00, Kij72, NFG^{+}91, RS74, Bar10, KMS80, vO69]. testing [AP80, CM09]. Teubner [OSWR86]. Texte [Sti77]. Thales [Lak96]. their [Coo03, Pau24a]. Theorem [Pau28, PF37, DPS97, Nag10]. Theoretical [Ano58, ARK^{+}99, DH91, FW60, Meh73, EAW15, Pau24a, Pau20b]. Theoretische [MGG^{+}95, Pau29c, Pau51e, Pau51d, MKD^{+}91, Pau20b]. theoretischen [Pau24a]. Theorie [FP38, Fie39, KP23, Pau19c, Pau21a, Pau35a, PF38, Pau36a, Pau53d, PS32a, PS32b, Pau36b, Pau53b, Hei84, Pau29d, Pau33c, Pau52d]. Theorien [JP52, Krö54, Pau52b]. Theories [Pau41, PH45, Pau94g, CP92, GH75, Krö54, Pau32d, Pau43c, Pau47b, Pau52b, Pau53a]. Theorist [SM08]. Theory [Ber06, Dir27, Ein22, Fie39, Gam85, GM87, GM88, GM89a, GM89b, HP93, KK73, MKD^{+}91, Pau33c, PR36, Pau38a, Pau43a, PK43a, PK43b, Pau46b, Pau49, PV49, PE57, Pau58f, Pau75b, Pau81, Pau93, Pau94c, Pau94q, Pau94s, Pau94t, PG00, QS98, Sch01, SZ97, VNB08, BNP05, Bon81, Bor47, Cli65, Cli87, Fer34, FP38, Fre28, Gam66a, Gam66b, Gam72, Hei84, KP55, KP23, Łop91, Maj37, MS94, OS00, Pau19a, Pau19c, Pau20c, Pau21a, Pau24c, Pau26b, PS32a, PS32b, Pau33f, Pau35a, Pau36b, Pau36a, Pau38b, PF38, PD42b, Pau48b, Pau53d, Pau53b, Pau55b, Pau56c, PK56, Pau65b, Pau73b, Pau73e, Pau74, PE00b, PE00e, Pei60a, SB33, SB53, Woi06, dlPA71, Pau21b, Pau29d, Pau52d, Pau58e, Pau25a, Dys60]. there [Fra04]. thermal [Pau23a]. thermische [Pau23a]. Thermodynamical [RC88]. Thermodynamics [Pau73e, PE00e, Pau58h]. Thermodynamik [Pau58g, Pau58h]. Thirty [Gam66a, Gam66b, Gam72, Gam85]. Thomson [Lak96]. Thought [Gie90, MKD^{+}91, Pau94p, vM90, Lau88, Pau56f, Pau59]. Thoughts [Fis04, Fis11]. threat [Mas04]. Three [NMS84, Wei70, ASG05, FRKK95, NM91, RL77, SOCg82]. Three-body [NMS84, SOCg82]. three-cluster [FRKK95, SOCg82]. three-electron [RL77]. three-level [NM91]. Tidal [HEB^{+}80]. Time [BBMS08, Coo03, Pau94r, Tel03, CP84, Dys02, Enz85a, Enz02, Hei60, Pau36c, Pau55a, Pau70, Pau72a, Pau72b, Pau90a, BS12, Str04a]. Time-dependent [BBMS08]. Times [MGG^{+}95, Coo03, CM09, Kle05]. today [AF14]. toeval [Cas92]. Too [Hür22]. Toomer [HW87]. topics [Pau73c, PE00c]. Topological [MGG^{+}95]. total [Ein41]. Transformation [Nel72]. transformations [Pau40b]. Transition [SZ97, Suz73]. Translated [Dys60]. Translation [Enz94a]. transmission [SS75]. Triad [Kev00]. Trieste [Meh73]. Trudy [Pau75a, Pau77a]. true [Cas03]. Tübingen [Pau30d]. Twentieth [FW60]. Twins [Sul62]. Two [FM68b, PCC10, ASG05, Bes00, FM68a, GS04, Kij72, Lin04, MS94, NM91, Pau39]. two- [NM91]. Two-Center [FM68b, FM68a]. Two-Cooper-pair [PCC10]. two-dimensional [Bes00]. two-fermion [MS94]. two-matrix [Kij72]. two-valuedness [Pau39]. Type [SZ97, PZ88, RL77].

U [GRE^{+}01, ARK^{+}99, HEB^{+}80, MGG^{+}95, MKD^{+}91, OSWR86]. u.a [Hen86, HvMW79, vMHW85, vM93, vM96, vM99, vM01, vM05b]. u.a. [Eis00]. U.I.P [BT58, DT58]. Übersichtsartikel [BDR84]. Übungsbuch [BPP64]. Ultracold [MKD^{+}91]. ultraviolet [BNP05]. Unbewussten [Pau54b]. unbounded [Kae48]. uncertainty [CG14]. Unconscious [Pau94e, Pau54b]. Uncovered [Hür22]. underdetermination [Mas04]. Unfinished [DT18]. unfragmented [Boy14]. Unify [Ber06]. Union [BT58, DT58]. unitaire [PS32a, PS32b]. unitarity [CP92]. unitary [PS32a, PS32b]. Units [BT58, DT58]. Units/Misprints [BT58, DT58]. unity [Woi06]. University [Coo03, Dys02, EGO97, Sta07, Str04a]. unknown [Esp07, NM26, PB40, vdW73]. unmechanical [UG25]. unmechanischen [UG25]. Unmechanischer [Ser77]. Unravelling [FR13b]. unspeakable [Lau97]. upper [IiS03, KLL90]. upper-half [IiS03]. Uranspaltung [GRE^{+}01]. US$60.00 [Dys02, Str04a]. Useful [CMM^{+}07]. Using [Boy14, JBE^{+}00, YS93]. usw [Sch01].

V [All56, Bal65, Gie90, Mor56a, vM90, PB58, Pau65b, Bay96]. Validation [Kra09, Sta07, Mas05]. Validity [Cav00, CNT99, CNT00, FWS^{+}85, LL79]. Valued [Ano45a]. valuedness [Pau39]. values [Pau47c]. vanishing [Ein41, Pau40b]. variable [Sch99a, Sch99b]. variables [BC09, Pau53b]. variational [BG88]. varying [SS75]. Vector [Joh33, PK43a, PK43b, PH45, BPP64]. Vector-Pair [PH45]. Vectors [Nel72]. Vektorrechnung [BPP64]. Velocity [GR31, FWS^{+}85, Hen81b, Pau25f]. Verantwortung [GRE^{+}01]. Verhaltens [UG25]. Verita [AvMMH^{+}07]. Verlag [Sch01]. Verleihung [Ehr31, Pau31a]. Verletzung [Pau58c]. verschwindender [Pau40b]. Version [GRE^{+}01]. versus [CP84, Sza97, dR99]. vertauschbarer [ET60]. Vertrieb [ARK^{+}99]. via [FF87]. Victor [HW87, JG04]. Vierpoltheorie [Pau74]. view [Enz00, Mil07]. views [Ano85, Eis91]. viii [Coo03]. Vilenkin [MGG^{+}95]. Villars [Xue91]. Violation [BBB^{+}10, PBB^{+}11, GM87, GM88, GM89a, GM89b, GH99, Pau94u, ATY92, BMT99, BH09, CP92, CP93, Pau58c, RC88]. Violations [Pla89, KLL90, YS93]. VIP [BBB^{+}10, PBB^{+}11]. virial [Nag10]. Vischer [MKD^{+}91]. Visit [DGB^{+}01]. Vklad [Pau65b]. Vleck [Pau33f]. Vol [Bro95, Eis02, OSWR86]. Volume [Eis00, HW87, HvMW79, Sta09, vMHW85, vM93, vM96, vM99, vM01, vM05b, FW60, Rei72, BPP64]. vom [Pau26c, Pau28, Pau54b, UG25, Fis99]. Vorlesungen [Pau25b, Pau29c, Pau51e, Pau51d, PE57]. Vorsstellungen [JP52]. Vorstellungen [Krö54, Pau52b]. Vorträge [BDR84, Pau61b, Pau61a]. Voß [MGG^{+}95].

W [All56, Ano53, BT58, DT58, Dys60, JP52, MKD^{+}91, Mor56a]. W. [Ein22, Ein79, Enz73, Enz00, Krö54, Pau31c]. W.W [vB10]. während [vM05a]. Wahrscheinlichkeit [Pau52a, vdWPR52, Pau54c]. Walk [Web73]. Was [Gil09, Bel86]. Wasserstoff [Pau21c]. Wasserstoff-Molekülions [Pau21c]. Wasserstoffmolekülions [Pau22]. Wasserstoffspektrum [Pau26c]. Wave [Pau73f, PE00f, Bad11, DC86, FP39a, HP29, HP30, MS94, Pau32c, Pau33g, PW34, PF38, Pau39, Pau40b, Pau53b, Pau58b]. wavelength [FP38]. waves [Pau33d]. way [De 14]. weak [Pau23c, Pau43c]. Wechselwirkung [Hei84]. Weg [Sti77]. Wegbereiter [vM68]. Weinberg [MGG^{+}95]. Weird [Str11]. Weise [GRE^{+}01]. Weiss [MGG^{+}95]. Weisskopf [Bal65, HEB^{+}80, JG04, All56, Dit15, Esp07, HW87, Mor56a]. Weizsäcker [GRE^{+}01]. well [Gou76]. Wellenfelder [HP30, HP29]. Wellengleichung [Pau32c, PW34]. Wellengleichungen [Pau40b]. Wellenmechanik [Pau33g, Pau39, Pau58b]. Weltharmonie [Fis04, Fis11]. Wenske [MGG^{+}95]. Werk [Ble91b, MGG^{+}95]. Werk/Weinberg [MGG^{+}95]. Werke [BDR84, BDR93]. werkelijkheid [Cas92]. Werner [GRE^{+}01, MGG^{+}95, BDR84, BDR93]. Werten [Pau47c]. Western [Pau94p, Pau56f, Pau59]. Weyl [Cha68, Cha68, CP84, Han11, Mie81, Pau19a, Pau19c]. Weyls [Pau19a]. Where [Fin92, Pei92]. Who [Ano45a, Cli87]. Wie [Bel86]. wijsheid [vE95]. Wilhelm [MGG^{+}95]. win [Ano45a]. Windows [GRE^{+}01, Nat03]. winner [Enz05]. winners [dVNHdVH53, dVNHdVH71]. Wirken [EGO97, BS12]. Wirklichkeit [Rös91]. Wirklichkeitsauffassungen [Eis91]. wisdom [vE95]. Wismut [BG28]. Wismuts [ZBG30]. Wismutspektrums [GB27, Tho26b]. Wissenschaft [MGG^{+}95, Pau56f, Pau59, APWB95, FS06, Pau56c]. Wissenschaft/Cassidy [MGG^{+}95]. Wissenschaftlers [GRE^{+}01]. Wissenschaftlers/Dürr [GRE^{+}01]. Wissenschaftliche [BDR84, BDR93, MKD^{+}91]. wissenschaftlichen [RP79]. Wissenschaftlicher [ARK^{+}99, Eis00, GRE^{+}01, HEB^{+}80, Hen86, HvMW79, MGG^{+}95, OSWR86, vMHW85, vM93, vM96, vM99, vM01, vM05b, vM68, Eis02]. within [ST82, ST84]. WKBJ [SP77]. Woit [Ber06]. Wolfgang [Ano45a, Ano45b, Ano58, Ano93, ARK^{+}99, Bet00, Coo03, Dys02, Ehr31, Eis91, Eis00, Eis02, Enz05, FR13c, Fis99, Fis04, Fis11, Gie90, HEB^{+}80, HW87, Kem59, MGG^{+}95, MM11, Str04a, Tel03, vB10, vM90, vW59, Ano59a, Ano59b, Ano60b, AvMMH^{+}07, Bal65, Bet00, Ble91a, Ble91b, Bro95, DGB^{+}01, Dys02, EFK90, Eis91, Eis00, Enz85a, Enz85b, EM88, Enz94a, Enz94b, EvM94, EGO97, Enz00, Enz02, FW59, Fie59, FW60, Fis99, Fis04, Fis11, Gar63a, Gie05, Hei59, Hen86, HvMW79, Kle59, KW64, Lan59b, Lan59a, Lau80, Lau82, Lau85b, Lau85a, Lau88, Lau97, Mei92, Mey04, Mil09, Pai00a, Pei60b, RP79, Sch01, Smu90, Str00, Str08, Ter65, Uns00, YK73, vE98, vExx, vEA02, vM00, vMHW85, vM93, vM96, vM99, vMS01, vM01]. Wolfgang [vM05b, vW59]. Work [Enz73, Ble91a, Ble91b, Dys02, Enz85a, Pau24d, BS12, Pet11]. Works [BDR84, Cha68, BDR93, Bor63, EGO97, Pau77b]. World [Hür22, Del72, Fis04, Fis11, Haw11]. Wörterbuch [MGG^{+}95]. writer [BS12]. Writings [EvM94]. Wrong [Ber06, Fin92, Pei92, Woi06].

X [Bro86, Che85]. xi [Bro86]. xii [Sta07]. xiv [Sta07]. XV [Sch01]. xxix [Hen86]. xxv [vB10].

Year [FR13b, Pau85a, Pau01]. Years [ARK^{+}99, Gam85, LM85, Enz82, Gam66a, Gam66b, Gam72, Gou76, Kae48, RP79, Uhl76, vM05a]. York [Dys60, Dys02, Eis00, Hen86, Str04a, vB10].

Zahl [MKD^{+}91, MM11]. Zeeman [BG28, BU26, Pau23b, Pau25f]. Zeemaneffekt [BG28, Pau25f]. Zeemaneffektes [Pau23b]. Zeichnen [GRE^{+}01].

Zeilinger [Fis10, Fis12]. Zeit [Hei60, Pau36c, Pau70, Pau90a, BS12]. Zero [ET60]. Zero-point [ET60]. Zufall [Fis99]. zukunftige [Hei79b, Hei79a]. Zum [Pau77b, Fis99, Fis04, Fis10, Fis11, Fis12]. Zuordnung [Pau23c]. Zur [Pau24d, Ehr31, GRE^{+}01, Pau35a, Pau47c, Pau48b, Pau55b, Pau01, Fer26, FP38, GP31, HP29, HP30, JP28, KP23, Pau19c, Pau21a, Pau23c, Pau24a, Pau27b, PF38, Pau57c, Pau58g, Pau58h, Pau61d, ZBG30]. Zürich [EGO97, EGO97]. Zusammenarbeit [vM05a]. Zusammenfassung [Pau56c]. Zusammenhang [Pau25c]. Zusammenhänge [JP52, JP52]. Zwang [UG25, Ser77]. zweiatomiger [Pau21a]. Zweideutigkeit [Set09]. Zweiwertigkeit [Pau39]. zwischen [Fis99, Fis04, Fis11, Pau23a, Pau24d].

References

Amado:1990:SQM

[ACD90] R. D. Amado, F. Cannata, and J. P. Dedonder. Supersymmetric quantum mechanics, the Pauli Principle, and nucleonalpha scattering. Physical Review C (Nuclear Physics), 41(3): 1289–1291, March 1, 1990. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490x, 1538-4497. URL http://link.aps.

org/doi/10.1103/PhysRevC.41.1289.

Adams:1972:FGN

[Ada72] J. B. Adams, F.R.S. Four generations of nuclear physicists. Notes and Records of the Royal Society of London, 27(1):75– 94, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Atmanspacher:2014:PJC

[AF14] Harald Atmanspacher and Christopher A. Fuchs, editors. The Pauli–Jung conjecture and its impact today. Imprint Academic, Exeter, England, second edition, 2014. ISBN 1-84540-668-0, 1-84540-759-8 (e-book). 333 (est.) pp. LCCN QC16.P37 A384

2014. URL http://site.ebrary.com/lib/alltitles/Doc?
id=10890219.

Allison:1956:BBB

[All56] Samuel K. Allison. Books: Niels Bohr and the Development of Physics, edited by W. Pauli with the assistance of L. Rosenfeld and V. Weisskopf. Bulletin of the Atomic Scientists, 12(2): 61, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1938:PAS

[Ano38] Anonymous. Professor Arnold Sommerfeld. Journal of Applied Physics, 9(12):754–755, December 1938. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. Note for the 70th birthday of Arnold Sommerfeld.

Anonymous:1945:EAW

[Ano45a] Anonymous. Experts on atoms win Nobel Prizes: [Wolfgang] Pauli of Princeton Institute and [Otto] Hahn, who is believed to be here, are honored. Chemistry award to Finn. Gabriela Mistral of Chile gets Literature Prize, valued at about $28,950. New York Times, ??(??):19, November 16, 1945. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/107323828/pageviewPDF/13F38F649E969AA05FF.

Anonymous:1945:NPP

[Ano45b] Anonymous. Nobel Prize for Physics for 1945: Prof. Wolfgang Pauli. Nature, 156(3970):657–658, December 1, 1945. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v156/n3970/
pdf/156657c0.pdf.

Anonymous:1953:PWP

[Ano53] Anonymous. Prof. W. Pauli, For.Mem.R.S. Nature, 171(4361): 956, May 30, 1953. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v171/n4361/pdf/171956a0.pdf.

Anonymous:1956:SRN

[Ano56] Anonymous. Science: The real neutrino. Time, ??(??):??, July 2, 1956. CODEN TYMEA9. ISSN 0040-781x (print),

2169-1665 (electronic). URL http://www.time.com/time/
magazine/article/0,9171,891295,00.html. From the story:

“[Neutrinos] were reasoned into existence by Nobel Prizewinners Enrico Fermi and Wolfgang Pauli to fill a theoretical need, and the gnawing suspicion has long persisted that they do not exist. Last week from the Atomic Energy Commission came big news. Neutrinos do exist”.

Anonymous:1958:PWP

[Ano58] Anonymous. Prof. Wolfgang Pauli: a great theoretical physicist. London Times, ??(??):??, December 1958. URL http://www-

groups.dcs.st-and.ac.uk/~history/Obits/Pauli.html.

Anonymous:1959:WP

[Ano59a] Anonymous. Wolfgang Pauli: 25.4.1900–15.12.1958. Zeitschrift für Physik, 154(2):i, April 1959. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01338369;   http://www.
springerlink.com/content/r288314972u4h726/fulltext.
pdf.

Anonymous:1959:WPO

[Ano59b] Anonymous. Wolfgang Pauli [obituary]. Physics Today, 12 (2):60, February 1959. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v12/i2/p60/s1.

Anonymous:1960:PMR

[Ano60a] Anonymous. Pauli Memorial Room at CERN. Nature, 186 (4730):1017, June 25, 1960. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v186/n4730/pdf/1861017a0.pdf.

Anonymous:1960:WP

[Ano60b] Anonymous. Wolfgang Pauli. Zeitschrift für Physik, 158(5):i, February 1960. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic). URL http://link.springer.com/
article/10.1007/BF01366502;   http://www.springerlink.
com/content/m7r253k62m512k56/fulltext.pdf.

Anonymous:1964:NLI

[Ano64] Anonymous, editor. Nobel lectures, including presentation speeches and laureates’ biographies. Physics, 1942–1962. Elsevier, Amsterdam, The Netherlands, 1964. xiii + 619 pp. LCCN ????

Anonymous:1985:KVP

[Ano85] Anonymous. Kastler’s views on Pauli’s contribution. Nature, 316(6026):308, July 25, 1985. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v316/n6026/pdf/316308a0.pdf.

Anonymous:1993:OMP

[Ano93] Anonymous. Österreichische Mathematik und Physik: Wolfgang Gröbner, Richard von Mises, Wolfgang Pauli. (German) [Austrian mathematics and physics: …]. Die Zentralbibliothek, Wien, Austria, 1993. ISBN 3-900490-03-1. 71 pp. LCCN QA27.A9 O87 1993.

Anonymous:2005:CLB

[Ano05] Anonymous. Critique de livre: Il était sept Fois La révolution: Albert Einstein et les autres. La Recherche, ??(386):11–??, May 1, 2005. CODEN RCCHBV. ISSN 0029-5671 (print), 1625-9955 (electronic). URL http:/

/www.larecherche.fr/idees/livres/il-etait-sept-fois-
revolution-albert-einstein-autres-01-05-2005-73924.

Aoi:1971:EPP

[Aoi71] Koya Aoi. Effects of Pauli paramagnetism on superconducting fluctuations. Zeitschrift für Physik, 246(1):71–83, February 1971. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831

(electronic). URL http://link.springer.com/article/10.
1007/BF01402652;   http://www.springerlink.com/content/
q39551m102544335/fulltext.pdf.

Aoi:1973:EPP

[Aoi73] Koya Aoi. Effects of Pauli paramagnetism on superconducting fluctuations. Zeitschrift für Physik, 258(3):284, June 1973. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01392839;   http://www.springerlink.com/content/w462661660544273/
fulltext.pdf.

Amado:1980:CTP

[AP80] R. D. Amado and H. Primakoff. Comments on testing the Pauli Principle. Physical Review C (Nuclear Physics), 22(3):1338– 1340, September 1, 1980. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490x, 1538-4497. URL http://link.aps.org/

doi/10.1103/PhysRevC.22.1338.

Atmanspacher:2006:PIM

[AP06] Harald Atmanspacher and Hans Primas. Pauli’s ideas on mind and matter in the context of contemporary science. Journal of Consciousness Studies, 13(3):5–50, ???? 2006.

ISSN 1355-8250. URL http://www.igpp.de/english/tda/
pdf/paulijcs8.pdf.

Atmanspacher:1995:PJD

[APWB95] Harald Atmanspacher, Hans Primas, and E. (Eva) Wertenschlag-Birkhäuser, editors. Der Pauli–Jung-Dialog und seine Bedeutung für die moderne Wissenschaft. (German) [The Pauli–Jung Dialog and its Meaning for Modern Science]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-58518-4. viii + 365 pp. LCCN Q175.3 .P38

1995. URL http://www.zentralblattmath.org/zmath/en/
search/?an=0820.01012.

Arai:1990:AAA

[Ara90] Asao Arai. An asymptotic analysis and its application to the nonrelativistic limit of the Pauli–Fierz and a spin-boson model. Journal of Mathematical Physics, 31(11):2653–2663, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v31/i11/p2653_s1.

Atmanspacher:1999:BSM

[ARK^{+}99] Harald Atmanspacher, Helmut Rechenberg, Horst Kant, Rolf Hempelmann, Reinhold Blümel, Gerhard Börner, Houjun Mo, Philippe Blanchard, Marius Grundmann, Dietrich Stauffer, and Christian Uebing. Bücher und Software: Meÿenn: Wolfgang Pauli. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a., Band IV/Teil II: 1953–1954/Mladjenovic: The Defining Years in Nuclear Physics. 1932–1960s/Frenkel: Professor Friedrich Houtermans Arbeiten, Leben, Schicksal/Kärger and Heitjans: Diffusion in Condensed Matter/Friedrich: Theoretical Atomic Physics/Peacock: Cosmological Physics/Grosse u Martin: Particle Physics and the Schrödinger Equation/ Gaponenko: Optical Properties of Semiconductor Nanocrystals/Gershenfeld: The Nature of Mathematical Modeling/ Origin 6.0: Vertrieb. Physikalische Blätter, 55(9):79–83, September 1999. CODEN PHBLAG. ISSN 0031-9279 (print),

1521-3722 (electronic). URL http://onlinelibrary.wiley.
com/doi/10.1002/phbl.19990550919/abstract.

Adam:2005:SDP

[ASG05] C. Adam and J. Sánchez-Guillén. The symmetries of the Dirac– Pauli equation in two and three dimensions. Journal of Mathematical Physics, 46(5):052304, May 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/
i5/p052304_s1.

Akama:1992:SVP

[ATY92] Keiichi Akama, Hidezumi Terazawa, and Masaki Yasuè. Superficial violation of the Pauli Principle due to the possible substructure of electrons. Physical Review Letters, 68(12):1826– 1829, March 23, 1992. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.

aps.org/doi/10.1103/PhysRevLett.68.1826.

Atmanspacher:2007:WPP

[AvMMH^{+}07] Harald Atmanspacher, Karl von Meyenn, Ulrich Müller-Herold, Reinhard Nesper, and Hans Primas, editors. Wolfgang Pauli’s Philosophical Ideas and Contemporary Science: May 20–25, 2007, Monte Verita, Ascona (Switzerland). ????, ????, 2007.

ISBN ???? LCCN ???? URL http://www.solid.ethz.ch/
pauli-conference/thematic.htm.

Badnell:2011:BPD

[Bad11] N. R. Badnell. A Breit–Pauli distorted wave implementation for autostructure. Computer Physics Communications, 182(7):1528–1535, July 2011. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465511001160.

Baggott:2011:QSH

[Bag11] J. E. Baggott. The quantum story: a history in 40 moments. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-956684-4 (hardcover). xix + 469 + 16 pp. LCCN QC173.98 .B34 2011.

Balazs:1965:BRW

[Bal65] Nandor L. Balazs. Book review: Wolfgang Pauli, R. Kronig and V. F. Weisskopf, Collected Scientific Papers. Physics Today, 18(6):51–52, June 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v18/i6/p51/s2.

Barker:1979:PR

[Bar79] William A. Barker. Pauli remembered. Physics Today, 32 (2):11, February 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v32/i2/p11/s1.

Barabash:2010:ETP

[Bar10] A. S. Barabash. Experimental test of the Pauli Exclusion Principle. Foundations of Physics, 40(7):703–718, July 2010. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/s10701-
009-9336-3.

Baylis:1996:PAC

[Bay96] W. E. Baylis. Pauli-algebra calculations in MAPLE V. In Clifford algebras with numeric and symbolic computations, pages 69–82. Birkhäuser Boston Inc., Cambridge, MA, USA, 1996.

Bartalucci:2010:VEL

[BBB^{+}10] S. Bartalucci, S. Bertolucci, M. Bragadireanu, M. Cargnelli, C. Curceanu (Petrascu), S. Di Matteo, J.-P. Egger, C. Guaraldo, M. Iliescu, T. Ishiwatari, M. Laubenstein, J. Marton, E. Milotti, D. Pietreanu, T. Ponta, A. Romero Vidal, D. L. Sirghi, F. Sirghi, L. Sperandio, O. Vazquez Doce, E. Widmann, and J. Zmeskal. The VIP experimental limit on the Pauli Exclusion Principle violation by electrons. Foundations of Physics, 40(7):765–775, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/s10701-009-9346-1.

Bouguerra:2008:TDP

[BBMS08] Y. Bouguerra, A. Bounames, M. Maamache, and Y. Saadi. Time-dependent Pauli equation in the presence of the Aharonov– Bohm effect. Journal of Mathematical Physics, 49(4):042107, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i4/p042107_s1.

Born:1969:AP

[BBSR69] Max Born, R. J. (Roger John) Blin-Stoyle, and J. M. Radcliffe. Atomic Physics. Dover books on physics and chemistry. Dover, New York, NY, USA, eighth edition, 1969. ISBN 0-486-65984-4. xiv + 495 + 11 pp. LCCN QC776 .B5713 1989.

US$11.95. URL http://www.loc.gov/catdir/description/
dover031/89012033.html.

Brode:1951:RMA

[BC51] Robert B. Brode and E. U. Condon. R. A. Millikan and Arnold Sommerfeld elected honorary members. Journal of the Optical Society of America, 41(2):63, February 1, 1951. CODEN JOSAAH. ISSN 0030-3941. URL https://

www.osapublishing.org/josa/abstract.cfm?uri=josa-41-
2-63&origin=search.

Bacciagaluppi:2009:HSP

[BC09] Guido Bacciagaluppi and Elise Crull. Heisenberg (and Schrödinger, and Pauli) on hidden variables. Studies in History and Philosophy of Modern Physics, 40(4):374–382, December 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000501.

Baguer:1989:EFP

[BCP89]  N. Baguer, R. Capote, and R. Pedrosa.  Exact for-

mulation of particle-hole state densities in the equidistant spacing model with Pauli and pairing corrections. Zeitschrift für Physik, 334(4):397–402, December 1989. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01294745;
 http://www.springerlink.com/content/j847p858gr076774/
fulltext.pdf.

Babusci:2011:NED

[BDQR11] D. Babusci, G. Dattoli, M. Quattromini, and P. E. Ricci. A note on the extension of the Dirac method. Applied Mathematics and Computation, 218(4):1495–1497, October 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300311008010.

Blum:1984:WHGb

[BDR84] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke. Abt. B: Wissenschaftliche Übersichtsartikel, Vorträge und Bücher. (German) [Werner Heisenberg: Collected Works: Series B: Scientific review papers, talks, and books]. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 3-540-13020-9, 0-387-13020-9. x + 937 pp. LCCN QC3 .H33 1984.

Blum:1993:WHG

[BDR93] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke. Abteilung A. Teil III. Wissenschaftliche Originalarbeiten. (German) [Werner Heisenberg: Collected works. Series A. Part III. Original scientific papers]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-13848-X. x + 700 pp. Annotated by Rolf Hagedorn, Heinz Koppe, H.-P. Dürr and H. Rechenberg.

Belloni:1982:PNH

[Bel82] Lanfranco Belloni. Pauli’s 1924 note on hyperfine structure. American Journal of Physics, 50(5):461–464, May 1982. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Beller:1986:BRW

[Bel86] Mara Beller. Book review: Wie es eigentlich gewesen? [How was it actually?] The Creation of Quantum Mechanics and the Bohr–Pauli Dialogue. Isis, 77(1):107–109, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232508.

Belloni:1994:FRF

[Bel94] Lanfranco Belloni. On Fermi’s route to Fermi–Dirac statistics. European Journal of Physics, 15(3):102–109, May 1994. CO-DEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic).

URL http://iopscience.iop.org/0143-0807/15/3/002.

Beller:1999:QDM

[Bel99] Mara Beller. Quantum dialogue: the making of a revolution. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1999. ISBN 0-226-04181-6 (hardcover), 0-226-04182-4. xv + 365 + 8 pp. LCCN QC174.13 .B45 1999. URL http://www.loc.

gov/catdir/description/uchi052/99035499.htm;   http:
//www.loc.gov/catdir/toc/uchi052/99035499.htm.

Bergman:2006:BRE

[Ber06] Aaron Bergman. Book review: Not Even Wrong: The Failure of String Theory and the Continuing Challenge to Unify the Laws of Physics, Peter Woit, Jonathan Cape, London 2006. World Wide Web document, August 18, 2006. URL http://zippy.

ph.utexas.edu/~abergman/Review.pdf.

Besch:2000:ESG

[Bes00] Alexander Besch. Eigenvalues in spectral gaps of the twodimensional Pauli operator. Journal of Mathematical Physics, 41(12):7918–7931, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bethe:1972:PPP

[Bet72] Hans A. Bethe. Pauli Principle and pion scattering. Report LA-5052, US Atomic Energy Commission, Washington, DC, USA,

October 1972. URL http://www.osti.gov/accomplishments/
documents/fullText/ACC0103.pdf.

Bethe:2000:RWP

[Bet00] Hans A. Bethe. Remembering Wolfgang Pauli / Erinnerungen an Wolfgang Pauli. In Unser [Uns00], pages 85–90. ISBN 3-9521386-2-2. LCCN QC16.P37 W635 2000.

Back:1928:KZW

[BG28] E. Back and S. Goudsmit. Kernmoment und Zeemaneffekt von Wismut. (German) [Nuclear moment and Zeeman effect in bismuth]. Zeitschrift für Physik, 47(3–4):174–183, February 1928. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02055794;   http://www.springerlink.com/content/rnwt7303k7677336/

.

Baretty:1988:MBP

[BG88] Reinaldo Baretty and Carmelo Garcia. Modified Breit–Pauli Hamiltonian suitable for variational calculations. International Journal of Quantum Chemistry, 34(S22):425–430, March 12– 19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Badrinarayanan:1999:PPC

[BGTJ99] R. Badrinarayanan, A. Góngora-T, and Jorge V. José. Pauli principle and chaos in a magnetized disk. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 60(2):1590–1599, August 1, 1999. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://

link.aps.org/doi/10.1103/PhysRevE.60.1590.

Brown:1982:BEP

[BH82] Laurie M. Brown and Lillian Hoddeson. The birth of elementary-particle physics. Physics Today, 35(4):36–43, April 1982. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/

1/phtoad/v35/i4/p36_s1. This is a much-abridged version of [BH83a].

Brown:1983:BEP

[BH83a] Laurie M. Brown and Lillian Hoddeson. The birth of elementary particle physics. In The Birth of Particle Physics [BH83b], pages 3–36. ISBN 0-521-24005-0 (hardcover), 0-521-33837-9 (paperback). LCCN QC793 .B57 1983.

Brown:1983:BPP

[BH83b] Laurie M. Brown and Lillian Hoddeson, editors. The Birth of Particle Physics. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-24005-0 (hardcover), 0-521-33837-9 (paperback). LCCN QC793 .B57 1983.

Bohata:2009:MVB

[BH09] Martin Bohata and Jan Hamhalter. Maximal violation of Bell’s inequalities and Pauli spin matrices. Journal of Mathematical Physics, 50(8):082101, August 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/
i8/p082101_s1.

Bleuler:1991:WPH

[Ble91a] K. Bleuler. Wolfgang Pauli: his scientific work and his ideas on the foundations of physics. In Debrus and Hirshfeld [DH91], pages 304–310. ISBN 3-540-53570-5 (Berlin), 0-387-53570-5 (New York). LCCN QC20.7.G44 G47 1991.

Bleuler:1991:WPS

[Ble91b] K. Bleuler. Wolfgang Pauli: über sein Werk und seine Ideen zu den Grundlagen der Physik. (German) [Wolfgang Pauli: his scientific work and his ideas on the foundations of physics]. In Debrus and Hirshfeld [DH91], pages 298–303. ISBN 3-540-53570-5 (Berlin), 0-387-53570-5 (New York). LCCN QC20.7.G44 G47 1991.

Barbaroux:2005:QEE

[BLV05] Jean-Marie Barbaroux, Helmut Linde, and Semjon Vugalter. Quantitative estimates on the enhanced binding for the Pauli– Fierz operator. Journal of Mathematical Physics, 46(12):122103, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i12/p122103_s1.

Bethe:2004:CAS

[BM04] Hans A. Bethe and N. David Mermin. A conversation about solid-state physics. Physics Today, 57(6):53–56, June 2004. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://ifup.cit.cornell.edu/bethe/;
 http://link.aip.org/link/phtoad/v57/i6/p53/s1. Also

available as a videorecording, made 25 February 2003, archived with the Cornell University Library.

Baron:1999:LPP

[BMT99] E. Baron, Rabindra N. Mohapatra, and Vigdor L. Teplitz. Limits on Pauli Principle violation by nucleons. Physical Review D (Particles and Fields), 59(3):036003:1–036003:4, January 8, 1999. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.

org/doi/10.1103/PhysRevD.59.036003.

Bertini:2005:RDR

[BNP05] Massimo Bertini, Diego Noja, and Andrea Posilicano. Rigorous dynamics and radiation theory for a Pauli–Fierz model in the ultraviolet limit. Journal of Mathematical Physics, 46(10): 102305, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i10/p102305_s1.

Bohr:1921:ASb

[Boh21] Niels Bohr. Atomic structure. Nature, 108(2711):208–209, October 13, 1921. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v108/n2711/pdf/108208b0.pdf.

Bohr:1926:SES

[Boh26] Niels Bohr. Spinning electrons and the structure of spectra. Nature, 117(2938):265, February 20, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v117/n2938/
pdf/117265a0.pdf.

Bohr:1960:F

[Boh60] Niels Bohr. Foreword. In Markus Fierz and Victor F. (Victor Frederick) Weisskopf, editors, Theoretical physics in the twentieth century: a memorial volume to Wolfgang Pauli, pages 1–4. Interscience Publishers, New York, NY, USA, 1960. LCCN QC3F54T 1960.

Bonchev:1981:ITI

[Bon81] D. Bonchev. Information theory interpretation of the Pauli principle and Hund rule. International Journal of Quantum Chemistry, 19(4):673–679, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bonolis:2005:BPS

[Bon05] Luisa Bonolis. Bruno Pontecorvo: From slow neutrons to oscillating neutrinos. American Journal of Physics, 73(6):487– 499, 2005. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/73/6/10.1119/1.1852540.

Born:1947:BRP

[Bor47] Max Born. Book review: Pauli: Meson theory of nuclear forces. Nature, 160(??):418–??, ???? 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1963:AAGb

[Bor63] Max Born. Ausgewählte Abhandlungen. (German) [Selected works], volume 2. Vandenhoeck & Ruprecht, Göttingen, West Germany, 1963. viii + 706 pp. LCCN ???? Mit einem Verzeichnis der wissenschaftlichen Schriften (German) [With a table of contents of scientific writings].

Boyce:2014:ULD

[Boy14] Conal Boyce. Using logic to define the Aufbau–Hund–Pauli relation: a guide to teaching orbitals as a single, natural, unfragmented rule-set. Foundations of Chemistry, 16(2):93–106, July 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/article/10.
1007/s10698-012-9176-7;   http://link.springer.com/
content/pdf/10.1007/s10698-012-9176-7.pdf.

Born:1922:QGM

[BP22] Max Born and W. Pauli, Jr. Über die Quantelung gestörter mechanischer Systeme. (German) [On the quantization of constrained mechanical systems]. Zeitschrift für Physik, 10(1): 137–158, December 1922. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.

com/article/10.1007/BF01332555;   http://www.springerlink.
com/content/l3686r1744052961/fulltext.pdf.

Born:1963:QGM

[BP63] Max Born and W. Pauli, Jr. Über die Quantelung gestörter mechanischer Systeme. (German) [On the quantization of constrained mechanical systems]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63], chapter 47, pages 1–22. LCCN ???? Reprinted from Z. Physik 10, 137–158 (1922).

Borg:2003:PAS

[BP03] J. L. Borg and J. V. Pulé. Pauli approximations to the selfadjoint extensions of the Aharonov–Bohm Hamiltonian. Journal of Mathematical Physics, 44(10):4385–4410, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bewert:1964:LUM

[BPP64] Fritz Bewert, Heinz Pester, and Wolfgang Pauli. Lehr- und Übungsbuch der Mathematik für Ingenieur- und Fachschulen. Band 3: Analytische Geometrie, Vektorrechnung und Infinitesimalrechnung. (German) [Teaching and practice of mathematics book for engineering and technical schools. Volume 3: analytical geometry, vector algebra and calculus.]. Harri Deutsch, Frankfurt/Main, Germany and Zürich, Switzerland, 1964. 651 pp. LCCN ????

Brandt:2009:HCD

[Bra09] Siegmund Brandt. The harvest of a century: discoveries of modern physics in 100 episodes. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-954469-7 (hardcover). xiv + 500 pp. LCCN QC7 .B64 2009.

Brown:1978:IN

[Bro78] Laurie M. Brown. The idea of the neutrino. Physics Today, 31(9):23–28, September 1978. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

nobelprize.org/nobel_prizes/physics/laureates/1935/;
 http://www.physicstoday.org/resource/1/phtoad/v31/

i9/p23_s1. Page 27 contains an English translation of Pauli’s famous letter proposing the existence of an as-yet-undiscovered particle, which he called the ‘neutron’. Fermi later renamed it ‘neutrino’ because of its small (and possibly zero) mass. The real neutron was first discovered by James Chadwick in 1932, a result for which he received the 1935 Nobel Prize in Physics.

Brown:1986:HJC

[Bro86] Harvey R. Brown. Book review: Hendry, John [1984]: The Creation of Quantum Mechanics and the Bohr–Pauli Dialogue. D. Reidel Publishing Company. xi + 177 pp. (ISBN 9O-277-1648-X). British Journal for the Philosophy of Science, 37(4): 497–506, December 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Brown:1995:BRK

[Bro95] Laurie M. Brown. Book review: Karl von Meyenn, Wolfgang Pauli: Scientific Correspondence with Bohr, Einstein, Heisenberg a. o.; Vol. III: 1940–1949. Physics Today, 48(4):86, April 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v48/
i4/p86/s1.

Blin-Stoyle:1973:PP

[BS73] R. J. Blin-Stoyle. Pauli’s physics. Nature, 245(5426):432, October 26, 1973. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v245/n5426/pdf/245432b0.pdf.

Pauli:2012:BRZ

[BS12] Susanne Blumesberger and Ernst Seibert, editors. “Eine Brücke über den Riss der Zeit … ”: das Leben und Wirken der Journalistin und Schriftstellerin Hertha Pauli (1906–1973). (German) [“A bridge over the crack of time … ”: the life and work of journalist and writer Hertha Pauli (1906–1973)], volume 10 of BiografiA. Neue Ergebnisse der Frauenbiografieforschung. Praesens Verlag, Wien, Austria, 2012. ISBN 3-7069-0462-4 (paperback). LCCN ????

Barz:1982:EFA

[BSI^{+}82] H. W. Barz, H. Schulz, H. Iwe, L. Münchow, G. Röpke, M. Schmidt, T. S. Biró, and J. Zimanyi. Expanding fireballs and the account of the Pauli quenching for the deuteron production. Zeitschrift für Physik, 308(2):187–188, June 1982. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01413012;
 http://www.springerlink.com/content/k62074q254lr59g3/
fulltext.pdf.

Bas:1955:BGB

[BSRD^{+}55] E. B. Bas, T. Stutz, E. Roth-Desmeules, E. Stiefel, W. Pauli, et al. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 6(1):76–80, January 1955. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL

http://link.springer.com/article/10.1007/BF01600738;
 http://www.springerlink.com/content/k2x4416g24577333/
fulltext.pdf.

Boer:1958:RIU

[BT58] J. De Boer and F. Trendelenburg. Recommendations of the International Union of Pure and Applied Physics Commission for Symbols, Units and Nomenclature/Additions to Ch. I: General Principles and Recommendations/Additions to Ch. II: Symbols for Physical Quantities/Addition to Ch. III: Mathematical Operations and Symbols/Additions to Ch. IV: Units/Misprints to Be Corrected in U.I.P. 6 (1955)/Max-Planck-Medaille an W. Pauli. Physikalische Blätter, 14(6):259–262, 1958. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Bichowsky:1926:PER

[BU26] F. R. Bichowsky and H. C. Urey. A possible explanation of the relativity doublets and anomalous Zeeman effect by means of a magnetic electron. Proceedings of the National Academy of Sciences of the United States of America, 12(2):80–85, February 15, 1926. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490

(electronic). URL http://www.jstor.org/stable/85075;
http://www.pnas.org/content/12/2/80.full.pdf+html.

Burgoyne:1958:CBS

[Bur58] N. Burgoyne. On the connection between spin and statistics. Il Nuovo Cimento (10), 8(4):607–609, May 16, 1958. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL

http://link.springer.com/article/10.1007/BF02828775.

Casimir:1983:HRH

[Cas83] H. B. G. (Hendrik Brugt Gerhard) Casimir. Haphazard reality: half a century of science. The Alfred P. Sloan Foundation series. Harper & Row, New York, NY, USA, 1983. ISBN 0-06-015028-9, 0-06-337031-X. xii + 356 pp. LCCN QC16.C36 A33 1983.

Casimir:1992:TVW

[Cas92] H. B. G. (Hendrik Brugt Gerhard) Casimir. Het toeval van de werkelijkheid: Een halve eeuw natuurkunde. (Dutch) [The coincidence of reality: a half-century physics]. Meulenhof, Amsterdam, The Netherlands, fifth edition, 1992. ISBN 90-290-9709-4. 414 pp. LCCN ????

Casti:2003:OTP

[Cas03] J. L. Casti. The one true platonic heaven: a scientific fiction on the limits of knowledge. Joseph Henry Press, Washington, DC, USA, 2003. ISBN 0-309-08547-0 (hardcover), 0-309-09510-7 (paperback). xviii + 160 pp. LCCN Q175 .C4339 2003. URL

http://www.loc.gov/catdir/toc/fy037/2003002279.html;
 http://www.nap.edu/catalog.php?record_id=10533.

Cavalcanti:2000:CVF

[Cav00] R. M. Cavalcanti. Comment on “Validity of Feynman’s prescription of disregarding the Pauli Principle in intermediate states”. Physical Review A (Atomic, Molecular, and Optical Physics), 62 (1):016101:1, June 8, 2000. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://link.

aps.org/doi/10.1103/PhysRevA.62.016101. See [CNT99]

and reply [CNT00].

Calvo:1983:CSP

[CC83] Miguel Calvo and Simón Codriansky. A class of solvable Pauli– Schrödinger Hamiltonians. Journal of Mathematical Physics, 24 (3):553–559, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i3/p553_s1.

Cockcroft:1934:SPN

[CCJ^{+}34] J. Cockcroft, J. Chadwick, F. Joliot, J. Joliot, N. Bohr, G. Gamov, P. A. M. Dirac, and W. Heisenberg, editors. Structure et propriétés des noyaux atomiques. Rapports et discussions du septième conseil de physique tenu à Bruxelles du 22 au 29 octobre 1933 sous les auspices de l’institut international de physique Solvay. (French) [Structure and properties of atomic nuclei. Reports and discussions of the Seventh Meeting on Physics held in Brussels from 22 to 29 October 1933 under the auspices of the Solvay International Institute of Physics]. Gauthier-Villars et cie, Paris, France, 1934. LCCN ???? Publiés par la commission administrative de l’institut.

Codriansky:1995:CSP

[CCS95] S. Codriansky, P. Cordero, and S. Salamó. On a class of solvable Pauli–Schrödinger Hamiltonians. Zeitschrift für Physik, 353(3):341–343, September 1995. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01292340;
 http://www.springerlink.com/content/m210574m70451675/
fulltext.pdf.

Crease:2014:QMH

[CG14] Robert P. Crease and Alfred S. Goldhaber. The quantum moment: how Planck, Bohr, Einstein, and Heisenberg taught us to love uncertainty. W. W. Norton & Co., New York, NY, USA, 2014. ISBN 0-393-06792-0 (hardcover). vii + 332 pp. LCCN QC174.123 .C74 2014.

Chadwick:1932:EN

[Cha32a] James Chadwick. The existence of a neutron. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 136(830):692–708, June 1, 1932. ISSN 0950-1207 (print), 2053-9150 (electronic). URL

http://www.jstor.org/stable/95816. This paper reports the

discovery of the neutron, for which Chadwick received the Nobel Prize in Physics in 1935.

Chadwick:1932:PEN

[Cha32b] James Chadwick. Possible existence of a neutron. Nature, 129 (3252):312, February 27, 1932. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v129/n3252/pdf/129312a0.pdf.

Chandrasekharan:1968:HWG

[Cha68] K. Chandrasekharan, editor. Hermann Weyl: Gesammelte Abhandlungen. (German) [Hermann Weyl: Collected Works]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. ???? pp.

Chevalley:1985:BRJ

[Che85] Catherine Chevalley. Book review: John Hendry, The Creation of Quantum Mechanics and the Bohr–Pauli Dialogue. Dordrecht/Boston/Lancaster: D. Reidel Publishing Company, 1984. Pp. 177. ISBN 90-277-1648-X. DF1 95. £23.50. British Journal for the History of Science, 18(3):362–364, November 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4026401.

Cline:1965:QPQ

[Cli65] Barbara Lovett Cline. The questioners: physicists and the quantum theory. Crowell, New York, 1965. vii + 274 pp. LCCN QC15

.C4. URL http://catalog.hathitrust.org/api/volumes/
oclc/372589.html.

Cline:1987:MWM

[Cli87] Barbara Lovett Cline. Men who made a new physics: physicists and the quantum theory. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-11027-3 (paperback). xii + 274 pp. LCCN QC15 .C4 1987. US$11.95. URL http://www.loc.

gov/catdir/description/uchi051/87010786.htm;   http:
//www.loc.gov/catdir/enhancements/fy0608/87010786-t.
htm;   http://www.loc.gov/catdir/enhancements/fy0609/
87010786-b.htm.

Condon:1930:IPE

[CM30] E. U. Condon and J. E. Mack. An interpretation of Pauli’s Exclusion Principle. Physical Review (2), 35(6):579–582, March 15, 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.35.579.

Curceanu:2009:PPF

[CM09] Catalina Curceanu and Edoardo Milotti. The Pauli principle faces testing times. CERN Courier: International Journal of High-Energy Physics, 49(1):25–26, January/February 2009. CO-DEN CECOA2. ISSN 0304-288X (print), 2077-9550 (electronic).

URL https://cds.cern.ch/record/1734323.

Cavalcanti:2007:UEP

[CMM^{+}07] D. Cavalcanti, L. M. Malard, F. M. Matinaga, M. O. Terra Cunha, and M. França Santos. Useful entanglement from the Pauli Principle. Physical Review B: Condensed Matter and Materials Physics, 76(11):113304:1–113304:4, September 28, 2007. CODEN PRBMDO. ISSN 1098-0121. URL http://link.aps.

org/doi/10.1103/PhysRevB.76.113304.

Coutinho:1999:VFP

[CNT99] F. A. B. Coutinho, Y. Nogami, and Lauro Tomio. Validity of Feynman’s prescription of disregarding the Pauli Principle in intermediate states. Physical Review A (Atomic, Molecular, and Optical Physics), 59(4):2624–2630, April 1, 1999. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-

4519. URL http://link.aps.org/doi/10.1103/PhysRevA.

59.2624. See comments [Cav00, CNT00].

Coutinho:2000:RCV

[CNT00] F. A. B. Coutinho, Y. Nogami, and Lauro Tomio. Reply to “Comment on ‘Validity of Feynman’s prescription of disregarding the Pauli Principle in intermediate states’ ”. Physical Review A (Atomic, Molecular, and Optical Physics), 62(1):016102:1– 016102:2, June 8, 2000. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://link.

aps.org/doi/10.1103/PhysRevA.62.016102. See [CNT99,

Cav00].

Cook:2003:ERT

[Coo03] Alan Cook. Essay review: The times of their lives, Pauli and company [No Time to Be Brief. A Scientific Biography of Wolfgang Pauli. By C. P. Enz. (Oxford University Press, 2002.). Pp. viii + 573. £35.00 (hardback). ISBN 0-19-856479-1. Scope: biography. Level: non-specialist]. Contemporary Physics, 44(4): 361–364, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cowling:1984:BVT

[CP84] Michael G. Cowling and John F. Price. Bandwidth versus time concentration: the Heisenberg–Pauli–Weyl inequality. SIAM Journal on Mathematical Analysis, 15(1):151–165, January 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cougo-Pinto:1992:CBC

[CP92] M. V. Cougo-Pinto. Connection between charge conjugation, unitarity, and statistics in theories of possible small violation of the Pauli Principle. Physical Review D (Particles and Fields), 46 (2):858–860, July 15, 1992. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.46.858.

Cougo-Pinto:1993:AIP

[CP93] M. V. Cougo-Pinto. The algebras of inhibited parafermionic violation of the Pauli principle. Journal of Mathematical Physics, 34(3):1110–1124, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v34/i3/p1110_s1.

Cowan:1956:DFN

[CRH^{+}56] C. L. Cowan, Jr., F. Reines, F. B. Harrison, H. W. Kruse, and A. D. McGuire. Detection of the free neutrino: a confirmation. Science, 1224(3212):103–104, July 20, 1956. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:/

/www.sciencemag.org/content/124/3212/103.extract. See

also the first report of detection of the neutrino [RC53].

Casimir:2010:HRH

[CS10] H. B. G. (Hendrik Brugt Gerhard) Casimir and Frans Saris. Haphazard reality: half a century of science. Amsterdam University Press, Amsterdam, The Netherlands, 2010. ISBN 90-8964-200-5. xii + 356 pp. LCCN ???? With a new preface by Frans Saris.

Cotaescu:2002:HDP

[CV02] Ion I. Cotaescu and Mihai Visinescu. Hierarchy of Dirac, Pauli, and Klein–Gordon conserved operators in Taub–NUT background. Journal of Mathematical Physics, 43(6):2978–2987, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deBeauregard:1986:AOBc

[dB86] O. Costa de Beauregard. Analyses d’ouvrages: The creation of Quantum Mechanics and the Bohr–Pauli Dialogue, (^{≪}Studies in the History of Modern Science ^{≫}, 14) par John Hendry. Revue d’Histoire des Sciences, 39(4):367–368, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632444.

Doughty:1986:MSM

[DC86] Noel A. Doughty and Graham P. Collins. Multispinor symmetries for massless arbitrary spin Fierz–Pauli and Rarita– Schwinger wave equations. Journal of Mathematical Physics, 27 (6):1639–1645, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i6/p1639_s1.

DeGregorio:2014:BWD

[De 14] Alberto De Gregorio. Bohr’s way to defining complementarity. Studies in History and Philosophy of Modern Physics, 45(??): 72–82, February 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219813000865.

Anton:1989:IQP

[Deg89] Peter Anton Degen. Interpretations of quantum physics, the mystical and the paranormal: Einstein, Schrödinger, Bohr, Pauli and Jordan. Ph.D. dissertation, Drew University, Madison, NJ, USA, December 1989. 319 pp.

DellAntonio:1961:CBS

[Del61] G. F. Dell’Antonio. On the connection between spin and statistics. Annals of Physics, 16(2):153–157, November 1961.

CODEN APNYA6. ISSN 0003-4916 (print), 1096-035X (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0003491661900318.

Delbruck:1972:W

[Del72] Max Delbrück. Out of this world. In Reines [Rei72], pages 280–288. ISBN 0-87081-025-1. LCCN QC780 .C65.

Dennison:1927:NSH

[Den27] David M. Dennison. A note on the specific heat of the hydrogen molecule. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 115(771):483–486, July 1, 1927. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.

jstor.org/stable/94849.

Dover:1974:PPPa

[DET74a] C. B. Dover, D. J. Ernst, and R. M. Thaler. Pauli Principle and the pion–nucleus interaction. Physical Review Letters, 32(10): 557–560, March 11, 1974. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.

aps.org/doi/10.1103/PhysRevLett.32.557.

Dover:1974:PPPb

[DET74b] C. B. Dover, D. J. Ernst, and R. M. Thaler. Pauli Principle and the pion-nucleus interaction. Physical Review Letters, 32 (16):913, April 22, 1974. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.

aps.org/doi/10.1103/PhysRevLett.32.913.

Dyson:2001:LEA

[DGB^{+}01] Freeman Dyson, Kurt Gottfried, Luis J. Boya, Karl von Meyenn, and Engelbert L. Schucking. Letters to the editor: Another visit with Wolfgang Pauli. Physics Today, 54(8):11–13, August 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v54/
i8/p11/s1.

deGuise:2000:GME

[dGPS00] H. de Guise, J. Patera, and R. T. Sharp. SL(3, C) generator matrix elements in a Pauli subgroup basis. Journal of Mathematical Physics, 41(7):4860–4880, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Debrus:1991:GTP

[DH91] Joachim Debrus and Allen C. Hirshfeld, editors. Geometry and Theoretical Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-540-53570-5 (Berlin), 0-387-53570-5 (New York). x + 323 pp. LCCN QC20.7.G44 G47 1991.

Dirac:1927:QTE

[Dir27] P. A. M. Dirac. The quantum theory of the emission and absorption of radiation. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 114(767): 243–265, March 1, 1927. CODEN PRLAAZ. ISSN 0080-4630.

URL http://www.jstor.org/stable/94746. Russian transla-

tion in [?, Vol. II, pp. 285ff].

Dittrich:2015:PWA

[Dit15] Walter Dittrich. On the Pauli–Weisskopf anti-Dirac paper. European Physical Journal H, 40(2):261–278, March 2015. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2015-60006-1.

Dyson:1967:SM

[DL67] Freeman J. Dyson and A. Lenard. Stability of matter. I. Journal of Mathematical Physics, 8(3):423–434, March 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i3/p423_s1.

Dover:1976:PNS

[DL76] Carl B. Dover and R. H. Lemmer. Pion–nucleus scattering and the Pauli Principle. Physical Review C (Nuclear Physics), 14 (6):2211–2225, December 1, 1976. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490x, 1538-4497. URL http://link.

aps.org/doi/10.1103/PhysRevC.14.2211.

delaPena-Auerbach:1971:STQ

[dlPA71] L. de la Peña-Auerbach. Stochastic theory of quantum mechanics for particles with spin. Journal of Mathematical Physics, 12 (3):453–461, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i3/p453_s1.

Duck:1997:PSS

[DPS97] Ian Duck, Wolfgang Pauli, and E. C. G. Sudarshan. Pauli and the spin-statistics theorem. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-3114-8. x + 512 pp. LCCN QC793.3.S6 D83 1997.

deRegt:1999:PVH

[dR99] Henk W. de Regt. Pauli versus Heisenberg: a case study of the heuristic role of philosophy. Foundations of Science, 4 (4):405–426, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1009643801616.

DeBoer:1958:RIU

[DT58] J. De Boer and F. Trendelenburg. Recommendations of the International Union of Pure and Applied Physics Commission for Symbols, Units and Nomenclature/Additions to Ch. I: General Principles and Recommendations/Additions to Ch. II: Symbols for Physical Quantities/Addition to Ch. III: Mathematical Operations and Symbols/Additions to Ch. IV: Units/Misprints to Be Corrected in U.I.P. 6 (1955)/Max-Planck-Medaille an W. Pauli. Physikalische Blätter, 14(6):259–262, 1958. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Dale-Trotter:2018:BRQ

[DT18] Iain Dale-Trotter. Book review: The quantum heretics: What is Real: the Unfinished Quest for the Meaning of Quantum Physics by Adam Becker. Physics World, 31(12):??, December 2018. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL https://physicsworld.com/a/the-
quantum-heretics/.

Heathcote:1953:NPW

[dVNHdVH53] Niels H. de V. (Niels Hugh de Vaudrey) Heathcote. Nobel Prize winners in physics, 1901-1950, volume 30 of The Life of science library. H. Schuman, New York, NY, USA, 1953. 473 pp. LCCN QC15 .H4. With a foreword by Herbert Dingle.

Heathcote:1971:NPW

[dVNHdVH71] Niels H. de V. (Niels Hugh de Vaudrey) Heathcote. Nobel Prize winners in physics, 1901–1950. Essay index reprint series. Books for Libraries Press, Freeport, NY, USA, 1971. ISBN 0-8369-2455-X. xvi + 473 pp. LCCN QC15 .H4 1971. With a foreword by Herbert Dingle.

Dyson:1960:BRBa

[Dys60] Freeman J. Dyson. Book review: Theory of Relativity. By W. Pauli. Translated by G. Field from 1921 German article, with 25 pp. of supplementary notes by Pauli. 241 pp. Pergamon Press, London & New York, 1958. $6.00. Physics Today, 13(6):46, June 1960. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v13/
i6/p46/s1.

Dyson:2002:BRC

[Dys02] Freeman J. Dyson. Book review: The conscience of physics: The life, work and dreams of Wolfgang Pauli [No time to be brief: a scientific biography of Wolfgang Pauli: Charles P. Enz; Oxford University Press, New York, 2002, pp. 581, price US$60.00, £35.00, ISBN 0-19-856479-1]. Nature, 420(6916):607–608, December 12, 2002. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://labs.adsabs.harvard.
edu/ui/abs/2002Natur.420..607D;   http://www.nature.
com/nature/journal/v420/n6916/full/420607a.html.

Esposito:2015:PEM

[EAW15] Salvatore Esposito, E. (Evgeny) Akhmedov, and Frank Wilczek. The physics of Ettore Majorana: phenomenological, theoretical, and mathematical. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-04402-2 (hardcover), 1-316-19108-7 (PDF ebook). xi + 382 pp. LCCN QC19.6 .E87 2015.

Eckart:1926:OCS

[Eck26] Carl Eckart. Operator calculus and the solution of the equations of quantum dynamics. Physical Review, 28(4):711–726, October 1926. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRev.28.711.

Eckart:1970:CEP

[Eck70] Carl Eckart. Carl Eckart papers, 1921–1973 (bulk 1935–1970). 1970. 8 (4002413) pp. LCCN Oversize 0113K; 0810D. URL

http://hdl.loc.gov/loc.mss/eadmss.ms001014.

Eckart:1973:CEP

[Eck73] Carl Eckart. Carl Eckart papers, 1921–1973 (bulk 1935–1970). US Library of Congress archival manuscript material (collec-

tion)., 1973. URL http://hdl.loc.gov/loc.mss/eadmss.
ms001014.

Eisenberg:1990:WPM

[EFK90] W. Eisenberg, M. Füting, and E. Krause. Wolfgang Pauli — mehr als nur ein Klassiker der modernen Physik. (German) [Wolfgang Pauli — more than just a classic of modern physics]. Wissenschaftliche Zeitschrift — Martin-Luther-Universität Halle-Wittenberg, Mathematisch-Naturwissenschaftliche Reihe, 39(2):127–136, ???? 1990. CODEN WMHMAP. ISSN 0138-1504.

Enz:1997:WPS

[EGO97] Charles P. (Charles Paul) Enz, Beat Glaus, and Gerhard Oberkofler, editors. Wolfgang Pauli und sein Wirken an der ETH Zürich: aus den Dienstakten der Eidgenössischen Technischen Hochschule. (German) [Wolfgang Pauli and his works at the Technical University of Zürich: from the Service of the Technical University]. VDF Hochschulverlag AG an der ETH Zürich, Zürich, Switzerland, 1997. ISBN 3-7281-2317-X. xi + 463 pp. LCCN QC16.P37 W66 1997.

Ehrenfest:1927:RBRb

[Ehr27a] P. Ehrenfest. Relation between the reciprocal impenetrability of matter and Pauli’s Exclusion Principle: a correction. Nature, 119(2999):602, April 23, 1927. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v119/n2999/pdf/119602c0.pdf. See

[Ehr27b].

Ehrenfest:1927:RBRa

[Ehr27b] Paul Ehrenfest. Relation between the reciprocal impenetrability of matter and Pauli’s Exclusion Principle. Nature, 119(2988): 196, February 5, 1927. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v119/n2988/pdf/119196a0.pdf. See correc-

tion [Ehr27a].

Ehrenfest:1931:AVL

[Ehr31] Paul Ehrenfest. Ansprache zur Verleihung der Lorentzmedaille an Professor Wolfgang Pauli am 31. Oktober 1931. (German) [Speech on the award of the Lorentz Medal to Professor Wolfgang Pauli on 31 October 1931]. Versl. Akad. Amsterdam, 43 (??):121–126, ???? 1931.

Einstein:1922:WPR

[Ein22] Albert Einstein. W. Pauli: Relativitätstheorie. (German) [Review of W. Pauli: The Theory of Relativity]. Naturwissenschaften, 10(??):184–185, 1922. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Einstein:1941:DNE

[Ein41] A. Einstein. Demonstration of the non-existence of gravitational fields with a non-vanishing total mass free of singularities. Univ. Nac. Tucumán. Revista A., 2:5–15, 1941.

Einstein:1979:LWP

[Ein79] Albert Einstein. Letter to W. Pauli, 19 September 1938, AE 19-175. In Hermann et al. [HvMW79], page ?? ISBN 0-387-08962-4, 3-540-08962-4. LCCN QC16.P37 W64. URL http://

www.springer.com/physics/book/978-3-540-08962-9.

Eisenberg:1991:WPW

[Eis91] W. Eisenberg. Zu den Wirklichkeitsauffassungen der Physiker Wolfgang Pauli und Ludwig Boltzmann. (German) [The views of reality: Physicists Wolfgang Pauli and Ludwig Boltzmann]. In Röseberg [Rös91], pages 47–54. ISBN 3-05-501349-

2. LCCN ???? URL http://www.gbv.de/dms/hbz/toc/
ht003858637.PDF;   http://www.zentralblatt-math.org/
zmath/en/search/?an=0751.00006.

Eisner:2000:RBP

[Eis00] Werner Eisner. Rezension: Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Herausgegeben von Karl von Meyenn (Band IV, Teil I: 1950–1952; Teil II: 1953–1954). Berlin/Heidelberg/New York: Springer 1996 und 1999; 968 und 1100 Seiten. (German) [Wolfgang Pauli: Scientific Correspondence with Bohr, Einstein, Heisenberg, and others. Edited by Karl of Meyenn (Volume IV, part I: 1950–1952, Part II: 1953–1954). Berlin / Heidelberg / New York: Springer 1996 and 1999, 968 and 1100 pages]. Berichte zur Wissenschaftsgeschichte, 23(1):59–62, ???? 2000. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Eisner:2002:RWB

[Eis02] Werner Eisner. Rezension: Wissenschaftlicher Briefwechsel. Bd IV, Teil III: 1955–1956 von Wolfgang Pauli. (German) [Review: Scientific correspondence. Vol. IV, Part III: 1955–1956 by Wolfgang Pauli]. Berichte zur Wissenschaftsgeschichte, 25(3): 193–194, September 2002. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Ellis:1948:CSN

[Ell48] Cecil B. Ellis. On closed shells in nuclei. Physical Review (2), 74(10):1547, November 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.74.1547.

Enz:1988:WPG

[EM88] Charles P. (Charles Paul) Enz and K. V. Meyenn, editors. Wolfgang Pauli: das Gewissen der Physik. (German) [Wolfgang Pauli: the conscience of physics]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1988. ISBN 3-528-08993-8. xii + 546 pp. LCCN QC16.P37 W65 1988.

Enz:1973:WPS

[Enz73] Charles P. Enz. W. Pauli’s scientific work. In Mehra [Meh73], pages 766–799. ISBN 90-277-0345-0, 90-277-2536-5. LCCN

QC173.96 .S95 1972. URL http://www.springer.com/us/
book/9789027703453.

Enz:1981:YAP

[Enz82] Charles P. Enz. 50 years ago Pauli invented the neutrino. Helvetica Physica Acta, 54(3):411–418, ???? 1981–1982. CODEN HPACAK. ISSN 0018-0238. URL http://retro.seals.ch/

digbib/view?rid=hpa-001:1981:54::700&id=hitlist.

Enz:1985:STF

[Enz85a] Charles P. Enz. The space, time and field concepts in Wolfgang Pauli’s work. In Lahti and Mittelstaedt [LM85], pages 127–145. ISBN 9971-5-0004-3. LCCN QC173.96 .S96 1985.

Enz:1985:WPP

[Enz85b] Charles P. Enz. Wolfgang Pauli, physicist and philosopher. In Lahti and Mittelstaedt [LM85], pages 241–255. ISBN 9971-5-0004-3. LCCN QC173.96 .S96 1985.

Enz:1994:PHT

[Enz94a] Charles P. Enz. Preface: The history of this translation: Paul Rosbaud, friend and publisher of Wolfgang Pauli. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], pages 1–

6. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York).

LCCN QC6.2 .P38 1994. URL http://www.loc.gov/catdir/
enhancements/fy0812/94015098-d.htm.

Enz:1994:WPB

[Enz94b] Charles P. Enz. Wolfgang Pauli (1900–1958): A biographical introduction. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], pages 13–26. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2 .P38 1994. URL http:

//www.loc.gov/catdir/enhancements/fy0812/94015098-d.
htm.

Enz:2000:WPB

[Enz00] C. P. Enz. Wolfgang Pauli’s 100th birthday: a personal view by Prof. C. P. Enz, W. Pauli’s last assistant. Europhysics News, 31 (4):12, July/August 2000. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (electronic).

Enz:2002:NTB

[Enz02] Charles P. (Charles Paul) Enz. No time to be brief: a scientific biography of Wolfgang Pauli. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-856479-1. viii + 573 pp. LCCN QC16.P37 E59 2002. URL http://www.loc.

gov/catdir/enhancements/fy0614/2002726902-d.html;
http://www.loc.gov/catdir/enhancements/fy0614/2002726902-
t.html.

Enz:2005:PGB

[Enz05] Charles P. (Charles Paul) Enz. ,,Pauli hat gesagt“: eine Biografie des Nobelpreisträgers Wolfgang Pauli, 1900–1958. (German) [Pauli Said: a Biography of the Nobel Prize winner Wolfgang Pauli, 1900–1958]. Verlag Neue Zürcher Zeitung, Zürich, Switzerland, 2005. ISBN 3-03823-144-4. 167 pp. LCCN QC16.P37 E593 2005.

Einstein:1943:NER

[EP43] Albert Einstein and Wolfgang Pauli. On the non-existence of regular stationary solutions of relativistic field equations. Annals of Mathematics (2), 44(2):131–137, April 1943. CODEN ANMAAH. ISSN 0003-486X (print), 1939-

8980 (electronic). URL http://links.jstor.org/sici?
sici=0003-486X%28194304%292%3A44%3A2%3C131%3AOTNORS%
3E2.0.CO%3B2-Q.

Esposito:2007:USM

[Esp07] Salvatore Esposito. An unknown story: Majorana and the Pauli–Weisskopf scalar electrodynamics. Annalen der Physik (1900), 16(12):824–841, December 2007. ISSN 1521-3889.

Enz:1960:NAV

[ET60] C. P. Enz and A. Thellung. Nullpunktsenergie und Anordnung nicht vertauschbarer Faktoren im Hamiltonoperator. (German) [Zero-point energy and arrangement of noncommuting factors in the Hamiltonian]. Helvetica Physica Acta, 33(8): 839–848, ???? 1960. CODEN HPACAK. ISSN 0018-

0238. URL http://retro.seals.ch/digbib/view?rid=hpa-

001:1960:33::1103. Comments about Pauli’s views of the zero-point energy. Intended as a birthday gift to Pauli, but not completed before his death.

EONR:1969:CRS

[Eur69] European Organization for Nuclear Research . Catalogue of reprints of scientific papers in the Pauli collection, volume 3 of CERN bibl. CERN, Geneva, Switzerland, 1969. v + 470 pp. LCCN Z7145 .E87.

Enz:1994:WPW

[EvM94] Charles P. (Charles Paul) Enz and K. von Meyenn. Wolfgang Pauli: Writings on physics and philosophy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). vi + 289 pp. LCCN QC6.2 .P38 1994. URL http://www.loc.

gov/catdir/enhancements/fy0812/94015098-d.htm.

Fara:2001:GPI

[Far01] Patricia Fara. Group portraits IV — the Seventh Solvay Conference [Bruxelles, 22–29 October 1933]. Endeavour, 25(4):

137–138, December 2001. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Fermi:1924:CSQ

[Fer24] Enrico Fermi. Considerazioni sulla quantizzazione dei sistemi che contengono degli elementi identici. (Italian) [Considerations on the quantization of systems that contain identical elements]. Il Nuovo Cimento (8), 1(1):145–152, ???? 1924. CODEN NU-CIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL

http://link.springer.com/article/10.1007/BF02956745;
 http://www.springerlink.com/content/97203r302q8w7unj/
fulltext.pdf.

Fermi:1926:QIE

[Fer26] Enrico Fermi. Zur Quantelung des idealen einatomigen Gases. (German) [On the quantization of an ideal monatomic gas]. Zeitschrift für Physik, 36(11–12):902–912, November 1926. CO-DEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01400221;   http://www.springerlink.com/content/k763270092273181/

.

Fermi:1934:TTR

[Fer34] Enrico Fermi. Tentativo di una teoria dei raggi β. (Italian) [Draft of a theory of β rays]. Il Nuovo Cimento (8), 11 (1):1–19, ???? 1934. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://inspirehep.net/

record/25382;   http://www.springerlink.com/content/
2t3p38675v840182/.

Friedrichs:1987:SPM

[FF87] Mark S. Friedrichs and Richard A. Friesner. Solution of the Pauli master equation via the Lanczos algorithm. Chemical Physics Letters, 137(3):285–290, June 12, 1987. CODEN CH-PLBC. ISSN 0009-2614 (print), 1873-4448 (electronic).

Fierz:1939:RTK

[Fie39] M. Fierz. Über die relativistische theorie kräftefreier teilchen mit beliebigem spin. (German) [On the relativistic theory of force-free particles with arbitrary spin]. Helvetica Physica Acta, 12(1):3–17, ???? 1939. CODEN HPACAK. ISSN 0018-

0238. URL http://retro.seals.ch/digbib/view?rid=hpa-
001:1939:12::630&id=hitlist.

Fierz:1959:WP

[Fie59] Marcus E. Fierz. Wolfgang Pauli (1900–1958). Nuclear Physics, 10:1–5, February/May 1959. CODEN NUPHA7. ISSN 0029-5582 (print), 1873-1546 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0029558259901786.

Finke:1949:BSP

[Fin49] W. Finke. über ein symmetrisches Periodensystem der chemischen Elemente nach dem Pauli-Prinzip. (German) [On a symmetric periodic table of chemical elements according to the Pauli Principle]. Zeitschrift für Physik, 126(1–2):106–107, January 1949. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831

(electronic). URL http://link.springer.com/article/10.
1007/BF01331045;   http://www.springerlink.com/content/
t0pw3kj711518202/fulltext.pdf.

Finegold:1992:WDP

[Fin92] Leonard X. Finegold. Where did Pauli make his ‘wrong’ remark? Physics Today, 45(9):103, September 1992. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/link.aip.org/link/phtoad/v45/i9/p103/s1. See response

[Pei92].

Fischbeck:1999:SIZ

[Fis99] Hans-Jürgen Fischbeck, editor. Vom Sinn im Zufall: zum Dialog zwischen Wolfgang Pauli und C. G. Jung; 8. bis 10. Mai 1998: Tagung. (German) [In the sense of coincidence: the dialogue between Wolfgang Pauli and C. G. Jung; 8 to 10 May 1998: Conference], volume 98,5 of Begegnungen Mit dem heutigen Wissen den Glauben denken. Evang. Akad., Mülheim an der Ruhr, Germany, 1999. ISBN 3-926188-86-3. LCCN ????

Fischer:2004:BKW

[Fis04] Ernst Peter Fischer. Brücken zum Kosmos: Wolfgang Pauli — Denkstoffe und Nachtträume zwischen Kernphysik und Weltharmonie. (German) [Bridge to the Cosmos: Wolfgang Pauli — Thoughts and dreams between nuclear physics and world harmony]. Libelle-Verlag, Lengwil, Switzerland, 2004. ISBN 3-909081-44-4. 219 pp. LCCN ???? ca. EUR 18.80, ca. EUR

AT 19.50, ca. CHF 32.50. URL http://www.dandelon.com/
intelligentSEARCH.nsf/alldocs/CDA0A492E9F8C293C1256F9600570686/

.

Fischer:2010:HQE

[Fis10] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger]. Herbig, München, Germany, 2010. ISBN 3-7766-2643-7. 350 pp. LCCN ????

Fischer:2011:BKW

[Fis11] Ernst Peter Fischer. Brücken zum Kosmos: Wolfgang Pauli — Denkstoffe und Nachtträume zwischen Kernphysik und Weltharmonie. (German) [Bridge to the Cosmos: Wolfgang Pauli — Thoughts and dreams between nuclear physics and world harmony]. Libelle-Verlag, Lengwil, Switzerland, second edition, 2011. ISBN 3-909081-44-4. 219 pp. LCCN ????

Fischer:2012:HQE

[Fis12] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger], volume 19406 of Fischer. Fischer-Taschenbuch-Verlag, Frankfurt am Main, Germany, 2012. ISBN 3-596-19406-7. 350 pp. LCCN ????

Fleming:2007:EPE

[Fle07] Gordon N. Fleming. The evolution of Pauli’s exclusion principle. Studies in History and Philosophy of Modern Physics, 38(1): 202–208, March 2007. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219806000712.

Fontana:1968:OTE

[FM68a] P. R. Fontana and W. J. Meath. One- and two-center expansions of the Breit–Pauli Hamiltonian. Journal of Mathematical Physics, 9(9):1357–1364, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fontana:1968:OTC

[FM68b] Peter R. Fontana and William J. Meath. One- and two-center expansions of the Breit–Pauli Hamiltonian. Journal of Mathematical Physics, 9(9):1357–1364, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i9/p1357_s1.

Fierz:1938:TEL

[FP38] M. Fierz and Wolfgang Pauli. Zur Theorie der Emission langwelliger Lichtquanten. (German) [On the theory of emission of long-wavelength light quanta]. Nuovo Cimento, 15(??):167–188, ???? 1938.

Fierz:1939:RWE

[FP39a] M. Fierz and W. Pauli. On relativistic wave equations for particles of arbitrary spin in an electromagnetic field. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 173:211–232, 1939. CODEN PRLAAZ. ISSN 0080-4630.

Fierz:1939:RFT

[FP39b] M. Fierz and Wolfgang Pauli. Über relativistische Feldgleichungen von Teilchen mit beliebigem Spin im elektromagnetischen Feld. (German) [On relativistic field equations of particles with arbitrary spin in an electromagnetic field]. Helvetica Physica Acta, 12(4):297–300, ???? 1939. CODEN HPACAK. ISSN

0018-0238. URL http://retro.seals.ch/digbib/view?rid=
hpa-001:1939:12::287&id=&id2=&id3=.

Fernandez:2013:SPP

[FR13a] Bernard Fernandez and Georges Ripka. 1925: Spin and the Pauli Principle. In Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956 [FR13b], page 121. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). LCCN QC773 .F47 2013.

Fernandez:2013:UMA

[FR13b] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896– 1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). xviii + 522 pp. LCCN QC773 .F47 2013.

Fernandez:2013:WP

[FR13c] Bernard Fernandez and Georges Ripka. Wolfgang pauli. In Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956 [FR13b], page 121. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). LCCN QC773 .F47 2013.

Franklin:2004:TRN

[Fra04] Allan Franklin. Are there really neutrinos?: an evidential history. Perseus Publishers, Cambridge, MA, USA, 2004. ISBN 0-7382-0265-7, 0-8133-4128-0 (paperback). ix + 371 pp. LCCN

QC793.5.N42 F73 2004. URL http://www.loc.gov/catdir/
enhancements/fy0837/2008530710-d.html;   http://www.
loc.gov/catdir/toc/fy0805/2008530710.html.

Franzen:2005:PO

[Fra05] Wolfgang Franzen. Pauli oscillations. Physics Today, 58(2): 14, February 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v58/i2/p14/s1.

Frenkel:1928:APF

[Fre28] J. Frenkel. Anwendung der Pauli-Fermischen Elektronengastheorie auf das Problem der Kohäsionskräfte. (German) [Application of the Pauli–Fermi electron gas theory to the problem of the cohesive forces]. Zeitschrift für Physik, 50(3–4):234–248, March 1928. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831

(electronic). URL http://link.springer.com/article/10.
1007/BF01328867;   http://www.springerlink.com/content/
l1549828472413v0/fulltext.pdf.

Frenkel:1930:CFP

[Fre30] J. Frenkel. On the correct formulation of Pauli’s Exclusion Principle. Nature, 125(3146):235–236, February 15, 1930. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v125/n3146/
pdf/125235b0.pdf.

FreireJunior:2015:QDR

[Fre15] Olival Freire Junior. The Quantum Dissidents: Rebuilding the Foundations of Quantum Mechanics (1950–1990). Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-662-44661-8, 3-662-44662-6 (e-book). xvi + 356 pp. LCCN QC173.98 .F74 2015.

Filippov:1995:CBP

[FRKK95] G. F. Filippov, I. Yu. Rybkin, S. V. Korennov, and K. Katō. On the complete basis of Pauli-allowed states of three-cluster systems in the Fock–Bargmann space. Journal of Mathematical Physics, 36(9):4571–4589, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fischer:2006:EPF

[FS06] Ernst Peter Fischer and Klaus Sander. Ernst Peter Fischer erzählt, Paarläufe der Wissenschaft. (German) [Ernst Peter Fischer tells the pair runs of science]. Suppoš, Köln, Germany, 2006. ISBN 3-932513-68-1. LCCN ???? URL http://deposit.

ddb.de/cgi-bin/dokserv?id=2758637\%26prov=M\%26dok\_
var=1\%26dok\_ext=htm. 4 CDs.

Fierz:1959:CWP

[FW59] Marcus E. Fierz and Victor F. Weisskopf. In commemoration of Wolfgang Pauli. Physics Today, 12(7):16–19, July 1959. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://link.aip.org/link/phtoad/v12/i7/
p16/s1.

Fierz:1960:TPT

[FW60] Markus Fierz and Victor Frederick Weisskopf. Theoretical physics in the Twentieth Century: a memorial volume to Wolfgang Pauli. Interscience Publishers, New York, NY, USA, 1960. x + 328 pp. LCCN QC3 .F52. See [Gou61] for comments on some inaccuracies in this book, and on the history of the discovery of the spin of the electron and the nucleus.

Feldman:2007:MSE

[FW07] Burton Feldman and Katherine Williams. 112 Mercer Street: Einstein, Russell, Gödel, Pauli, and the end of innocence in science. Arcade Publishing, New York, NY, USA, 2007. ISBN 1-55970-704-6. xx + 243 pp. LCCN Q141 .F345 2007. URL http:

//www.loc.gov/catdir/toc/ecip078/2007001194.html.

Farazdel:1985:VMV

[FWS^{+}85] Abbas Farazdel, William M. Westgate, Alfredo M. Simas, Robin P. Sagar, and Vedene H. Smith, Jr. Validity of the mass-velocity term in the Breit–Pauli Hamiltonian. International Journal of Quantum Chemistry, 28(S19):61–68, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gamow:1966:TYSa

[Gam66a] George Gamow. Thirty years that shook physics: the story of quantum theory. Doubleday, Garden City, NY, USA, 1966. xvi + 224 pp. LCCN QC174.1 .G3 1966b.

Gamow:1966:TYSb

[Gam66b] George Gamow. Thirty years that shook physics: the story of quantum theory, volume S45 of Science study series. Anchor Books, New York, NY, USA, 1966. xvi + 224 pp. LCCN QC174.1 .G3.

Gamow:1972:TYS

[Gam72] George Gamow. Thirty years that shook physics: the story of quantum theory, volume 38 of Science study series. Heinemann Educational, London, UK, 1972. ISBN 0-435-55071-3. viii + 224 + 12 pp. LCCN QC173.98 .G35 1972.

Gamow:1985:TYS

[Gam85] George Gamow. Thirty Years That Shook Physics: the Story of Quantum Theory. Dover, New York, NY, USA, 1985. ISBN 0-486-24895-X (paperback). xiv + 224 + 9 pp. LCCN QC174.12

.G35 1985. US$4.95. URL http://www.loc.gov/catdir/
description/dover032/85006797.html.

Garrett:1963:BCN

[Gar63a] Alfred Benjamin Garrett. Birth of the concept that neutrinos must exist: Wolfgang Pauli. In The flash of genius [Gar63b], pages 204–207. LCCN Q125 .G38.

Garrett:1963:FG

[Gar63b] Alfred Benjamin Garrett. The flash of genius. Van Nostrand, Princeton, NJ, USA, 1963. x + 249 pp. LCCN Q125 .G38.

Gauderis:2014:ENP

[Gau14] Tjerk Gauderis. To envision a new particle or change an existing law? Hypothesis formation and anomaly resolution for the curious case of the β decay spectrum. Studies in History and Philosophy of Modern Physics, 45(??):27–45, February 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000920.

Goudsmit:1927:FTW

[GB27] S. Goudsmit and E. Back. Feinstrukturen und Termordnung des Wismutspektrums. (German) [Fine structure and term order of the bismuth spectrum]. Zeitschrift für Physik, 43(5–6):321–334, May 1927. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/
10.1007/BF01397446;   http://www.springerlink.com/content/
h5w4160831u31v16/.

Goudsmit:1929:PBE

[GB29] S. Goudsmit and R. F. Bacher. The Paschen–Back effect of hyperfine structure. Physical Review (2), 34(11):1499–1500, December 1929. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.34.1499;   http://prola.aps.org/abstract/PR/
v34/i11/p1499_1.

Goudsmit:1930:PBE

[GB30] S. Goudsmit and R. F. Bacher. Der Paschen–Back-Effekt der Hyperfeinstruktur. (German) [The Paschen–Back effect in hyperfine structure]. Zeitschrift für Physik, 66(1–2):13–30, January 1930. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/
10.1007/BF01397522;   http://www.springerlink.com/content/
p763u480j8281v62/.

Galvagno:2005:PPC

[GG05] Mariano Galvagno and Gastón Giribet. The particle problem in classical gravity: a historical note on 1941. European Journal of Physics, 26(6):S97, 2005. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL http://stacks.iop.org/ 0143-0807/26/i=6/a=S03. Comments on [Ein41, EP43].

Gurtler:1975:CFD

[GH75] R. Gurtler and D. Hestenes. Consistency in the formulation of the Dirac, Pauli, and Schrödinger theories. Journal of Mathematical Physics, 16(3):573–584, March 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/
i3/p573_s1.

Greenberg:1999:QSC

[GH99] O. W. Greenberg and Robert C. Hilborn. Quon statistics for composite systems and a limit on the violation of the Pauli Principle for nucleons and quarks. Physical Review Letters, 83(22): 4460–4463, November 29, 1999. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://

link.aps.org/doi/10.1103/PhysRevLett.83.4460.

GarciaDoncel:1987:SPP

[GHMP87] Manuel García Doncel, A. Hermann, L. Michel, and A. Pais, editors. Symmetries in physics (1600–1980): proceedings of the 1st International Meeting on the History of Scientific Ideas held at Sant Feliu de Guíxols, Catalonia, Spain, September 20–26, 1983. Seminari d’História de les Ciéncies, Universitat Autónoma de Barcelona, Bellaterra, Barcelona, Spain, 1987. ISBN 84-7488-148-X. LCCN QC174.17.S9 I57 1983.

Gibson:2019:SIH

[Gib19] Susannah Gibson. The Spirit of Inquiry: How One Extraordinary Society Shaped Modern Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2019. ISBN 0-19-883337-7 (hardcover). xxi + 377 pp. LCCN Q41.C194 G537 2019.

Gieser:1990:BRB

[Gie90] Suzanne Gieser. Book review: Beyond the Atom: The Philosophical Thought of Wolfgang Pauli by K. V. Laurikainen. Isis, 81(2):377–378, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233777.

Gieser:2005:IKD

[Gie05] Suzanne Gieser. The innermost kernel: depth psychology and quantum physics: Wolfgang Pauli’s dialogue with C. G. Jung. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-20856-9. xii + 378 pp.

LCCN QC15 .G52 2005. URL http://www.loc.gov/catdir/
toc/fy0614/2004113652.html.

Gilder:2009:AEW

[Gil09] Louisa Gilder. The Age of Entanglement: When Quantum Physics Was Reborn. Vintage Books, New York, NY, USA, 2009. ISBN 1-4000-9526-3. xviii + 443 pp. LCCN QC174.12 .G528 2009.

Glauber:2001:RP

[Gla01] Roy Glauber. A remembrance of Pauli in 1950. Physics Today, 54(2):49, February 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v54/i2/p49/s1.

Gleason:1986:MP

[Gle86] W. B. Gleason. Memories of Pauli. Physics Today, 39(8):11, August 1986. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v39/i8/p11/s2.

Guerra:2014:WEC

[GLR14] Francesco Guerra, Matteo Leone, and Nadia Robotti. When energy conservation seems to fail: The prediction of the neutrino. Science & Education (Springer), 23(6):1339–1359, June 2014. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Greenberg:1987:LQF

[GM87] O. W. Greenberg and R. N. Mohapatra. Local quantum field theory of possible violation of the Pauli Principle. Physical Review Letters, 59(22):2507–2510, November 30, 1987. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.59.2507.

Greenberg:1988:LQF

[GM88] O. W. Greenberg and R. N. Mohapatra. Local quantum field theory of possible violation of the Pauli Principle. Physical Review Letters, 61(12):1432, September 19, 1988. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.61.1432.

Greenberg:1989:DLQa

[GM89a] O. W. Greenberg and R. N. Mohapatra. Difficulties with a local quantum field theory of possible violation of the Pauli Principle. Physical Review Letters, 62(7):712–714, February 13, 1989.

CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.62.712.

Greenberg:1989:DLQb

[GM89b] O. W. Greenberg and R. N. Mohapatra. Difficulties with a local quantum field theory of possible violation of the Pauli Principle. Physical Review Letters, 62(16):1927, April 17, 1989. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.62.1927.2.

GoeppertMayer:1948:CSN

[Goe48] Maria Goeppert Mayer. On closed shells in nuclei. Physical Review (2), 74(3):235–239, August 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

link.aps.org/doi/10.1103/PhysRev.74.235.

GoeppertMayer:1949:CSN

[Goe49] Maria Goeppert Mayer. On closed shells in nuclei. II. Physical Review (2), 75(12):1969–1970, June 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:

//link.aps.org/doi/10.1103/PhysRev.75.1969;   http:
//prola.aps.org/abstract/PR/v75/i12/p1969_1;   http:
//www.nobelprize.org/nobel_prizes/physics/laureates/
1963/.

Gottstein:1983:CGE

[Got83] Klaus Gottstein. Commentary: The Göttingen Eighteen. Bulletin of the Atomic Scientists, 39(8):62–63, October 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1961:PNS

[Gou61] S. A. Goudsmit. Pauli and nuclear spin. Physics Today, 14 (6):18–21, June 1961. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v14/i6/p18/s1. See [FW60] for the book whose

inaccuracies are addressed in this article, which also records the author’s involvement in the historical discovery of the spin of the electron and the nucleus.

Goudsmit:1971:DES

[Gou71] Samuel A. Goudsmit. The discovery of the electron spin. Web document, April 1971. URL http://www.lorentz.

leidenuniv.nl/history/spin/goudsmit.html. English trans-

lation by J. H. van der Waals of Goudsmit’s lecture in Dutch for the golden jubilee of the Dutch Physical Society in April 1971.

Goudsmit:1976:FYS

[Gou76] Samuel A. Goudsmit. Fifty years of spin: It might as well be spin. Physics Today, 29(6):40–43, June 1976. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/

phtoad/v29/i6/p40_s1. See correspondence [RG76].

Guttinger:1931:HLT

[GP31] P. Güttinger and W. Pauli. Zur Hyperfeinstruktur von Li^{+}. Teil II. (German) [On the hyperfine structure of Li^{+}. Part II]. Zeitschrift für Physik, 67(11–12):743–765, November 1931. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01390756;   http://www.springerlink.com/content/t331151k247157v2/
fulltext.pdf.

Giannetto:2000:NSS

[GP00] E. A. Giannetto and F. Pozzi. Non-separability and synchronicity: Pauli, Jung and a new historical, philosophical perspective on quantum physics. In Garola and Rossi [GR00], pages 251–

259. ISBN 981-02-4262-X. LCCN QC173.96 .F663 2000.

Gamow:1931:DVO

[GR31] George Gamow and Léon Rosenfeld. On the determination of the velocity of an object moving in a fluid on the basis of a single photograph. Originally written in German, and reproduced in English translation in [Del72, pages 285–287]. Submitted to the journal Physica, but rejected by editor Paul Ehrenfest., June 7, 1931.

Garola:2000:FQM

[GR00] Claudio Garola and Arcangelo Rossi, editors. The foundations of quantum mechanics: historical analysis and open questions: Lecce, Italy, 13–16 October 1998. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2000. ISBN 981-02-4262-X. LCCN QC173.96 .F663 2000.

Gornitz:2001:BSM

[GRE^{+}01] Thomas Görnitz, Helmut Rechenberg, Berthold-Georg Englert, Thomas Walcher, Katja Bammel, Thomas W. Beneke, Wolfgang W. Schwippert, and Filip Floecel. Bücher und Software: Meyenn: Wolfgang Pauli. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg U. A./Kraus: Von der Uranspaltung zur Göttinger Erklärung. Otto Hahn, Werner Heisenberg und Carl Friedrich von Weizsäcker und die Verantwortung des Wissenschaftlers/Dürr: Bohmsche Mechanik als Grundlage der Quantenmechanik/Thomas/Weise: The Structure of the Nucleon/Sube: Langenscheidts Fachwörterbuch Physik. Deutsch-Englisch/ConceptDraw 1.6 Zeichnen mit intelligenten Objekten /Rath: Quantenphysik für Windows Version 1.5. Physikalische Blätter, 57(11):82–86, November 2001. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
20010571123/abstract.

Geyler:2004:POA

[GS04] V. A. Geyler and P. St’ovícek. On the Pauli operator for the Aharonov–Bohm effect with two solenoids. Journal of Mathematical Physics, 45(1):51–75, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i1/p51_s1.

Goudsmit:1926:RES

[GU26] S. Goudsmit and G. E. Uhlenbeck. Over het roteerende electron en de structuur der spectra. (Dutch) [On the rotating electron and the structure of the spectra]. Physica, 6(??):273–290, ???? 1926. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Gulmanelli:1954:STD

[Gul54] P. Gulmanelli. Su una Teoria dello Spin Isotropico. Sezione di Milano dell istituto Nazionale di Fisica Nucleare, Casa Editrice Pleion, Milano, Italy, 1954.

Hall:1986:PEP

[Hal86] Peter Joseph Hall. The Pauli Exclusion Principle and the foundations of chemistry. Synthese, 69(3):267–272, December 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00413974.

Haldane:1991:FSA

[Hal91] F. D. M. Haldane. “fractional statistics” in arbitrary dimensions: A generalization of the Pauli Principle. Physical Review Letters, 67(8):937–940, August 19, 1991. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL

http://link.aps.org/doi/10.1103/PhysRevLett.67.937.

Han:2010:SFG

[Han10] Gang Han. The symmetries of the fine gradings of sl(n^{k}, C) associated with direct product of Pauli groups. Journal of Mathematical Physics, 51(9):092104, September 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/
i9/p092104_s1.

Han:2011:WGF

[Han11] Gang Han. The Weyl group of the fine grading of ∫↕(n, C) associated with tensor product of generalized Pauli matrices. Journal of Mathematical Physics, 52(4):042109, April 2011. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i4/p042109_s1.

Hawking:2011:DSM

[Haw11] Stephen Hawking, editor. The dreams that stuff is made of: the most astounding papers on quantum physics — and how they shook the scientific world. Running Press, Philadelphia, PA, USA, 2011. ISBN 0-7624-3434-1. xi + 1071 pp. LCCN QC173.98 .D74 2011.

Haensel:1980:BFK

[HEB^{+}80] R. Haensel, Karl Ehrlich, K. Bethge, H. Soffel, W. von Witsch, and H. Rechenberg. Bücher Filme: Kunz: Synchrotron Radiation Techniques and Applications/Lehmann: Interaction of Radiation with Solids and Elementary Defect Production/ Olson u. Schumacher: Collective Ion Acceleration/Brosche u Sündermann: Tidal Friction and the Earth’s Rotation/ Haas u. Koch: Physik-Lehrbuch für Pharmazeuten und Mediziner/Hermann, Meyenn u. Weisskopf: Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a., Bd. I: 1919–1929. Physikalische Blätter, 36(10):318–320, October 1980. CODEN PHBLAG. ISSN 0031-9279 (print),

1521-3722 (electronic). URL http://onlinelibrary.wiley.
com/doi/10.1002/phbl.19800361013/abstract.

Heisenberg:1956:HSI

[Hei56] W. Heisenberg. Hilbert space II and the “ghost” states of Pauli and Källén. Il Nuovo Cimento (10), 4(supplemento):743–747, 1956. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Heisenberg:1959:WP

[Hei59] Werner Heisenberg. Wolfgang Pauli 25.4.1900–15.12.1958. In ????, editor, Bayerische Akademie der Wissenschaften. Jahrbuch 1959. (German) [Bavarian Academy of Science, Yearbook 1959], pages 183–187. Bayerische Akademie der Wissenschaften, Munich, West Germany, 1959. LCCN ????

Heisenberg:1960:EZE

[Hei60] Werner Heisenberg. Erinnerungen an die Zeit der Entwicklung der Quantenmechanik. (German) [Memories of the time of the evolution of quantum mechanics]. In Theoretical physics in the twentieth century (Pauli memorial volume), pages 40–47. Interscience Publishers, New York, NY, USA, 1960.

Heisenberg:1979:AHBb

[Hei79a] Werner Heisenberg. Anlage zu [414] (Bisherige und zukunftige Quantenelektrodynamik). Ist eine deterministische Ergänzung der Quantenmechanik möglich?. (German) [Addendum to [414] (present and future quantum electrodynamics). Can quantum mechanics be completed deterministically?]. In Hermann et al. [HvMW79], pages 409–418. ISBN 0-387-08962-4, 3-540-08962-

4. LCCN QC16.P37 W64. URL http://www.springer.com/
physics/book/978-3-540-08962-9. Addendum to letter of

Heisenberg to Pauli of 25 April 1935.

Heisenberg:1979:AHBa

[Hei79b] Werner Heisenberg. Anlage zu Heisenbergs Briefe [407] (Bisherige und zukunftige Quantenelektrodynamik). (German) [Addendum to Heisenberg’s letter [407] (on present and future quantum electrodynamics)]. In Hermann et al. [HvMW79], pages 386–392. ISBN 0-387-08962-4, 3-540-08962-4. LCCN QC16.P37

W64. URL http://www.springer.com/physics/book/978-3-

540-08962-9. Addendum to letter of Heisenberg to Pauli of 25 April 1935.

Heisenberg:1984:BAE

[Hei84] Werner Heisenberg. Bericht über die allgemeinen Eigenschaften der Elementarteilchen. II. Die Wechselwirkung der Elementarteilchen. III. Die Grenzen der bisherighen Theorie. (German) [Report on the general properties of elementary particles. II. The interaction of elementary particles. III. The limits of current theory]. In Blum et al. [BDR84], pages 346–358. ISBN 3-540-13020-9, 0-387-13020-9. LCCN QC3 .H33 1984. Parts II and III were to be presented at the 8th Solvay Conference in Brussels, 22–29 October 1939. The conference was canceled because of the outbreak of World War II in September 1939, and the report was not published until the Collected Works were prepared. Wolfgang Pauli prepared Part I of this joint work; it was published in revised form in [Pau41].

Helman:1967:ESF

[Hel67] J. S. Helman. Effect of spin-flip processes on the Pauli susceptibility. Zeitschrift für Physik. B, Condensed matter, 6(4): 297–306, October 27, 1967. CODEN ZPCMDN. ISSN 0722-3277 (print), 1431-584X (electronic). URL http://link.springer.

com/article/10.1007/BF02422510;   http://www.springerlink.
com/content/47124t115p063056/fulltext.pdf.

Hendry:1981:PP

[Hen81a] John Hendry. Pauli as philosopher. British Journal for the Philosophy of Science, 32(3):277–282, September 1981. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL http://bjps.oxfordjournals.org/content/32/3/277.
full.pdf+html;   http://www.jstor.org/stable/687591.

Henneberger:1981:ABS

[Hen81b] Walter C. Henneberger. Aharonov–Bohm scattering and the velocity operator. Journal of Mathematical Physics, 22(1):116– 117, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i1/p116_s1.

Hendry:1984:CQM

[Hen84] John Hendry. The creation of quantum mechanics and the Bohr– Pauli dialogue, volume 14 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 90-277-1648-X. xi + 177 pp. LCCN QC173.98 .H46 1984.

Hendry:1986:BRW

[Hen86] John Hendry. Book review: Wolfgang Pauli. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Scientific Correspondence with Bohr, Einstein, Heisenberg a.o. Berlin, Heidelberg and New York: Springer, 1985. Pp. xxix + 783. ISBN 3-540-13609-6. DM 298.00. British Journal for the History of Science, 19(3):348, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4026519.

Hibbert:1991:GPCb

[HGF91] Alan Hibbert, Robert Glass, and Charlotte Froese Fischer. A general program for computing angular integrals of the Breit–Pauli Hamiltonian. Computer Physics Communications, 64(3):455–472, June 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190138B.

Hrivnak:2006:GCP

[HNPT06] J. Hrivnák, P. Novotný, J. Patera, and J. Tolar. Graded contractions of the Pauli graded sl(3, C). Linear Algebra and its Applications, 418(2–3):498–550, October 15, 2006. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Horn:2011:GPD

[Hor11] Martin Erik Horn. Grassmann, Pauli, Dirac: special relativity in the schoolroom. In Petsche [Pet11], pages 435–452. ISBN 3-0346-0404-1 (print), 3-0346-0405-X (e-book). LCCN

QA29.G73 F76 2011. URL http://www.springerlink.com/
content/978-3-0346-0405-5.

Heisenberg:1929:QWG

[HP29] W. Heisenberg and W. Pauli. Zur Quantendynamik der Wellenfelder. (German) [On the quantum dynamics of wave fields]. Zeitschrift für Physik, 56(1–2):1–61, January 1929. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01340129;
 http://www.springerlink.com/content/mr542891u38632q0/

.

Heisenberg:1930:QWI

[HP30] W. Heisenberg and W. Pauli. Zur Quantentheorie der Wellenfelder. II. (German) [On the quantum dynamics of wave fields.

II]. Zeitschrift für Physik, 59(3–4):168–190, March 1930. CO-DEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01341423;   http://www.springerlink.com/content/m516326152712757/

.

Huang:1970:RPP

[HP70] Kerson Huang and Frank E. Paige. Regge poles and the Pauli Principle. Physical Review D (Particles and Fields), 2(7):1351– 1352, October 1, 1970. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.2.1351.

Heisenberg:1993:IGT

[HP93] Werner Heisenberg and Wolfgang Pauli. On the isospin group in the theory of the elementary particles. In Blum et al. [BDR93], pages 337–351. ISBN 3-540-13848-X. Preprint of March 1958. See disagreement [Pau93] in which Pauli broke off collaborative work with Heisenberg in this area. Pauli died on 15 December 1958.

Havlicek:2002:AFG

[HPPT02] Miloslav Havlíček, Jiří Patera, Edita Pelantová, and Jiří Tolar. Automorphisms of the fine grading of sl(n, C) associated with the generalized Pauli matrices. Journal of Mathematical Physics, 43(2):1083–1094, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hristev:1971:CCC

[Hri71] Anatole P. Hristev. Conformal charge conjugation. Journal of Mathematical Physics, 12(1):118–124, January 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i1/p118_s1.

Hiroshima:2011:NBR

[HSS11] Fumio Hiroshima, Herbert Spohn, and Akito Suzuki. The nobinding regime of the Pauli–Fierz model. Journal of Mathematical Physics, 52(6):062104, June 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/
i6/p062104_s1.

Hughes:1989:SIQ

[Hug89] R. I. G. Hughes. The Structure and Interpretation of Quantum Mechanics. Harvard University Press, Cambridge, MA, USA, 1989. ISBN 0-674-84391-6 (hardcover), 0-674-84392-4 (paperback). xiii + 369 pp. LCCN QC174.12 .H82 1989.

Hurter:2022:TBS

[Hür22] Tobias Hürter. Too Big for a Single Mind: How the Greatest Generation of Physicists Uncovered the Quantum World. The Experiment, New York, NY, USA, 2022. ISBN 1-61519-920-9 (hardcover), 1-61519-921-7 (e-book). 357 pp. LCCN QC7 .H88813 2022. Translation to English by David Shaw of Das Zeitalter der Unschärfe, published by Klett-Cotta (2021).

Hermann:1979:WPW

[HvMW79] Armin Hermann, Karl von Meyenn, and Victor F. Weisskopf, editors. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band I: 1919–1929 Scientific Correspondence With Bohr, Einstein, Heisenberg, a.o. Volume I: 1919–1929, volume 2 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. ISBN 0-387-08962-4, 3-540-08962-4. xlvii + 577 pp. LCCN QC16.P37

W64. URL http://www.springer.com/physics/book/978-3-
540-08962-9.

Heinzmann:1987:BRS

[HW87] B. Heinzmann and S. L. Wolff. Book review: Sources in the History of Mathematics and Physical Sciences Volume 6: Teil II: 1930–1939 by G. J. Toomer, Wolfgang Pauli, Karl von Meyenn, Armin Hermann, Victor F. Weisskopf. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 71(1):120, ???? 1987. CODEN SUARAH. ISSN 0039-4564. URL http://www.jstor.

org/stable/20777123.

Inahama:2003:SPP

[IiS03] Yuzuru Inahama and Shin ichi Shirai. Spectral properties of Pauli operators on the Poincaré upper-half plane. Journal of Mathematical Physics, 44(6):2451–2462, June 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Javorsek:2000:NET

[JBE^{+}00] D. Javorsek II, M. Bourgeois, D. Elmore, E. Fischbach, D. Hillegonds, J. Marder, T. Miller, H. Rohrs, M. Stohler, and S. Vogt. New experimental test of the Pauli Exclusion Principle using accelerator mass spectrometry. Physical Review Letters, 85(13): 2701–2704, September 25, 2000. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://

link.aps.org/doi/10.1103/PhysRevLett.85.2701.

Jensen:2000:CCN

[Jen00] Carsten Jensen. Controversy and Consensus: Nuclear Beta Decay 1911–1934, volume 24 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2000. ISBN 3-0348-9569-0 (paperback), 3-7643-5313-9 (hardcover), 3-0348-8444-3 (e-book). xv + 217 pp. LCCN

QC793.5.B425 J46 2000. URL http://www.springerlink.
com/content/978-3-0348-8444-0. Carsten Jensen died of can-

cer a few months after presenting his doctoral dissertation in 1990 at the University of Copenhagen. Finn Aaserud, Helge Kragh, Erik Rüdinger, and Roger H. Stuewer produced this book as a slightly edited version of that work, supplying additional figures, but leaving the prose largely untouched.

Jackson:2004:VFW

[JG04] J. David Jackson and Kurt Gottfried. Victor Frederick Weisskopf: September 19, 1908–April 22, 2002. Biographical memoirs — National Academy of Sciences of the United States of America, 84:372–398, 2004. CODEN BMN-SAC. ISBN 0-309-52419-9. ISSN 0077-2933. URL http:

//www.nap.edu/catalog.php?record_id=10992;   http://
www.nap.edu/catalog/10992.html.

Jolos:1980:EPP

[JMS80] R. V. Jolos, J. L. Molina, and V. G. Soloviev. Effect of the Pauli principle on the excited states of doubly-even deformed nuclei. Zeitschrift für Physik, 295(2):147–152, June 1980. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01413072;
 http://www.springerlink.com/content/m38876v87q741xk1/
fulltext.pdf.

Johnson:1933:VMP

[Joh33] M. H. Johnson, Jr. The vector model and the Pauli Principle. Physical Review (2), 43(8):627–631, April 15, 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.43.627.

Jordan:1928:QLF

[JP28] P. Jordan and W. Pauli, Jr. Zur Quantenelektrodynamik ladungsfreier Felder. (German) [On quantum electrodynamics of charge-free fields]. Zeitschrift für Physik, 47(3–4):151–173, February 1928. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic). URL http://link.springer.com/
article/10.1007/BF02055793;   http://www.springerlink.
com/content/p2th5r124576672k/fulltext.pdf.

Jung:1952:NPS

[JP52] C. G. (Carl Gustav) Jung and Wolfgang Pauli, editors. Naturerklärung und Psyche Synchronizität als ein Prinzip akausaler Zusammenhänge [von C. G. Jung]. Der Einfluss archetypischer Vorsstellungen auf die Bildung naturwissenschaftlicher Theorien bei Kepler [von W. Pauli]: Synchronizität als ein Prinzip akausaler Zusammenhänge, volume 4 of Studien aus dem C. G. Jung-Institut, Zürich. Rascher, Zürich, Switzerland, 1952. 194 pp. LCCN BF1033 .N38. Available in English translation [JP55].

Jung:1955:INP

[JP55] C. G. (Carl Gustav) Jung and Wolfgang Pauli. The Interpretation of nature and the psyche, volume 51 of Bollingen series. Pantheon Books, New York, NY, USA, 1955. vii + 247 pp. LCCN BF1033 .N385. English translation by R. F. C. Hull of [JP52].

Jordan:1928:PAG

[JW28] P. Jordan and E. Wigner. Über das Paulische Äquivalenzverbot. (German) [On Pauli’s equivalence prohibition]. Zeitschrift für Physik, 47(9–10):631–651, ???? 1928. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01331938;   http://www.
springerlink.com/content/hx1t32272451437h.

Kacser:1974:PP

[Kac74] Claude Kacser. Pauli’s proof. Physics Today, 27(8):70, August 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v27/
i8/p70/s1.

Kaempffert:1948:RRB

[Kae48] Waldemar Kaempffert. The revolution that radium began: Fifty years after the Curies’ great discovery, nuclear physics is still a realm unbounded. New York Times, ??(??):SM13, SM25, SM27, December 26, 1948. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/108348269/.

Kalbach:1989:PPC

[Kal89] C. Kalbach. Pauli principle corrections in particle-hole state densities. Zeitschrift für Physik, 332(2):157–161, June 1989. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01289771;   http://www.springerlink.com/content/q13071475055x37n/
fulltext.pdf.

Kaplan:2002:PEP

[Kap02] I. G. Kaplan. Is the Pauli exclusive principle an independent quantum mechanical postulate? International Journal of Quantum Chemistry, 89(4):268–276, ???? 2002. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaplan:2013:PEP

[Kap13] I. G. Kaplan. The Pauli Exclusion Principle. Can it be proved? Foundations of Physics, 43(10):1233–1251, October 2013. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/s10701-
013-9742-4.

Kemmer:1959:PWP

[Kem59] N. Kemmer. Prof. Wolfgang E. Pauli, For.Mem.R.S. Nature, 183 (4668):1089, April 18, 1959. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v183/n4668/pdf/1831089a0.pdf.

Kemmer:1994:RS

[Kem94] Nicholas Kemmer. Robert Schlapp (1899–1991). In Wolfgang Pauli: Writings on physics and philosophy [EvM94], pages 7– 12. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York).

LCCN QC6.2 .P38 1994. URL http://www.loc.gov/catdir/
enhancements/fy0812/94015098-d.htm.

Kennard:1930:RPE

[Ken30] E. H. Kennard. On the reason for Pauli’s Exclusion Principle. Physical Review (2), 35(9):1127, May 1, 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.35.1127.

Keve:2000:TPA

[Kev00] Tom Keve. Triad: the physicists, the analysts, the kabbalists. Rosenberger and Krausz, London, UK, 2000. ISBN 0-9536219-0-1. ix + 362 pp. LCCN ????

Kuhn:1967:SHQ

[KHFA67] Thomas S. Kuhn, John L. Heilbron, Paul Forman, and Lini Allen. Sources for history of quantum physics: an inventory and report, volume 68 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1967. ix + 176 pp. LCCN QC174.1 .S66.

URL http://www.amphilsoc.org/guides/ahqp/;   http://
www.amphilsoc.org/guides/ahqp/s-t.htm#schrodinger.

Kijewski:1972:TMC

[Kij72] L. J. Kijewski. Test of the g-matrix condition on the twomatrix satisfying the Pauli restriction. International Journal of Quantum Chemistry, 6(S6):73–78, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Krejs:1973:TGS

[KK73] Franz Krejs and Abraham Klein. Theory of ground state correlations of closed shell nuclei: a density-matrix formulation. Journal of Mathematical Physics, 14(9):1155–1163, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v14/i9/p1155_s1.

Klein:1959:WPN

[Kle59] Oscar Klein. Wolfgang Pauli. Några minnesord. (Swedish) [Wolfgang Pauli. some reminiscences]. Kosmos, Fysika Uppsatzer, 37(??):9–12, ???? 1959. German translation in [Cha68, 49–52].

Klein:2005:ESF

[Kle05] Etienne Klein. Il était sept fois la révolution: Albert Einstein et les autres . . .. (French) [The seven-times revolution: Albert Einstein and the others …]. Flammarion, Paris, France, 2005. ISBN 2-08-210343-9. 237 pp. LCCN QC16.E5.

Kekez:1990:ULV

[KLL90] D. Kekez, A. Ljubiić, and B. A. Logan. An upper limit to violations of the Pauli Exclusion Principle. Nature, 348(6298): 224, November 15, 1990. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v348/n6298/pdf/348224a0.pdf.

Kozlowski:1989:ADE

[KM89] P. M. Kozlowski and N. H. March. Approximate density– external potential relation and the Pauli potential for systems with Coulombic interaction. International Journal of Quantum Chemistry, 36(6):741–748, December 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ktorides:1980:ERP

[KMS80] Christos N. Ktorides, Hyo Chul Myung, and Ruggero Maria Santilli. Elaboration of the recently proposed test of Pauli’s Principle under strong interactions. Physical Review D (Particles and Fields), 22(4):892–907, August 15, 1980. CODEN PRV-DAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.22.892.

Kramers:1923:TBG

[KP23] H. A. Kramers and W. Pauli. Zur Theorie der Bandenspektren. (German) [On the theory of band spectra]. Zeitschrift für Physik, 13(1):351–367, December 1923. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01328226;

 http://www.springerlink.com/content/j621635506g4586t/
fulltext.pdf.

Kallen:1955:MSD

[KP55] Gunnar Källén and Wolfgang Pauli. On the mathematical structure of T. D. Lee’s model of a renormalizable field theory: dedicated to Professor Niels Bohr on the occasion of his 70th birthday, volume 30(7) of Det Kongelige Danske videnskabernes selskab. Matematisk-fysiske meddelelser. Munksgaard, København, Danmark, 1955. 23 pp. LCCN AS281 .D215 bd. 30, no. 7.

Kramer:1965:AMP

[Kra65] Peter Kramer. Angular momenta and Pauli principle in nuclear reactions a+A → C^{∗} → b_{1}+b_{2}+b_{3}. Reviews of Modern Physics, 37(3):346–349, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.37.346;   http://rmp.
aps.org/abstract/RMP/v37/i3/p346_1.

Kragh:2009:BRM

[Kra09] Helge Kragh. Book review: Michela Massimi, Pauli’s Exclusion Principle: The Origin and Validation of a Scientific Principle. British Journal for the Philosophy of Science, 60(1):235–238, March 2009. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/60/1/235.full.pdf+html;
http://www.jstor.org/stable/25591996.

Kroner:1954:WPE

[Krö54] Franz Kröner. W. Pauli: Der Einfluss archetypischer Vorstellungen auf die Bildung naturwissenschaftlicher Theorien bei Kepler. (German) [W. Pauli: The influence of archetypal ideas on the formation of scientific theories of Kepler]. Dialectica: International Review of Philosophy of Knowledge, 8(2):173–180, June 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kircher:1981:PCS

[KS81] R. Kircher and E. W. Schmid. Pauli-correct separable potentials for composite particle interactions. Zeitschrift für Physik, 299(3):241–244, September 1981. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01443941;

 http://www.springerlink.com/content/qr2g2532874823p1/
fulltext.pdf.

Kronig:1964:CSP

[KW64] Ralph Kronig and Victor F. Weisskopf, editors. Collected Scientific Papers [Wolfgang Pauli]. Interscience Publishers, New York, NY, USA and London, UK, 1964. ???? pp. LCCN QC3 .P32 1964. Two volumes.

Lakhtakia:1996:MMH

[Lak96] A. (Akhlesh) Lakhtakia, editor. Models and Modelers of Hydrogen: Thales, Thomson, Rutherford, Bohr, Sommerfeld, Goudsmit, Heisenberg, Schrödinger, Dirac, Sallhofer. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1996. ISBN 981-02-2302-1. xv + 424 pp. LCCN

QC171.2 .M63 1996. URL http://www.worldscientific.com/
worldscibooks/10.1142/2788.

Landau:1959:WPOb

[Lan59a] L. D. Landau. Wolfgang Pauli [obituary]. Soviet Physics. Uspekhi, 2(4):624, 1959. CODEN SOPUAP. ISSN 0038-5670. URL

http://stacks.iop.org/0038-5670/2/i=4/a=M09.

Landau:1959:WPOa

[Lan59b] Lev D. Landau. Wolfgang Pauli [obituary]. Uspekhi Fizicheskikh Nauk, 68(7):556–559, July 1959. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.

ru/ru/articles/1959/7/j/.

Laurikainen:1980:WPJ

[Lau80] Kalervo Vihtori Laurikainen. Wolfgang Pauli ja hänen kirjeensä. (finnish) [Wolfgang Pauli and his letters]. Arkhimedes, 32(??): ??, ???? 1980. CODEN AKMDA5. ISSN 0004-1920.

Laurikainen:1982:WPP

[Lau82] Kalervo Vihtori Laurikainen. Wolfgang Pauli ja filosofia. (Finnish) [Wolfgang Pauli and philosophy]. Arkhimedes, 34(4): 206–222, ???? 1982. CODEN AKMDA5. ISSN 0004-1920.

Laurikainen:1985:WPCb

[Lau85a] Kalervo Vihtori Laurikainen. Wolfgang Pauli and the Copenhagen philosophy. In Lahti and Mittelstaedt [LM85], pages 273– 287. ISBN 9971-5-0004-3. LCCN QC173.96 .S96 1985.

Laurikainen:1985:WPCa

[Lau85b] Kalervo Vihtori Laurikainen. Wolfgang Pauli’s conception of reality. In Lahti and Mittelstaedt [LM85], pages 209–228. ISBN 9971-5-0004-3. LCCN QC173.96 .S96 1985.

Laurikainen:1988:BAP

[Lau88] Kalervo Vihtori Laurikainen. Beyond the atom: the philosophical thought of Wolfgang Pauli. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-19456-8 (US). xix + 234 pp. LCCN QC16.P37 L3813 1988.

Laurikainen:1997:MAE

[Lau97] Kalervo Vihtori Laurikainen. The message of the atoms: essays on Wolfgang Pauli and the unspeakable. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-61754-X, 3-642-64457-0, 3-642-60560-5 (e-book). ix + 203 pp. LCCN QC174.12 .L39 1997.

Lenard:1968:SMI

[LD68] A. Lenard and Freeman J. Dyson. Stability of matter. II. Journal of Mathematical Physics, 9(5):698–711, May 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i5/p698_s1.

Lindorff:2004:PJM

[Lin04] David P. Lindorff. Pauli and Jung: the meeting of two great minds. Quest Books, Wheaton, IL, USA, 2004. ISBN 0-8356-0837-9. xiv + 299 pp. LCCN BF109.J8 L55 2004.

Logan:1979:VPE

[LL79] B. A. Logan and A. Ljubičić. Validity of the Pauli Exclusion Principle for nucleons. Physical Review C (Nuclear Physics), 20(5):1957–1958, November 1, 1979. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490x, 1538-4497. URL http://link.

aps.org/doi/10.1103/PhysRevC.20.1957.

Lewis:1974:PL

[LM74] H. R. Lewis and S. Margulies. Pauli lectures. Physics Today, 27(7):13, July 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v27/i7/p13/s1.

Lahti:1985:SFM

[LM85] Pekka (Pekka Johannes) Lahti and Peter Mittelstaedt, editors. Symposium on the Foundations of Modern Physics: 50 Years of the Einstein–Podolsky–Rosen Gedankenexperiment, Joensuu, Finland, 16–20 June 1985. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1985. ISBN 9971-5-0004-3. LCCN QC173.96 .S96 1985.

Leon:2017:PO

[LM17] Juan Leon and Lorenzo Maccone. The Pauli objection. Foundations of Physics, 47(12):1597–1608, December 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Lopuszanski:1991:APL

[Łop91] Jan Łopuszański. On the action of the Pauli–Lubanski Casimir operator in a relativistic supersymmetric quantum field theory. Journal of Mathematical Physics, 32(11):3189–3194, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v32/i11/p3189_s1.

Lauber:1956:BGB

[LPB^{+}56] A. Lauber, M. Plancherel, W. Baumgartner, W. Pauli, H. Labhart, et al. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 7(5):465–472, September 1956. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039

(electronic). URL http://link.springer.com/article/10.
1007/BF01606331;   http://www.springerlink.com/content/
q104578154215007/fulltext.pdf.

Lieb:1991:SMA

[LT91] Elliott H. Lieb and Walter E. Thirring, editors. The stability of matter: from atoms to stars: selecta of Elliott H. Lieb. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-540-53039-8 (Berlin), 0-387-53039-8 (N.Y.). viii + 565 pp. LCCN QC173.4.T48 L54 1991. With a preface by F. Dyson.

Lieb:1997:SMA

[LT97] Elliott H. Lieb and Walter E. Thirring, editors. The stability of matter: from atoms to stars: selecta of Elliott H. Lieb. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1997. ISBN 3-540-61565-2 (Berlin: hardcover). xi + 675 pp. LCCN QC173.4.T48 L54 1997. With a preface by F. Dyson.

Lieb:2001:SMA

[LT01] Elliott H. Lieb and Walter E. Thirring, editors. The stability of matter: from atoms to stars: selecta of Elliott H. Lieb. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2001. ISBN 3-540-42083-5. xiii + 812 pp. LCCN QC173.4.T48 L54 2001. With a preface by F. Dyson.

Lieb:2005:SMA

[LT05] Elliott H. Lieb and Walter E. Thirring, editors. The stability of matter: from atoms to stars: selecta of Elliott H. Lieb. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2005. ISBN 3-540-22212-X. xv + 932 pp. LCCN QC173.4.T48 L54 2005. URL http://www.loc.gov/

catdir/enhancements/fy0663/2004108033-d.html. With a

preface by F. Dyson.

Luders:1958:CBS

[LZ58] Gerhart Lüders and Bruno Zumino. Connection between spin and statistics. Physical Review, 110(6):1450–1453, June 1958. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v110/i6/
p1450_1.

Majorana:1937:TSD

[Maj37] Ettore Majorana. Teoria simmetrica dell elettrone e del positrone. (Italian) [Symmetrical theory of the electron and the positron]. Il Nuovo Cimento (8), 14(4):171–184, April 1937. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://en.wikipedia.org/

wiki/Majorana_mass;   http://nl.wikipedia.org/wiki/

Majorana-deeltje. In this paper, Majorana predicted the existence of a new type of particle, now called a Majorana fermion, which is its own antiparticle, and whose existence may have finally been confirmed by experiment seventy years later [?, ?, ?]. See also comments in [?]. Esposito [?] reports about this paper “With amazing farsightedness Majorana suggested that the neutrino, which had just been postulated by Wolfgang Pauli and Fermi to explain puzzling features of radioactive beta decay, could be such a particle. This would make the neutrino unique among the elementary particles and, moreover, enable it to have mass. Today many experiments are still devoted to detect these peculiar properties, which include the phenomenon of neutrino oscillations: we have not yet succeeded to find a definite answer to Majorana’s proposal.”.

Margenau:1944:EPP

[Mar44] Henry Margenau. The exclusion principle and its philosophical importance. Philosophy of Science, 11(4):187–208, October 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184797.

Massimi:2004:WDI

[Mas04] Michela Massimi. What demonstrative induction can do against the threat of underdetermination: Bohr, Heisenberg, and Pauli on spectroscopic anomalies (1921–24). Synthese, 140(3):243– 277, June 2004. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/
content/pdf/10.1023/B%3ASYNT.0000031319.64615.49.pdf.

Massimi:2005:PEP

[Mas05] Michela Massimi. Pauli’s exclusion principle: the origin and validation of a scientific principle. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-83911-4 (hardcover). xiv + 211 pp. LCCN QC174.17.P3 M37 2005. URL http://www.loc.

gov/catdir/enhancements/fy0632/2005296620-d.html;
http://www.loc.gov/catdir/enhancements/fy0632/2005296620-
t.html;   http://www.loc.gov/catdir/enhancements/fy0733/
2005296620-b.html.

McClintock:1986:LTP

[McC86] Peter McClintock. Low-temperature physics: Pauli Principle in a gas. Nature, 323(6091):756–757, October 30, 1986. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v323/n6091/
pdf/323756b0.pdf.

Meier:2001:AAP

[MEF01] C. A. (Carl Alfred) Meier, Charles P. (Charles Paul) Enz, and Markus Fierz, editors. Atom and archetype: the Pauli/ Jung letters, 1932–1958. Princeton University Press, Princeton, NJ, USA, 2001. ISBN 0-691-01207-5. lx + 250 pp. LCCN

QC16.P37 A42 2001. URL http://www.loc.gov/catdir/
bios/prin051/2001016323.html;   http://www.loc.gov/
catdir/description/prin021/2001016323.html;   http:/
/www.loc.gov/catdir/samples/prin031/2001016323.html;
 http://www.loc.gov/catdir/toc/prin031/2001016323.html.

Mehra:1973:PCN

[Meh73] Jagdish Mehra, editor. The physicist’s conception of nature: Symposium on the Development of the Physicist’s Conception of Nature in the 20th century. Held at the International Centre for Theoretical Physics, Miramare, Trieste, Italy, 18–25 September 1972. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1973. ISBN 90-277-0345-0, 90-277-2536-5. LCCN QC173.96 .S95 1972. URL http://www.

springer.com/us/book/9789027703453.

Mehra:1975:SCP

[Meh75] Jagdish Mehra, editor. The Solvay conferences on physics: aspects of the development of physics since 1911. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1975. ISBN 90-277-0635-2. xxxii + 415 pp. LCCN QC1.S792 M43.

Meier:1992:WPC

[Mei92] Carl Alfred Meier. Wolfgang Pauli and Carl Gustav Jung: Ein Briefwechsel: 1932–1958. (German) [Wolfgang Pauli and Carl Gustav Jung: Correspondence: 1932–1958]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-54663-4. 275 pp. LCCN QC16.P37. URL

http://www.gbv.de/dms/ilmenau/toc/022349073.PDF.

Meyer:2004:RWP

[Mey04] Horst Meyer. Reminiscence of Wolfgang Pauli. Physics Today, 57(6):16, June 2004. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v57/i6/p16/s1.

Martienssen:1995:BAF

[MGG^{+}95] W. Martienssen, S. Großmann, P. Grassberger, D. Meissner, W. Sandhas, H. Nicolai, D. Hoffmann, H. Kant, P. Richter, H. Atmanspacher, M. Eckert, H. Rechenberg, H. Maier-Leibnitz, and D. Haarer. Buchbesprechungen: Argyris/Faust: Die Erforschung des Chaos/Plaschko/Brod: Nichtlineare Dynamik, Bifurkation und Chaotische Systeme/Bunde/Havlin: Fractals in Science/Goetzberger/Voß/Knobloch: Sonnenenergie: Photovoltaik/Lindner: Grundkurs Theoretische Physik/Vilenkin/Shellard: Cosmic Strings and other Topological Defects/Weiss: Großforschung in Deutschland Geschichte des Hahn-Meitner-Instituts/Fölsing: Wilhelm Conrad Röntgen Aufbruch ins Innere der Materie/Brandmüller: Galilei und die Kirche Ein Fall und seine Lösung/Meÿenn: Wolfgang Pauli Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a. Band 3: 1940 49/Atmanspacher/Primas: Der Pauli-Jung-Dialog und seine Bedeutung für die moderne Wissenschaft/Cassidy: Werner Heisenberg Leben und Werk/Weinberg: The First Nuclear Era The Life and Times of a Technological Fixer/Wenske: Wörterbuch Chemie/Dictionary of Chemistry. Physikalische Blätter, 51(11):1101–1106, November 1995. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
19950511117/abstract.

Mielke:1981:TES

[Mie81] Eckehard W. Mielke. Toward exact solutions of the nonlinear Heisenberg–Pauli–Weyl spinor equation. Journal of Mathematical Physics, 22(9):2034–2039, September 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/
i9/p2034_s1.

Milotti:2007:EFV

[Mil07] Edoardo Milotti. Enrico Fermi’s view of identical particles. arxiv.org, ??(??):9, May 9, 2007. URL http://arxiv.org/abs/ 0705.1363. See [Fer24, Fer26] for the original Italian and German versions.

Miller:2009:DCN

[Mil09] Arthur I. Miller. Deciphering the Cosmic Number: the Strange Friendship of Wolfgang Pauli and Carl Jung. W. W. Norton & Co., New York, NY, USA, 2009. ISBN 0-393-06532-4 (hardcover). xxv + 336 + 8 pp. LCCN QC16.P37 M55 2009.

Miller:2010:JPP

[Mil10] Arthur I. Miller. 137: Jung, Pauli, and the Pursuit of a Scientific Obsession. W. W. Norton & Co., New York, NY, USA, 2010. ISBN 0-393-33864-9. xxiii + 336 pp. LCCN QC16.P37 M55 2010.

Methfessel:1991:RMC

[MKD^{+}91] S. Methfessel, H. Klar, D. Dubbers, W. Sandhas, K. Kromphardt, H. F. Ebel, E. Filos, H. Rechenberg, Th. Görnitz, and K. Stierstadt. [Rezensionen:] Mott: Can Scientists Believe?/Friedrich: Theoretische Atomphysik/Ignatovich: The Physics of Ultracold Neutrons/Belyaev: Lectures on the Theory of Few-Body Systems/Elster: Humanökologie als Aufgabe für Natur- u. Geisteswissenschaften/Vischer: Plane deinen Ruhm!/Fritzsch, Sucher: Von und über Physiker/Kay: Managing Creativity in Science and Hi-Tech/Scribal: Max Planck: Wissenschaftliche Selbstbiographie/Laurikainen: Beyond the Atom: The Philosophical Thought of W. Pauli/Falk: Physik Zahl und Realität. Physikalische Blätter, 47(7):650–653, 1991. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Miller:2011:CGJ

[MM11] Arthur I. Miller and Hubert Mania. 137: C. G. Jung, Wolfgang Pauli und die Suche nach der kosmischen Zahl. (German) [137: Jung, Pauli, and the search for the cosmic number]. Dt. Verl.- Anst., München, Germany, 2011. ISBN 3-421-04290-X. 411 pp. LCCN ????

Morrison:1956:BRB

[Mor56a] P. Morrison. Book review: Niels Bohr and the Development of Physics: Essays Dedicated to Niels Bohr on the Occasion of his Seventieth Birthday, edited by W. Pauli, L. Rosenfeld, and V. Weisskopf. Physics Today, 9(8):32, August 1956. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://link.aip.org/link/phtoad/v9/i8/p32/s1. See

[PRW55].

Morrison:1956:N

[Mor56b]  Philip Morrison.  The neutrino.  Nature, 194(??):58–

69, ???? 1956. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

scientificamerican/journal/v194/n1/pdf/scientificamerican0156-
58.pdf.

Morrison:2007:SAW

[Mor07] Margaret Morrison. Spin: All is not what it seems. Studies in History and Philosophy of Modern Physics, 38(3):529– 557, September 2007. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219806000931.

Mourad:1994:TFR

[MS94] J. Mourad and H. Sazdjian. The two-fermion relativistic wave equations of constraint theory in the Pauli–Schrödinger form. Journal of Mathematical Physics, 35(12):6379–6406, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v35/i12/p6379_s1.

Nagy:2010:PPD

[Nag10] Á. Nagy. The Pauli potential from the differential virial theorem. International Journal of Quantum Chemistry, 110(12): 2117–2120, ???? 2010. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

NRCNFAC:2003:NBN

[Nat03] National Research Council (U.S.). Neutrino Facilities Assessment Committee. Neutrinos and beyond: new windows on nature. National Academy Press, Washington, DC, USA, 2003. ISBN 0-309-08716-3 (Book), 0-309-50634-4 (PDF). xiv + 89 pp. LCCN QB464.2 .N36 2003. URL http://www.nap.edu/

catalog.php?record_id=10583.

Nauenberg:2013:ESB

[Nau13] Michael Nauenberg. Edmund Stoner and the Bohr atom. Physics Today, 66(4):10, April 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Navarro:2006:EAD

[Nav06] Jaume Navarro. Early attempts to detect the neutrino at the Cavendish Laboratory. Physics in Perspective (PIP), 8 (1):64–82, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-005-0249-z.

Nelson:1972:ASK

[Nel72] C. A. Nelson. Algebraic solution for the Källén–Pauli state vectors in the V θ sector by a congruence transformation. Journal of Mathematical Physics, 13(7):1051–1056, July 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i7/p1051_s1.

Nolte:1991:TPE

[NFG^{+}91] E. Nolte, T. Faestermann, A. Gillitzer, G. Korschinek, D. Müller, V. M. Novikov, A. A. Pomansky, A. Ljubicic, and D. Miljanic. Test of the Pauli exclusion principle for nucleons and atomic electrons by accelerator mass spectrometry. Zeitschrift für Physik, 340(4):411–413, December 1991. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01290329;
 http://www.springerlink.com/content/h81121p1197q1734/
fulltext.pdf.

Nagaoka:1926:U

[NM26] H. Nagaoka and T. Mishima. [unknown]. Proc. Imp. Acad. Tokyo, 2(??):249–??, ???? 1926.

Nagy:1991:EFP

[NM91] A. Nagy and N. H. March. The exact form of the Pauli potential for the ground state of two- and three-level atoms and ions. International Journal of Quantum Chemistry, 39(4):615– 623, April 1991. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nakaichi-Maeda:1984:TBP

[NMS84] S. Nakaichi-Maeda and E. W. Schmid. Three-body Pauli potential in the npα system. Zeitschrift für Physik, 318(2):171–177, June 1984. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/
10.1007/BF01413465;   http://www.springerlink.com/content/
ht374n2451w70877/fulltext.pdf.

Niederle:1999:ESS

[NN99] J. Niederle and A. G. Nikitin. Extended supersymmetries for the Schrödinger–Pauli equation. Journal of Mathematical Physics, 40(3):1280–1293, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nouicer:2006:PHP

[Nou06] Khireddine Nouicer. Pauli–Hamiltonian in the presence of minimal lengths. Journal of Mathematical Physics, 47(12):122102, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i12/p122102_s1.

Oppenheimer:19xx:JRO

[OBxx] J. Robert Oppenheimer and Niels Bohr. J. Robert Oppenheimer papers (1799–1980) (bulk 1947–1967) and Niels Henrik David Bohr papers. 19xx. ???? pp. LCCN 0456G; Vault 0202A; Micro-

film 16,646-1P. URL http://hdl.loc.gov/loc.mss/eadmss.
ms998007.

Orear:2001:EFM

[Ore01] Jay Orear. Enrico Fermi, the man. Il Nuovo Saggiatore, 17(5–6): 30–38, September/December 2, 2001. ISSN 0393-4578 (print),

1827-6148 (electronic). URL http://prometeo.sif.it:8080/
papers/online/sag/2001/05-06/pdf/05.pdf. Special issue in

honor of the centennial of the birth of Enrico Fermi.

ORaifeartaigh:2000:GTH

[OS00] Lochlainn O’Raifeartaigh and Norbert Straumann. Gauge theory: Historical origins and some modern developments. Reviews of Modern Physics, 72(1):1–23, January 2000. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
72.1;   http://rmp.aps.org/abstract/RMP/v72/i1/p1_1.

Oechsner:1986:BKP

[OSWR86] H. Oechsner, A. Schlachetzki, H. Walther, and H. Rechenberg. Buchbesprechungen: Kirschner: Polarized Electrons at Surfaces/Paul, Teubner: Optoelektronische Halbleiterbauelemente/Stitch, Bass: Laser Handbook, Vol. 4/Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a.: Teil II: 1930–1939. Physikalische Blätter, 42(7):252–253, July 1986. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-

3722 (electronic). URL http://onlinelibrary.wiley.com/
doi/10.1002/phbl.19860420726/abstract.

Pagnamenta:1965:SKP

[Pag65] Antonio Pagnamenta. Solution of the Källén–Pauli equation. Journal of Mathematical Physics, 6(6):955–965, June 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i6/p955_s1.

Pais:2000:WEP

[Pai00a] A. Pais. Wolfgang Ernst Pauli. In The Genius of Science: a Portrait Gallery [Pai00b], pages 210–262. ISBN 0-19-850614-7 (hardcover). LCCN Q141 .P29 2000. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/;   http://www.loc.
gov/catdir/toc/fy02/99046603.html.

Pais:2000:GSP

[Pai00b] Abraham Pais. The Genius of Science: a Portrait Gallery. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850614-7 (hardcover). 356 pp.

LCCN Q141 .P29 2000.  URL ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/toc/fy02/99046603.html.

Pauli:1919:MSW

[Pau19a] W. Pauli, Jr. Merkurperihelbewegung und Strahlenablenkung in Weyls Gravitationstheorie. (German) [Mercury perihelion and beam deflection in Weyl’s theory of gravitation]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 21(21–22):742–750, ???? 1919.

Pauli:1919:EGG

[Pau19b] W. Pauli, Jr. Über die Energiekomponenten des Gravitationsfeldes. (German) [On the energy component of the gravitational field]. Physikalische Zeitschrift, 20(1):25–27, January 1, 1919. CODEN PHZTAO. ISSN 0369-982X. URL http://hdl.

handle.net/2027/mdp.39015009220727.

Pauli:1919:TGE

[Pau19c] Wolfgang Pauli, Jr. Zur Theorie der Gravitation und der Elektrizität von Hermann Weyl. (German) [On the theory of gravitation and electricity of Hermann Weyl]. Physikalische Zeitschrift, 20(20):457–467, October 15, 1919. CODEN PHZ-

TAO. ISSN 0369-982X. URL http://hdl.handle.net/2027/
mdp.39015009220727.

Pauli:1920:ALB

[Pau20a] W. Pauli. Die Ausbreitung des Lichtes in bewegten Medien. (German) [The propagation of light in moving media]. Mathematische Annalen, 82(1-2):113–119, 1920. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Pauli:1920:TBD

[Pau20b] W. Pauli. Theoretische Bemerkungen über den Diamagnetismus einatomiger Gase. (German) [Theoretical remarks on the diamagnetism of monatomic gases]. Zeitschrift für Physik, 2(3): 201–205, June 1920. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.

com/article/10.1007/BF01328724;   http://www.springerlink.
com/content/g021l2j464g43j6h/fulltext.pdf.

Pauli:1920:QMG

[Pau20c] Wolfgang Pauli, Jr. Quantentheorie und Magneton. (German) [Quantum theory and magneton]. Physikalische Zeitschrift, 21 (21–22):615–617, November 1–15, 1920. CODEN PHZTAO.

ISSN 0369-982X. URL http://hdl.handle.net/2027/mdp.
39015086723080?urlappend=%3Bseq=681.

Pauli:1921:TDZ

[Pau21a] W. Pauli. Zur Theorie der Dielektrizitätskonstante zweiatomiger Dipolgase. (German) [On the theory of the dielectric constant of diatomic dipole gases]. Zeitschrift für Physik, 6(1):319–327, December 1921. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/
10.1007/BF01327993;   http://www.springerlink.com/content/
p50pg1rp960l3676/fulltext.pdf.

Pauli:1921:RGT

[Pau21b] Wolfgang Pauli. Relativitätstheorie. (German) [Theory of Relativity]. In Arnold Sommerfeld, editor, Encyklopädie der mathematischen Wissenschaften, volume 5(2), pages iv + 540–775. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1921.

LCCN ???? URL http://resolver.sub.uni-goettingen.

de/purl?PPN360709672. Foreword by Arnold Sommerfeld. New Edition with Preface and Additions (1958), with English

[Pau58f] and Italian [Pau58e] translations.

Pauli:1921:BMW

[Pau21c] Wolfgang Pauli. Über das Modell des Wasserstoff-Molekülions. (German) [On the model of the hydrogen molecular ion]. Dr. phil. thesis, Ludwig-Maximilians-Universität München, Munich, Germany, 1921.

Pauli:1922:MWG

[Pau22] W. Pauli. Über das Modell des Wasserstoffmolekülions. (German) [On the model of the hydrogen molecular ion]. Annalen der Physik (1900), 373(11):177–240, 1922. ISSN 1521-3889. Journal publication of Pauli’s thesis [Pau21c].

Pauli:1923:TGZ

[Pau23a] W. Pauli, Jr. Über das thermische Gleichgewicht zwischen Strahlung und freien Elektronen. (German) [On the thermal equilibrium between radiation and free electrons]. Zeitschrift für Physik, 18(1):272–286, December 1923. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01327708;
 http://www.springerlink.com/content/28h14wq1m6k67844/
fulltext.pdf.

Pauli:1923:GAZ

[Pau23b] W. Pauli, Jr. Über die Gesetzmäßigkeiten des anomalen Zeemaneffektes. (German) [On the laws of the anomalous Zeeman effect]. Zeitschrift für Physik, 16(1):155–164, December 1923. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01327386;   http://www.springerlink.com/content/v77p5184406p8135/
fulltext.pdf.

Pauli:1923:FZK

[Pau23c] W. Pauli, Jr. Zur Frage der Zuordnung der Komplexstrukturterme in starken und in schwachen äußeren Feldern. (German) [On the question of assignment of the complex structure terms in strong and in weak external fields]. Zeitschrift für Physik, 20(1):371–387, December 1923. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01327950;
 http://www.springerlink.com/content/q115q8ttg15k8351/
fulltext.pdf.

Pauli:1924:FTD

[Pau24a] W. Pauli. Zur Frage der theoretischen Deutung der Satelliten einiger Spektrallinien und ihrer Beeinflussung durch magnetische Felder. (German) [On the question of the theoretical interpretation of some satellite spectral lines and their influence by magnetic fields]. Naturwissenschaften, 12(37):741–743, September 1924. CODEN NATWAY. ISSN 0028-1042 (print), 1432-

1904 (electronic). URL http://link.springer.com/article/
10.1007/BF01504828;   http://www.springerlink.com/content/
q414121432406233/fulltext.pdf. See [GB27, BG28] for the

first, and independent, interpretation of experimental data for the hyperfine structure of bismuth as being due to a nuclear magnetic moment (spin). See [Gou61] for corrections to the confusing and often misattributed history of the discovery of [electron and nuclear] spin.

Pauli:1924:BRB

[Pau24b] Wolfgang Pauli. Book review: E. Buchwald, Das Korrespondenzprinzip. Naturwissenschaften, 12(??):??, ???? 1924. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1924:SGP

[Pau24c] Wolfgang Pauli. Störungstheorie. (German) [Perturbation theory]. In ????, editor, Physikalisches Handwörterbuch, pages 752–

756. Berliner und Scheel, ????, 1924. LCCN ????

Pauli:1924:BAD

[Pau24d] W. Pauli, Jr. Bemerkungen zu den Arbeiten Die Dimension der Einsteinschen Lichtquanten und Zur Dynamik des Stoßes zwischen einem Lichtquant und einem Elektron von L. S. Ornstein und H. C. Burger. (German) [Comments on the work “The dimension of Einstein’s light quanta and on the dynamics of the collision between a photon and an electron” by LS Ornstein and HC Burger]. Zeitschrift für Physik, 22(1):261–265, December 1924. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01328129;
 http://www.springerlink.com/content/wh618655kn066784/
fulltext.pdf.

Pauli:1925:BRP

[Pau25a] Wolfgang Pauli. Book review: E. P. Adams, The Quantum theory. Naturwissenschaften, 12(??):412–413, ???? 1925. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1925:BRM

[Pau25b] Wolfgang Pauli. Book review: M. Born, Vorlesungen über Atommechanik. 1. Teil. Naturwissenschaften, 13(??):487–488, ????

1925. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1925:ZAE

[Pau25c] Wolfgang Pauli. Über den Zusammenhang des Abschlusses der Elektronengruppen im Atom mit der Komplexstruktur der Spektren. (German) [On the relation of the completion of groups of electrons in the atom with the complex structure of spectra]. Zeitschrift für Physik, 31(1):765–783, February 1925. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF02980631;
 http://www.springerlink.com/content/20w1m0vr050j033r/

fulltext.pdf. This is the paper in which Pauli introduced the famous Exclusion Principle, for which he received the Nobel Prize in Physics 1945 [Pau45b].

Pauli:1925:ARK

[Pau25d] Wolfgang Pauli. Über die Absorption der Reststrahlen in Kristallen. (German) [On the absorption of residual rays in crystals]. Verh. Dtsch. Phys. Ges. (3), 6(??):10–11, ???? 1925.

Pauli:1925:IIE

[Pau25e] Wolfgang Pauli. Über die Intensitäten der im elektrischen Felde erscheinenden Kombinationslinien. (German) [On the intensities of the combination lines appearing in electric fields]. Dan. Mat. Fys. Medd., 7(3):??, ???? 1925.

Pauli:1925:EGE

[Pau25f] Wolfgang Pauli, Jr. Über den Einfluß der Geschwindigkeitsabhängigkeit der Elektronenmasse auf den Zeemaneffekt. (German) [On the influence of the velocity dependence of electron mass on the Zeeman effect]. Zeitschrift für Physik, 31(1):373–385, February 1925. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831

(electronic). URL http://link.springer.com/article/10.
1007/BF02980592;   http://www.springerlink.com/content/
e6631774u53u8841/.

Pauli:1926:BRE

[Pau26a] Wolfgang Pauli. Book review: A. S. Eddington, Relativitätstheorie in mathematischer Behandlung. Naturwissenschaften, 14(??):273–274, ???? 1926. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1926:QGQ

[Pau26b] Wolfgang Pauli. Quantentheorie. (German) [Quantum theory]. In ????, editor, Handbuch der Physik, volume 23, pages 1–278. Geiger and Scheel, ????, 1926. LCCN ????

Pauli:1926:WSN

[Pau26c] W. Pauli, Jr. Über das Wasserstoffspektrum vom Standpunkt der neuen Quantenmechanik. (German) [On the hydrogen spectrum, from the standpoint of the new quantum mechanics]. Zeitschrift für Physik, 36(5):336–363, May 1926. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01450175;
 http://www.springerlink.com/content/n14x167240x812ru/
fulltext.pdf.

Pauli:1927:GPG

[Pau27a] W. Pauli, Jr. Über Gasentartung und Paramagnetismus. (German) [On gas degeneracy and paramagnetism]. Zeitschrift für Physik, 41(6–7):81–102, June 1927. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01391920;   http://www.
springerlink.com/content/n25735226785627j/fulltext.
pdf.

Pauli:1927:QME

[Pau27b] W. Pauli, Jr. Zur Quantenmechanik des magnetischen Elektrons. (German) [On the quantum mechanics of the magnetic electron]. Zeitschrift für Physik, 43(9–10):601–623, September 1927. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831

(electronic). URL http://link.springer.com/article/10.
1007/BF01397326;   http://www.springerlink.com/content/
ww10g288m777t400/fulltext.pdf.

Pauli:1928:TAE

[Pau28] Wolfgang Pauli. Über das H-Theorem vom Anwachsen der Entropie vom Standpunkt der neuen Quantenmechanik. (German) [On the H-theorem of entropy increase from the standpoint of the new quantum mechanics]. In ????, editor, Probleme der modernen Physik, Arnold Sommerfeld zum 60. Geburtstage, gewidmet von seinen Schülern. (German) [Problems of modern physics, Arnold Sommerfeld’s 60th Birthday, dedicated by his students], pages 30–45. ????, Leipzig, Germany, 1928. LCCN ????

Pauli:1929:AGQ

[Pau29a] Wolfgang Pauli. Allgemeine grundlagen der quantentheorie des atombaues. In Müller–Pouillets Lehrbuch, volume 2(2), pages 1709–1842. ????, ????, 11 edition, 1929. LCCN ????

Pauli:1929:BRB

[Pau29b] Wolfgang Pauli. Book review: Ergebnisse der exakten Naturwissenschaften. Band 7. Naturwissenschaften, 17(??):257–259, ???? 1929. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1929:BRH

[Pau29c] Wolfgang Pauli. Book review: H. A. Lorentz, Vorlesungen über theoretische Physik. Naturwissenschaften, 17(??):279, ???? 1929. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1929:TSS

[Pau29d] Wolfgang Pauli. Theorie der schwarzen Strahlung. (German) [Theory of black-body radiation]. In ????, editor, Müller– Pouillets Lehrbuch, volume 2(2), pages 1483–1553. ????, ????, 1929. LCCN ????

Pauli:1930:BRB

[Pau30a] Wolfgang Pauli. Book review: Ergebnisse der exakten Naturwissenschaften, Band 8. Naturwissenschaften, 18(??):568–570, ???? 1930. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1930:BRM

[Pau30b] Wolfgang Pauli. Book review: M. Born and P. Jordan, Elementare Quantenmechanik. Naturwissenschaften, 18(??):602, ???? 1930. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1930:BRP

[Pau30c] Wolfgang Pauli. Book review: P. A. M. Dirac, The principles of quantum mechanics. Naturwissenschaften, 19(??):188, ???? 1930. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1930:LTC

[Pau30d] Wolfgang Pauli. Letter to Tübingen conference participants. Web document., December 4, 1930. English translation in [Bro78, page 27].

Pauli:1931:AVL

[Pau31a] Wolfgang Pauli. Antwort bei Verleihung der Lorentzmedaille. (German) [Response in the Lorentz Medal Ceremony]. Proc. Acad. Sci. Amst., 40(??):125–126, ???? 1931.

Pauli:1931:BRM

[Pau31b] Wolfgang Pauli. Book review: A. March, Die Grundlagen der Quantenmechanik. Naturwissenschaften, 19(??):867, ???? 1931. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1931:BRW

[Pau31c] Wolfgang Pauli. Book review: W. Heisenberg, Die physikalischen Prinzipien der Quantentheorie. Naturwissenschaften, 19 (??):188–189, ???? 1931. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1931:NP

[Pau31d] Wolfgang Pauli. [the neutrino postulate]. Talk presented at The Present Status of the Problem of Nuclear Structure, a conference held in Pasadena, CA, USA. Pauli was still uncertain about the existence of the neutrino, so he refused to allow his paper to be published in the proceedings [see [?, page 231]]. His first publication on the neutrino is in [CCJ^{+}34] (contents and titles not yet located)., June 1931.

Pauli:1932:BRB

[Pau32a] Wolfgang Pauli. Book review: Ergebnisse der exakten Naturwissenschaften, Band 10. Naturwissenschaften, 20(??):186–187, ???? 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1932:BRG

[Pau32b] Wolfgang Pauli. Book review: G. Gamow, Der Bau des Atomkernes und die Radioaktivität. Naturwissenschaften, 20(??):582, ???? 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1932:DWE

[Pau32c] Wolfgang Pauli. Diracs Wellengleichung des Elektrons und geometrische Optik. (German) [Dirac’s wave equation of the electron and geometric optics]. Helvetica Physica Acta, 5(3):179– 199, ???? 1932. CODEN HPACAK. ISSN 0018-0238.

Pauli:1932:TQD

[Pau32d] Wolfgang Pauli. Les theories quantiques du magnetisme in ‘L’électron magnetique’. (French) [Quantum theories of magnetism in ‘the magnetic electron’]. In ????, editor, Sixieme Conseil de Physique Solvay, Bruxelles, 1930, pages 175–238, discussion 239–280. ????, ????, 1932. LCCN ????

Pauli:1933:EQB

[Pau33a] W. Pauli. Einige die Quantenmechanik betreffenden Erkundigungsfragen. (German) [Some exploration of issues relevant to quantum mechanics]. Zeitschrift für Physik, 80(9–10):573–586, September 1933. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic). URL http://link.springer.com/
article/10.1007/BF01335695;   http://www.springerlink.
com/content/x711540mlj227376/fulltext.pdf.

Pauli:1933:PEO

[Pau33b] W. Pauli. Paul Ehrenfest [obituary]. Naturwissenschaften, 21 (48):841–843, ???? 1933. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://link.springer.

com/article/10.1007/BF01504522;   http://www.springerlink.
com/content/u0nx823747245437/fulltext.pdf.

Pauli:1933:FNFa

[Pau33c] W. Pauli. Über die Formulierung der Naturgesetze mit fünf homogenen Koordinaten Teil I: Klassische Theorie. (German) [On the formulation of the laws of nature with five homogeneous coordinates, Part I: Classical theory]. Annalen der Physik (Series 5), 410(3):305–336, 1933. CODEN ANPYA2. ISSN 1521-3889.

Pauli:1933:FNFb

[Pau33d] W. Pauli. Über die Formulierung der Naturgesetze mit fünf homogenen Koordinaten. Teil II: Die Diracschen Gleichungen für die Materiewellen. (German) [On the formulation of the laws of nature with five homogeneous coordinates. Part II: The Dirac equations for matter waves]. Annalen der Physik (Series 5), 410 (4):337–372, 1933. CODEN ANPYA2. ISSN 1521-3889.

Pauli:1933:BRB

[Pau33e] Wolfgang Pauli. Book review: Ergebnisse der exakten Naturwissenschaften, Band 11. Naturwissenschaften, 21(??):301–302, ???? 1933. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1933:BRJ

[Pau33f] Wolfgang Pauli. Book review: J. H. van Vleck, The theory of electric and magnetic susceptibilities. Naturwissenschaften, 21(??):239, ???? 1933. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1933:APW

[Pau33g] Wolfgang Pauli. Die allgemeinen Prinzipien der Wellenmechanik. (German) [The general principles of wave mechanics]. In Handbuch der Physik, volume XXIV(1), pages 83– 272. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1933. According to [De 14, page 76, column 1], this may be the publication where “wave–particle duality” was first defined.

Pauli:1933:ISB

[Pau33h] Wolfgang Pauli. Über die Intensität der Streustrahlung bewegter freier Elektronen. (German) [On the intensity of the scattered radiation of moving free electrons]. Helvetica Physica Acta, 6(4):279–286, ???? 1933. CODEN HPACAK. ISSN 0018-0238.

Pauli:1934:DPM

[Pau34] Wolfgang Pauli. Discussion of paper by M. Heisenberg on “La Structure du Noyau”. In Cockcroft et al. [CCJ^{+}34], pages 324– 325. LCCN ???? Publiés par la commission administrative de l’institut.

Pauli:1935:BMT

[Pau35a] Wolfgang Pauli. Beitrage zur mathematischen Theorie der Dirac’schen Matrizen. (German) [Contributions to the mathematical theory of the Dirac matrices]. In ????, editor, Zeeman Verhandelingen (Den Haag), pages 31–43. ????, ????, 1935. LCCN ????

Pauli:1935:BRP

[Pau35b] Wolfgang Pauli. Book review: P. Debye, Kernphysik. Naturwissenschaften, 23(??):772–773, ???? 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1936:TQR

[Pau36a] W. Pauli. Théorie quantique relativiste des particules obéissant à la statistique de Einstein–Bose. (French) [Relativistic quantum theory of particles obeying Einstein–Bose statistics]. Annales de l’Institut Henri Poincaré, 6(2):137–152, 1936. CODEN AIHPA2. ISSN 0365-320x (print), 2400-4855 (electronic). URL

http://www.numdam.org/item?id=AIHP_1936__6_2_137_0.

Pauli:1936:CMT

[Pau36b] Wolfgang Pauli. Contributions mathématiques à la théorie des matrices de Dirac. (French) [Mathematical contributions to the theory of Dirac matrices]. Annales de l’Institut Henri Poincaré, 6(2):109–136, ???? 1936. CODEN AIHPA2. ISSN 0365-320x (print), 2400-4855 (electronic). URL http://www.numdam.org/

item?id=AIHP_1936__6_2_109_0.

Pauli:1936:RZK

[Pau36c] Wolfgang Pauli. Raum, Zeit und Kausalitat in der modernen Physik. (German) [Space, time and causality in modern physics]. Scientia, 59(??):65–76, ???? 1936.

Pauli:1938:ASF

[Pau38a] Wolfgang Pauli. On asymptotic series for functions in the theory of diffraction of light. Physical Review (2), 54(11):924– 931, December 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v54/i11/p924_1.

Pauli:1938:SCP

[Pau38b] Wolfgang Pauli. Some considerations of principle on the theory of P decay. (Russian). Bull. Acad. Sci. URSS (Série physique), ??(??):149–152, ???? 1938.

Pauli:1939:KZE

[Pau39] W. Pauli. Über ein Kriterium für Ein- oder Zweiwertigkeit der Eigenfunktionen in der Wellenmechanik. (German) [On a criterion for one- or two-valuedness of eigenfunctions in wave mechanics]. Helvetica Physica Acta, 12(2):147–168, ???? 1939. CODEN HPACAK. ISSN 0018-0238.

Pauli:1940:CBS

[Pau40a] W. Pauli. The connection between spin and statistics. Physical Review (2), 58(8):716–722, October 15, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v58/i8/p716_1.

Pauli:1940:IDW

[Pau40b] W. Pauli. Über die Invarianz der Dirac’schen Wellengleichungen gegenüber Ähnlichkeitstransformationen des Linienelementes im Fall verschwindender Ruhmasse. (German) [On the invariance of the Dirac wave equations compared to similarity transformations of the line element in the case of vanishing rest mass]. Helvetica Physica Acta, 13(3):204–208, ???? 1940. CODEN HPACAK. ISSN 0018-0238.

Pauli:1941:RFT

[Pau41] W. Pauli. Relativistic field theories of elementary particles. Reviews of Modern Physics, 13(3):203–232, July 1, 1941. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.13.203;   http://rmp.aps.org/
abstract/RMP/v13/i3/p203_1.

Pauli:1943:ALP

[Pau43a] W. Pauli. On applications of the λ-limiting process to the theory of the meson field. Physical Review (2), 64(11–12):332– 344, December 1, 1943. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v64/i11-12/p332_1.

Pauli:1943:DNM

[Pau43b] W. Pauli. On Dirac’s new method of field quantization. Reviews of Modern Physics, 15(3):175–207, July 1, 1943. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.15.175;   http://rmp.aps.org/
abstract/RMP/v15/i3/p175_1.

Pauli:1943:SCW

[Pau43c] Wolfgang Pauli. On strong coupling and weak coupling theories of the meson field. Physical Review (2), 63(5–6):221, March 1943. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v63/

i5-6/p215_1. Bull. Am. Phys. Soc. New York Meeting, Jan. 22–23, 1943, Abstract No. 25.

Pauli:1945:NBH

[Pau45a] W. Pauli. Niels Bohr on his 60th birthday. Reviews of Modern Physics, 17(2–3):97–101, April 1, 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.17.
97;   http://rmp.aps.org/abstract/RMP/v17/i2-3/p97_1.

Pauli:1945:NPP

[Pau45b] Wolfgang Pauli. The Nobel Prize in Physics 1945. Nobel-

prize.org, 1945. URL http://www-groups.dcs.st-and.ac.
uk/~history/Extras/Pauli_Exclusion_Principle.html;
http://www.nobelprize.org/nobel_prizes/physics/laureates/

1945/pauli-bio.html. The Nobel Prize in Physics 1945 was awarded to Wolfgang Pauli “for the discovery of the Exclusion Principle, also called the Pauli Principle”.

Pauli:1946:DFE

[Pau46a] Wolfgang Pauli. Diracs Feldquantisierung und Emission von Photonen kleiner Frequenzen. (German) [Dirac field quantization and emission of photons of lower frequencies]. Helvetica Physica Acta, 19(4):234–237, ???? 1946. CODEN HPACAK.

ISSN 0018-0238. URL http://retro.seals.ch/digbib/view?
rid=hpa-001:1946:19::236&id=hitlist.

Pauli:1946:MTN

[Pau46b] Wolfgang Pauli. Meson Theory of Nuclear Forces. Interscience Publishers, New York, NY, USA, 1946. 69 pp.

Pauli:1946:RHE

[Pau46c] Wolfgang Pauli. Remarks on the history of the Exclusion Principle. Science, 103(2669):213–215, February 22, 1946. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/103/
2669/213.full.pdf.

Pauli:1947:BRH

[Pau47a] Wolfgang Pauli. Book review: H. Reichenbach, Philosophic Foundations of Quantum Mechanics. Dialectica: International Review of Philosophy of Knowledge, 1(??):176–178, ???? 1947. ISSN 0012-2017 (print), 1746-8361 (electronic).

Pauli:1947:DFT

[Pau47b] Wolfgang Pauli. Difficulties of field theories and of field quantization. In ????, editor, Phys. Soc. Cambridge Conf. Report, pages 5–10. ????, ????, 1947. LCCN ????

Pauli:1947:MBO

[Pau47c] Wolfgang Pauli. Eine Methode zur Bestimmung von Oszillatorenstärken (f-Werten) aus dem Starkeffekt. (German) [A method for determining oscillator strengths (f-values) from the Stark effect]. Zeitschrift für Physik, 124(1–2):121–128, January 1947. CODEN ZEPYAA. ISSN 0044-3328. URL https://

link.springer.com/article/10.1007/BF01374927.

Pauli:1947:EPQ

[Pau47d] Wolfgang Pauli. Exclusion Principle and quantum mechanics: Discours prononcé à la réception du prix Nobel de physique 1945. (French) [Lecture given on receiving the 1945 Nobel Prize in Physics]. Dialectica: International Review of Philosophy of Knowledge, 1(2):204, 1947. ISSN 0012-2017 (print), 1746-8361 (electronic).

Pauli:1947:SMG

[Pau47e] Wolfgang Pauli. Statistische Mechanik. (German) [Statistical Mechanics]. Verein der Matematiker und Physiker an der ETH, Zurich, Switzerland, 1947. ???? pp. LCCN ???? Prepared by M. R. Schafroth.

Pauli:1948:EIK

[Pau48a] W. Pauli. Editorial: Die Idee der Komplementarität. (German) [The concept of complementarity]. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):307–311, 1948. ISSN 0012-2017 (print), 1746-8361 (electronic).

Pauli:1948:SBQ

[Pau48b] W. Pauli. Sommerfelds beiträge zur Quantentheorie. (German) [Sommerfeld’s contributions to quantum theory]. Naturwissenschaften, 35(5):129–132, ???? 1948. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://link.springer.com/article/10.1007/BF00631594;

 http://www.springerlink.com/content/r6v820q252r80222/
fulltext.pdf.

Pauli:1949:ECQ

[Pau49] Wolfgang Pauli. Einstein’s contributions to quantum theory. In Schilpp [Sch49a], chapter 4, pages 147–160. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Pauli:1950:CBS

[Pau50a] W. Pauli. On the connection between spin and statistics. Progress of Theoretical Physics, 5:526–543, 1950. CODEN PTP-KAV. ISSN 0033-068X (print), 1347-4081 (electronic).

Pauli:1950:PBI

[Pau50b] Wolfgang Pauli. Die philosophische Bedeutung der Idee der Komplementaritat. (German) [The philosophical significance of the concept of complementarity]. Experientia, 6(2):72–81, ???? 1950. CODEN EXPEAM. ISSN 0014-4754.

Pauli:1951:AS

[Pau51a] W. Pauli. Arnold Sommerfeld. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 2(4):301, ???? 1951. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF02579693;
 http://www.springerlink.com/content/c8110842238148n6/
fulltext.pdf.

Pauli:1951:ASOa

[Pau51b] Wolfgang Pauli. Arnold Sommerfeld [obituary]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 2(??):301–??, ???? 1951. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Pauli:1951:ASOb

[Pau51c] Wolfgang Pauli. Arnold Sommerfeld [obituary]. Naturforsch., 6a(??):301–??, ???? 1951.

Pauli:1951:BRS

[Pau51d] Wolfgang Pauli. Book review: A. Sommerfeld, Vorlesungen iuber theoretische Physik, Band IV. Optik. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 2(3):215, May 1951. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Pauli:1951:BRA

[Pau51e] Wolfgang Pauli. Book review: Arnold Sommerfeld, Vorlesungen iuber theoretische Physik, Band IV. Optik. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 2(3):215, May 1951. CODEN ZAM-PDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL

http://link.springer.com/article/10.1007/BF02579762;
 http://www.springerlink.com/content/13817655g46060p8/
fulltext.pdf.

Pauli:1952:BWS

[Pau52a] Wolfgang Pauli. Der Begriff der Wahrscheinlichkeit und seine Rolle in den Naturwissenschaften. (German) [The concept of probability and its role in the natural sciences]. Verh. Schweiz. Naturf. Ges., ??(??):76–79, ???? 1952. Aarau.

Pauli:1952:EAV

[Pau52b] Wolfgang Pauli. Der Einfluss archetypischer Vorstellungen auf die Bildung naturwissenschaftlicher Theorien bei Kepler. (German) [The influence of archetypal ideas on the formation of scientific theories of Kepler]. In Jung and Pauli [JP52], pages 109–194. LCCN BF1033 .N38. Available in English translation [JP55].

Pauli:1952:GPS

[Pau52c] Wolfgang Pauli. Die Geschichte des periodischen Systems der Elemente. (German) [The history of the Periodic Table of the Elements]. Vjschr. naturf. Ges. Zurich, 97(??):138–139, ???? 1952.

Pauli:1952:TEG

[Pau52d] Wolfgang Pauli. Theorie und Experiment. (German) [Theory and experiment]. Dialectica: International Review of Philosophy of Knowledge, 6(2):141–142, June 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pauli:1953:HSN

[Pau53a] W. Pauli. On the Hamiltonian structure of non-local field theories. Il Nuovo Cimento (9), 10:648–667, 1953. CODEN NU-CIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Pauli:1953:RPP

[Pau53b] W. Pauli. Remarques sur le problème des paramètres cachés dans la mécanique quantique et sur la théorie de l’onde pilote. (French) [Remarks on the problem of hidden variables in quantum mechanics and on the pilot-wave theory]. In Louis de Broglie, physicien et penseur, pages 33–42. Editions Albin Michel, Paris, 1953.

Pauli:1953:DRT

[Pau53c] Wolfgang Pauli. Discussion remark to the talk Isotopic spin and mass quantization by A. Pais. Physica, 19(??):887, ???? 1953. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Pauli:1953:EAT

[Pau53d] Wolfgang Pauli. Etat actuel de la theorie quantique des champs. La renormalization. (French) [Current state of quantum field theory. Renormalization]. In ????, editor, Particules fondamentales et noyaux: Paris 1950. Colloques internationaux du Centre National de la Recherche Scientifique, volume 38, pages 67–77. ????, Paris, France, 1953. LCCN ????

Pauli:1954:M

[Pau54a] Wolfgang Pauli. Matter. In H. Muschel, editor, Man’s Right to Knowledge, pages 10–18. Columbia University Press, New York, NY, USA, 1954. LCCN ????

Pauli:1954:NEA

[Pau54b] Wolfgang Pauli. Naturwissenschaftliche und erkenntnistheoretische Aspekte der Ideen vom Unbewussten. (German) [Natural sciences and epistemological aspects of the ideas of the unconscious]. Dialectica: International Review of Philosophy of Knowledge, 8(4):283–301, December 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pauli:1954:WPG

[Pau54c] Wolfgang Pauli. Wahrscheinlichkeit und Physik. (German) [Probability and physics]. Dialectica: International Review of Philosophy of Knowledge, 8(2):112–118; Diskussion 118–124, June 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pauli:1955:EPL

[Pau55a] W. Pauli. Exclusion principle, Lorentz group and reflection of space-time and charge. In W. Pauli, L. Rosenfeld, and V. Weisskopf, editors, Niels Bohr and the development of physics: Essays Dedicated to Niels Bohr on the Occasion of His Seventieth Birthday, pages 30–51. McGraw-Hill, New York, NY, USA, 1955.

Pauli:1955:EBQ

[Pau55b] Wolfgang Pauli. Einsteins Beitrag zur Quantentheorie. (German) [Einstein’s contribution to quantum theory]. In ????, editor, Albert Einstein als Philosoph und Naturforscher. (German) [Albert Einstein as philosopher and scientist], pages 74–83. ????, Stuttgart, West Germany, 1955. LCCN ????

Pauli:1955:RPS

[Pau55c] Wolfgang Pauli. Rydberg and the periodic system of the elements. In ????, editor, Proceedings of the Rydberg Centennial Conference on Atomic Spectroscopy, Lund, 1954, pages 22–26. ????, Lund, Sweden, 1955. LCCN ????

Pauli:1956:RPC

[Pau56a] W. Pauli. Remarks on problems connected with the renormalization of quantized fields. Il Nuovo Cimento (10), 4 (supplemento):703–710, 1956. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Pauli:1956:A

[Pau56b] Wolfgang Pauli. Announcement. CERN Symposium, 1956.

Pauli:1956:ADP

[Pau56c] Wolfgang Pauli. Ansprache durch den Präsident der Konferenz; Zusammenfassung und Schlußwort durch den Präsident der Konferenz; Relativitätstheorie und Wissenschaft. (German) [Opening remarks by the Conference President; closing remarks by the Conference President: Relativity theory and science]. Helvetica Physica Acta, 29(Suppl. IV):27, 261–267, 282–286, 1956. CODEN HPACAK. ISSN 0018-0238. Fünfzig Jahre Relativitätstheorie’, Bern 1955.

Pauli:1956:BRD

[Pau56d] Wolfgang Pauli. Book review: D. ter Haar, Elements of statistical mechanics. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 7(??): 467, ???? 1956. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Pauli:1956:CGQ

[Pau56e] Wolfgang Pauli. Continuous groups in quantum mechanics. Report CERN 56-31, CERN, Geneva, Switzerland, 1956.

Pauli:1956:WAD

[Pau56f] Wolfgang Pauli. Die Wissenschaft und das abendländische Denken. (German) [Science and Western thought]. In M. Göhring, editor, ‘Europa — Erbe und Aufgabe’. Internationaler Gelehrtenkongress, Mainz, 1955. (German) [Europe heritage and task. International scholars congress], pages 71–79. ????, Wiesbaden, West Germany, 1956. LCCN ???? Reprinted in [Pau59].

Pauli:1957:CLC

[Pau57a] W. Pauli. On the conservation of the lepton charge. Il Nuovo Cimento (10), 6:204–215, 1957. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Pauli:1957:PPR

[Pau57b] Wolfgang Pauli. Phänomen und Physikalische Realität. (German) [Phenomena and physical reality]. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):36–48, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). Lecture of 24 August 1954 at the International Philosophy Congress in Zurich, Switzerland.

Pauli:1957:ANG

[Pau57c] Wolfgang Pauli. Zur älteren und neueren Geschichte des Neutrinos. (German) [On old and new history of the neutrino]. Vierteljahresschrift der Naturforschenden Gesellschaft Zürich, 102(??): 387–388, ???? 1957. Abstract of lecture at Zurich, 21 January 1957. Reprinted in [Pau61d].

Pauli:1958:AEE

[Pau58a] Wolfgang Pauli. Albert Einstein in der Entwicklung der Physik. (German) [Albert Einstein in the development of physics]. Neue Zürcher Zeitung, 89(??):??, January 12, 1958.

Pauli:1958:APW

[Pau58b] Wolfgang Pauli. Die allgemeinen Prinzipien der Wellenmechanik. (German) [The general principles of wave mechanics]. In Siegfried Flügge, editor, Handbuch der Physik, volume 5(1), pages 1–168. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1958. LCCN ????

Pauli:1958:VSS

[Pau58c] Wolfgang Pauli. Die Verletzung von Spiegelungs-Symmetrien in den Gesetzen der Atomphysik. (German) [The violation of reflection symmetries in the laws of atomic physics]. Experientia, 14(1):1–5, ???? 1958. CODEN EXPEAM. ISSN 0014-4754.

Pauli:1958:IMC

[Pau58d] Wolfgang Pauli. The indefinite metric with complex roots. In ????, editor, 1958 Annual International Conference on High Energy Physics at CERN, Proceedings, pages 127–128. ????, ????, 1958. LCCN ????

Pauli:1958:TDR

[Pau58e] Wolfgang Pauli. Teoria della Relatività. (Italian) [Theory of Relativity]. Paolo Boringhieri, Torino, Italy, 1958. xiv + 327 pp. LCCN ???? Translation by Paolo Gulmanelli from the German original [Pau21b].

Pauli:1958:TR

[Pau58f] Wolfgang Pauli. Theory of Relativity. Pergamon, New York, NY, USA, 1958. xiv + 241 pp. LCCN ???? English translation by G. Field from the German original [Pau21b], with supplementary notes by the author.

Pauli:1958:TDGa

[Pau58g] Wolfgang Pauli. Zur Thermodynamik dissoziierter Gleichgewichtsgemische in äusseren Kraftfeldern. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 9b:490–497, 1958. CODEN ZAM-PDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Pauli:1958:TDGb

[Pau58h] Wolfgang Pauli. Zur Thermodynamik dissoziierter Gleichgewichtsgemische in äusseren Kraftfeldern. (German) [On the thermodynamics of equilibrium mixtures of dissociated in external force fields]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 9(5–6):490–497, March 1958. CODEN ZAM-PDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL

http://link.springer.com/article/10.1007/BF02424768;
 http://www.springerlink.com/content/b3xj217k4518667x/
fulltext.pdf.

Pauli:1959:WAD

[Pau59] Wolfgang Pauli. Die Wissenschaft und das abendländische Denken. (German) [Science and Western thought]. Schweize. Bauzeitung, 77(??):1–4, ???? 1959. Reprint of [Pau56f].

Pauli:1961:AVP

[Pau61a] Wolfgang Pauli. Aufsätze und Vorträge über Physik und Erkenntnistheorie. (German) [Essays and lectures on physics and epistemology], volume 115 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1961. 183 pp. LCCN QC6 .P317.

Pauli:1961:AVB

[Pau61b] Wolfgang Pauli. Aufsätze und Vorträge über Physik und Erkenntnistheorie. (German) [Essays and lectures on physics and epistemology], volume 115 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1961. 183 pp. LCCN ????

Pauli:1961:IAE

[Pau61c] Wolfgang Pauli. Impressionen über Albert Einstein. (German) [Impressions of Albert Einstein]. In Aufsätze und Vorträge über Physik und Erkenntnistheorie. (German) [Essays and lectures on physics and epistemology] [Pau61b], pages 81–?? LCCN ????

Pauli:1961:ANG

[Pau61d] Wolfgang Pauli. Zur älteren und neueren Geschichte des Neutrinos. (German) [On old and new history of the neutrino]. In Aufsätze und Vorträge über Physik und Erkenntnistheorie. (German) [Essays and lectures on physics and epistemology] [Pau61b], pages 156–180. LCCN ???? Reprint of [Pau57c].

Pauli:1964:CSP

[Pau64a] Wolfgang Pauli. Collected Scientific Papers. Interscience Publishers, New York, NY, USA, 1964. xxvii + 1133 (volume 1), xiii + 1408 (volume 2) pp. LCCN QC3 P32. Edited by Ralph Kronig and Victor Frederick Weisskopf.

Pauli:1964:EPQ

[Pau64b] Wolfgang Pauli. Exclusion principle and quantum mechanics. In Anonymous [Ano64], pages 27–43. LCCN ????

Pauli:1965:CGQ

[Pau65a] W. Pauli. Continuous groups in quantum mechanics. Ergebnisse der Exakten Naturwissenschaften, 37:85–104, 1965. CODEN EENAA3. ISSN 0367-0325.

Pauli:1965:VEV

[Pau65b] Wolfgang Pauli. Vklad Einsteina v kvantovuyu teoriyu. (Russian) [The contribution of Einstein’s quantum theory]. Uspekhi Fizicheskikh Nauk, 86(3):413–420, July 1965. CODEN UF-NAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1965/7/c/.

Pauli:1970:RZG

[Pau70] Hertha Ernestine Pauli. Der Riß der Zeit geht durch mein Herz. (German) [The tearing of time pierces my heart]. Paul Zsolnay, Vienna, Austria and Hamburg, Germany, 1970. 270 pp. LCCN

PT2631.A77 Z5. URL http://www.univie.ac.at/biografiA/
PauliTagung/HerthaPauli.htm.

Pauli:1972:BT

[Pau72a] Hertha Pauli. Break of time. Hawthorn Books, New York, NY, USA, 1972. vii + 239 pp. LCCN D811.5 .P3513 1972.

Pauli:1972:DDT

[Pau72b] Hertha Pauli. La Déchirure du temps: roman (French) [The tearing of time: novel]. Presses de la cité, Paris, France, 1972. 246 pp. LCCN ????

Pauli:1973:E

[Pau73a] Wolfgang Pauli, editor. Electrodynamics, volume 1 of Pauli lectures on physics. MIT Press, Cambridge, MA, USA, 1973. ISBN 0-262-16046-3. x + 160 pp. LCCN QC3 .P35 vol. 1. Translated by S. Margulies and H. R. Lewis. Foreword by Victor F. Weisskopf.

Pauli:1973:OTE

[Pau73b] Wolfgang Pauli, editor. Optics and the theory of electrons, volume 2 of Pauli lectures on physics. MIT Press, Cambridge, MA, USA, 1973. ISBN 0-262-16047-1. x + 159 pp. LCCN QC3 .P35 vol. 2. Translated by S. Margulies and H. R. Lewis. Foreword by Victor F. Weisskopf.

Pauli:1973:STF

[Pau73c] Wolfgang Pauli, editor. Selected topics in field quantization, volume 6 of Pauli lectures on physics. MIT Press, Cambridge, MA, USA, 1973. ISBN 0-262-16051-X. x + 188 pp. LCCN QC3 .P35 vol. 6. Translated by S. Margulies and H. R. Lewis. Foreword by Victor F. Weisskopf.

Pauli:1973:SM

[Pau73d] Wolfgang Pauli, editor. Statistical mechanics, volume 4 of Pauli lectures on physics. MIT Press, Cambridge, MA, USA, 1973. ISBN 0-262-16049-8. x + 121 pp. LCCN QC3 .P35 vol. 4. Translated by S. Margulies and H. R. Lewis. Foreword by Victor F. Weisskopf.

Pauli:1973:TKT

[Pau73e] Wolfgang Pauli, editor. Thermodynamics and the kinetic theory of gases, volume 3 of Pauli lectures on physics. MIT Press, Cambridge, MA, USA, 1973. ISBN 0-262-16048-X. x + 138 pp. LCCN QC3 .P35 vol. 3. Translated by S. Margulies and H. R. Lewis. Foreword by Victor F. Weisskopf.

Pauli:1973:WM

[Pau73f] Wolfgang Pauli, editor. Wave mechanics, volume 5 of Pauli lectures on physics. MIT Press, Cambridge, MA, USA, 1973. ISBN 0-262-16050-1. xv + 205 pp. LCCN QC3 .P35 vol. 5. Translated by S. Margulies and H. R. Lewis. Foreword by Victor F. Weisskopf.

Pauli:1974:VIA

[Pau74] Wolfgang Pauli. Vierpoltheorie und ihre Anwendung auf elektronische Schaltungen. (German) [Quadrapole theory and its application to electronic circuits]. Akademie-Verlag, Berlin, Germany, 1974. 234 pp. Wissenschaftliche Taschenbücher, Band 135.

Pauli:1975:TPK

[Pau75a] Vol′fgang Pauli. Trudy po kvantovoi teorii. Izdat. “Nauka”, Moscow, USSR, 1975. 687 + 1 pp. Kvantovaya teoriya. Obshchie printsipy volnovoi mekhaniki. Stati 1920–1928. [Quantum theory. General principles of wave mechanics. Papers 1920–1928], Edited by Ja. A. Smorodinskiĭ, Translated from the German by Ju. A. Danilov and A. A. Sazykin, Klassiki Nauki. [Classics of Science Series].

Pauli:1975:QTM

[Pau75b] Wolfgang Pauli. 6. quantum theory of magnetism: The magnetic electron. In Mehra [Meh75], chapter 7, pages 194–199. ISBN 90-277-0635-2. LCCN QC1.S792 M43.

Pauli:1977:TPK

[Pau77a] Vol′fgang Pauli. Trudy po kvantovoi teorii. Izdat. “Nauka”, Moscow, USSR, 1977. 695 pp. Stati 1928–1958. [Papers 1928– 1958], Edited by Ja. A. Smorodinskiĭ, Translated from the German and French by Ju. A. Danilov and A. A. Sazykin, Seriya “Klassiki Nauki” [“Classics in Science” series].

Pauli:1977:FAA

[Pau77b] Wolfgang Pauli. Fünf Arbeiten Zum Ausschliessungsprinzip Und Zum Neutrino. (German) [Five works on the Exclusion Principle and the neutrino], volume 27 of Texte zur Forschung. Wissenschaftliche Buchgesellschaft, ????, 1977. ISBN 3-534-06733-9. 121 pp. LCCN ????

Pauli:1980:GPQ

[Pau80] Wolfgang Pauli. General principles of quantum mechanics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 3-540-09842-9. xii + 212 pp. Translated from the German by P. Achuthan and K. Venkatesan, With an introduction by Charles P. Enz.

Pauli:1981:TR

[Pau81] Wolfgang Pauli. Theory of Relativity. Dover, New York, NY, USA, 1981. ISBN 0-486-64152-X. xiv + 241 pp. LCCN QC173.55 .P3813 1981. Translation by G. Field from the German original [Pau21b].

Pauli:1985:JNG

[Pau85a] Wolfgang Pauli. Das Jahr 1930 Die Neutrinohypothese. (German) [The year 1930: The neutrino hypothesis]. In Sources in the History of Mathematics and Physical Sciences [Pau85b], pages 1–48. ISBN 3-540-78801-8.

Pauli:1985:SHM

[Pau85b] Wolfgang Pauli, editor. Sources in the History of Mathematics and Physical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 3-540-78801-8. ???? pp.

Pauli:1990:RZG

[Pau90a] Hertha Pauli. Der Riß der Zeit geht durch mein Herz: Erlebtes, Erzähltes. (German) [The tearing of time pierces my heart: experienced, narrated], volume 30243 of Ullstein-Buch Die Frau in der Literatur. Ullstein, Frankfurt am Main and Berlin, Germany, 1990. ISBN 3-548-30243-2. 280 pp. LCCN ???? URL

http://www.gbv.de/dms/hbz/toc/ht003736613.pdf.

Pauli:1990:DGP

[Pau90b] V. Pauli, Jr. Degeneration of a gas, and paramagnetism. In Einstein collection, 1986–1990 (Russian), pages 158–180. “Nauka”, Moscow, USSR, 1990. Translated from the German by A. N. Temchin.

Pauli:1991:EMR

[Pau91] Wolfgang Pauli. On the earlier and more recent history of the neutrino. In Winter [Win91], pages 1–25. ISBN 0-521-36452-3. LCCN QC793.5.N42 N47 1990.

Pauli:1993:PIG

[Pau93] Wolfgang Pauli. Postscript: “on the isospin group in the theory of the elementary particles”. In Blum et al. [BDR93], pages 337– 351. ISBN 3-540-13848-X. The author reports: “As essential parts of the preprint [[HP93]] with the above title don’t any longer agree with my opinion, I am forced to give up the plan to publish a common paper with Heisenberg on the subject in question.”.

Pauli:1994:AED

[Pau94a] Wolfgang Pauli. Albert Einstein and the development of physics. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 13, pages 117–124. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2

.P38 1994. URL http://www.loc.gov/catdir/enhancements/
fy0812/94015098-d.htm.

Pauli:1994:AS

[Pau94b] Wolfgang Pauli. Arnold Sommerfeld. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 6, pages 69– 72. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York).

LCCN QC6.2 .P38 1994. URL http://www.loc.gov/catdir/
enhancements/fy0812/94015098-d.htm.

Pauli:1994:ECQ

[Pau94c] Wolfgang Pauli. Einstein’s contribution to quantum theory. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 9, pages 85–94. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2

.P38 1994. URL http://www.loc.gov/catdir/enhancements/
fy0812/94015098-d.htm.

Pauli:1994:EPQ

[Pau94d] Wolfgang Pauli. Exclusion principle and quantum mechanics. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 18, pages 165–182. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2

.P38 1994. URL http://www.loc.gov/catdir/enhancements/
fy0812/94015098-d.htm.

Pauli:1994:IUS

[Pau94e] Wolfgang Pauli. Ideas of the unconscious from the standpoint of natural science and epistemology. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 17, pages 149– 164. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York).

LCCN QC6.2 .P38 1994. URL http://www.loc.gov/catdir/
enhancements/fy0812/94015098-d.htm. English translation

of [Pau54b].

Pauli:1994:IAE

[Pau94f] Wolfgang Pauli. Impressions of Albert Einstein. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 12, pages 113–116. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2 .P38 1994. URL http://www.loc.

gov/catdir/enhancements/fy0812/94015098-d.htm.

Pauli:1994:IAI

[Pau94g] Wolfgang Pauli. The influence of archetypal ideas on the scientific theories of Kepler. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 21, pages 219–280. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2

.P38 1994. URL http://www.loc.gov/catdir/enhancements/
fy0812/94015098-d.htm.

Pauli:1994:M

[Pau94h] Wolfgang Pauli. Matter. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 1, pages 27– 34. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York).

LCCN QC6.2 .P38 1994. URL http://www.loc.gov/catdir/
enhancements/fy0812/94015098-d.htm.

Pauli:1994:NBH

[Pau94i] Wolfgang Pauli. Niels Bohr on his 60th birthday. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 4, pages 49–58. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2 .P38 1994. URL http://www.loc.gov/

catdir/enhancements/fy0812/94015098-d.htm.

Pauli:1994:EMR

[Pau94j] Wolfgang Pauli. On the earlier and more recent history of the neutrino. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 20, pages 193–218. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2

.P38 1994. URL http://www.loc.gov/catdir/enhancements/
fy0812/94015098-d.htm.

Pauli:1994:PE

[Pau94k] Wolfgang Pauli. Paul Ehrenfest. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 8, pages 79– 84. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York).

LCCN QC6.2 .P38 1994. URL http://www.loc.gov/catdir/
enhancements/fy0812/94015098-d.htm.

Pauli:1994:PPR

[Pau94l] Wolfgang Pauli. Phenomenon and physical reality. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 15, pages 127–136. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2 .P38 1994. URL http://www.loc.

gov/catdir/enhancements/fy0812/94015098-d.htm.

Pauli:1994:PSI

[Pau94m] Wolfgang Pauli. The philosophical significance of the idea of complementarity. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], pages 35–42. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2 .P38

1994. URL http://www.loc.gov/catdir/enhancements/

fy0812/94015098-d.htm. English translation of [Pau50b].

Pauli:1994:PP

[Pau94n] Wolfgang Pauli. Probability and physics. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 3, pages 43–48. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2 .P38 1994. URL http://www.loc.

gov/catdir/enhancements/fy0812/94015098-d.htm. English

translation of [Pau54c].

Pauli:1994:RPS

[Pau94o] Wolfgang Pauli. Rydberg and the periodic system of the elements. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 7, pages 73–78. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2

.P38 1994. URL http://www.loc.gov/catdir/enhancements/
fy0812/94015098-d.htm.

Pauli:1994:SWT

[Pau94p] Wolfgang Pauli. Science and western thought. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 16, pages 137–148. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2 .P38 1994. URL http://www.loc.

gov/catdir/enhancements/fy0812/94015098-d.htm.

Pauli:1994:SCQ

[Pau94q] Wolfgang Pauli. Sommerfeld’s contributions to quantum theory. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 5, pages 59–68. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2

.P38 1994. URL http://www.loc.gov/catdir/enhancements/
fy0812/94015098-d.htm.

Pauli:1994:STC

[Pau94r] Wolfgang Pauli. Space, time and causality in modern physics. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 10, pages 95–106. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2 .P38 1994. URL http:

//www.loc.gov/catdir/enhancements/fy0812/94015098-d.
htm.

Pauli:1994:TE

[Pau94s] Wolfgang Pauli. Theory and experiment. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 14, pages 125–126. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2 .P38 1994. URL http://www.loc.gov/

catdir/enhancements/fy0812/94015098-d.htm.

Pauli:1994:TRS

[Pau94t] Wolfgang Pauli. The theory of Relativity and science. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 11, pages 107–112. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2 .P38 1994. URL http:

//www.loc.gov/catdir/enhancements/fy0812/94015098-d.
htm.

Pauli:1994:VRS

[Pau94u] Wolfgang Pauli. The violation of reflection symmetries in the laws of atomic physics. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 19, pages 183–192. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2

.P38 1994. URL http://www.loc.gov/catdir/enhancements/

fy0812/94015098-d.htm. English translation of [Pau58c].

Pauli:2001:JBR

[Pau01] Wolfgang Pauli. Das Jahr 1955 Der Berner Relativitätskongreß und der Beitrag zur Bohr-Festschrift. (German) [The year 1955: The Bern Relativity Conference and the contribution to the Bohr Anniversary Publication]. In von Meyenn [vM01], pages 1–461. ISBN 3-540-67591-4, 3-540-78805-0. ISSN 0172-6315. LCCN QC16.P37 A34.

Pauli:1927:TKA

[PB27] W. Pauli and W. Baade. Über den auf die Teilchen in den Kometenschweifen ausgeübten Strahlungsdruck. (German) [On the radiation pressure exerted on particles in comet tails]. Naturwissenschaften, 15(2):49–51, ???? 1927. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://link.springer.com/article/10.1007/BF01504621;

 http://www.springerlink.com/content/t78801822t881576/
fulltext.pdf.

Pauli:1940:SBK

[PB40] W. Pauli and F. J. Belinfante. On the statistical behavior of known and unknown elementary particles. Physica, 7:177–192, 1940. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Pauli:1958:NBR

[PB58] V. Pauli and Niels Bohr, editors. Nil’s Bor i razvitie fiziki: sbornik, posvjascennyj Nil’su Boru v svjazi s ego semidesjatiletiem. (Russian) [Niels Bohr and the development of physics]. Izdatelstvo Inostrannoj Literatury, Moskva, USSR, 1958. 258 pp. LCCN ????

Curceanu:2011:NEL

[PBB^{+}11] C. Curceanu (Petrascu), S. Bartalucci, S. Bertolucci, M. Bragadireanu, M. Cargnelli, S. Di Matteo, J.-P. Egger, C. Guaraldo, M. Iliescu, T. Ishiwatari, M. Laubenstein, J. Marton, E. Milotti, D. Pietreanu, T. Ponta, A. Romero Vidal, D. L. Sirghi, F. Sirghi, L. Sperandio, O. Vazquez Doce, E. Widmann, and J. Zmeskal. New experimental limit on the Pauli Exclusion Principle violation by electrons — the VIP experiment. Foundations of Physics, 41(3):282–287, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-009-9378-6.

Pogosov:2010:TCP

[PCC10] Walter V. Pogosov, Monique Combescot, and Michel Crouzeix. Two-Cooper-pair problem and the Pauli Exclusion Principle. Physical Review B: Condensed Matter and Materials Physics, 81(17):174514:1–174514:9, May 13, 2010. CODEN PRBMDO.

ISSN 1098-0121. URL http://link.aps.org/doi/10.1103/
PhysRevB.81.174514.

Pauli:1942:PMF

[PD42a] W. Pauli and S. M. Dancoff. The pseudoscalar meson field with strong coupling. Physical Review (2), 62(3–4):85–108, August 1, 1942. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v62/
i3-4/p85_1.

Pauli:1942:SCM

[PD42b] W. Pauli and S. M. Dancoff. Strong coupling mesotron theory of nuclear interactions [abstract only]. Physical Review, 61(5– 6):387–??, March 1942. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Pauli:1957:VOE

[PE57] Wolfgang Pauli and Paul Erdős. Vorlesungen über Optik Und Elektronentheorie. (German) [Lectures on Optics and Electron Theory]. Verlag Des Vereins Der Mathematiker Und Physiker An Der ETH, Zurich, Switzerland, second edition, 1957. 90 pp. LCCN ????

Pauli:2000:E

[PE00a] Wolfgang Pauli and Charles P. (Charles Paul) Enz, editors. Electrodynamics, volume 1 of Pauli lectures on physics. Dover, New York, NY, USA, dover edition, 2000. ISBN 0-486-41457-4 (paperback). x + 160 pp. LCCN QC3 .P35

2000 vol. 1. URL http://www.loc.gov/catdir/description/

dover031/00031582.html. Translated by S. Margulies and H. R. Lewis. Foreword by Victor F. Weisskopf.

Pauli:2000:OTE

[PE00b] Wolfgang Pauli and Charles P. (Charles Paul) Enz, editors. Optics and the theory of electrons, volume 2 of Pauli lectures on physics. Dover, New York, NY, USA, 2000. ISBN 0-486-41458-2 (paperback). x + 159 pp. LCCN QC3 .P35

2000 vol. 2. URL http://www.loc.gov/catdir/description/

dover031/00031577.html. Translated by S. Margulies and H. R. Lewis. Foreword by Victor F. Weisskopf.

Pauli:2000:STF

[PE00c] Wolfgang Pauli and Charles P. (Charles Paul) Enz, editors. Selected topics in field quantization, volume 6 of Pauli lectures on physics. Dover, New York, NY, USA, dover edition, 2000. ISBN 0-486-41459-0 (paperback). x + 188 pp. LCCN QC3 .P35

2000 vol. 6. URL http://www.loc.gov/catdir/description/

dover033/00031581.html. Translated by S. Margulies and H. R. Lewis. Foreword by Victor F. Weisskopf.

Pauli:2000:SM

[PE00d] Wolfgang Pauli and Charles P. (Charles Paul) Enz, editors. Statistical mechanics, volume 4 of Pauli lectures on physics.

Dover, New York, NY, USA, dover edition, 2000. ISBN 0-486-41460-4 (paperback). x + 121 pp. LCCN QC3 .P35

2000 vol. 4. URL http://www.loc.gov/catdir/description/

dover032/00031579.html. Translated by S. Margulies and H. R. Lewis. Foreword by Victor F. Weisskopf.

Pauli:2000:TKT

[PE00e] Wolfgang Pauli and Charles P. (Charles Paul) Enz, editors. Thermodynamics and the kinetic theory of gases, volume 3 of Pauli lectures on physics. Dover, New York, NY, USA, dover edition, 2000. ISBN 0-486-41461-2 (paperback). x + 138 pp. LCCN QC3 .P35 2000 vol. 3. URL http://www.loc.gov/

catdir/description/dover032/00031578.html. Translated

by S. Margulies and H. R. Lewis. Foreword by Victor F. Weisskopf.

Pauli:2000:WM

[PE00f] Wolfgang Pauli and Charles P. (Charles Paul) Enz, editors. Wave mechanics, volume 5 of Pauli lectures on physics. Dover, New York, NY, USA, 2000. ISBN 0-486-41462-0 (paperback). xv + 205 pp. LCCN QC3 .P3513

2000 vol. 5. URL http://www.loc.gov/catdir/description/

dover031/00031580.html. Translated by S. Margulies and H. R. Lewis. Foreword by Victor F. Weisskopf.

Pecul:1993:DIS

[Pec93] Krzysztof Pecul. Decomposition and interpretation of the SCF interaction and deformation energies by the modified Pauli blockade method. International Journal of Quantum Chemistry, 47(2):145–153, July 15, 1993. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Peierls:1960:QTS

[Pei60a] R. E. Peierls. Quantum theory of solids. In Theoretical physics in the twentieth century (Pauli memorial volume), pages 140– 160. Interscience, New York, 1960.

Peierls:1960:WEP

[Pei60b] R. E. Peierls. Wolfgang Ernst Pauli: 1900–1958. Biographical Memoirs of Fellows of the Royal Society, 5:174–192, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL http://www.jstor.org/stable/769285.

Peierls:1985:AP

[Pei85a] Sir Rudolf E. (Rudolf Ernst) Peierls. Assistant to Pauli. In Bird of Passage: Recollections of a Physicist [Pei85b], chapter 3, pages 46–81. ISBN 0-691-08390-8, 0-691-02416-2 (paperback), 0-691-60220-4, 1-4008-5461-X (e-book). LCCN QC16.P375

A32 1985. URL http://site.ebrary.com/id/10897402;
https://www.jstor.org/stable/j.ctt7ztn4b.

Peierls:1985:BPR

[Pei85b] Sir Rudolf E. (Rudolf Ernst) Peierls. Bird of Passage: Recollections of a Physicist. Princeton legacy library. Princeton University Press, Princeton, NJ, USA, 1985. ISBN 0-691-08390-8, 0-691-02416-2 (paperback), 0-691-60220-4, 1-4008-5461-X (e-book). xii + 350 + 12 pp. LCCN QC16.P375

A32 1985. URL http://site.ebrary.com/id/10897402;
https://www.jstor.org/stable/j.ctt7ztn4b.

Peierls:1992:WPM

[Pei92] Rudolf Peierls. Where Pauli made his ‘wrong’ remark. Physics Today, 45(12):112, December 1992. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v45/i12/p112/s2.

Pesic:2005:BRC

[Pes05] Peter Pesic. Book review: C. A. Meier (Editor): Atom and Archetype: The Pauli/Jung Letters, 1932–1958. Isis, 96(1):148– 149, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
10.1086/433037.

Petsche:2011:PFG

[Pet11] Hans-Joachim Petsche, editor. From past to future: Graßmann’s work in context: Graßmann Bicentennial Conference, September 2009. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2011. ISBN 3-0346-0404-1 (print), 3-0346-0405-X (e-book). LCCN QA29.G73 F76 2011. URL http://

www.springerlink.com/content/978-3-0346-0405-5.

Pauli:1937:EQG

[PF37] Wolfgang Pauli and M. Fierz. Über das H-Theorem in der Quantenmechanik. (German) [On the H-theorem in quantum mechanics]. Zeitschrift für Physik, 106(9–10):572–587, September 1937. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831

(electronic). URL http://link.springer.com/article/10.
1007/BF01339897;   http://www.springerlink.com/content/
m77326kw8457u871/fulltext.pdf.

Pauli:1938:TEL

[PF38] Wolfgang Pauli and Marcus Eduard Fierz. Zur Theorie der Emission langwelliger Lichtquanten.. (German) [On the theory of emission of long-wave light quanta]. Nicola Zanichelli, Bologna, Italy, 1938. xvi + 22 pp. LCCN ????

Pauli:2000:RGR

[PG00] Wolfgang Pauli and Domenico Giulini. Relativitätstheorie. (German) [Relativity Theory]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67312-1. xv + 300 pp. LCCN ???? URL http://www.gbv.de/

dms/hebis-darmstadt/toc/90428919.pdf;   http://www.
zentralblatt-math.org/zmath/en/search/?an=0959.83001.

Reprint of [Pau21b], newly edited and commented by D. Giulini.

Pauli:1945:SCC

[PH45] W. Pauli and Ning Hu. On the strong coupling case for spin-dependent interactions in scalar- and vector-pair theories. Reviews of Modern Physics, 17(2–3):267–286, April 1, 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.17.267;   http://rmp.aps.org/
abstract/RMP/v17/i2-3/p267_1.

Pauli:1944:ADM

[PJ44] Wolfgang Pauli and J. M. Jauch. On the application of Dirac’s method of field-quantization to the problem of emission of low frequency photons. Physical Review (2), 65(7–8):255–256, April 1944. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v65/

i7-8/p252_1. Meeting in New York City, January 14–15, 1944.

Pauli:1943:TMPa

[PK43a] W. Pauli and S. Kusaka. On the theory of a mixed pseudoscalar and a vector meson field. Physical Review (2), 63(11– 12):400–416, June 1, 1943. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v63/i11-12/p400_1.

Pauli:1943:TMPb

[PK43b] W. Pauli and S. Kusaka. On the theory of a mixed pseudoscalar and a vector meson field. Physical Review (2), 63(63):400–416, June 1, 1943. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v63/i11-12/p400_1.

Pauli:1956:MSR

[PK56] Wolfgang Pauli and G. Källén. O matematicheskoi strukturye renormiruyemoi teorii polya li. (Russian) [On the mathematical structure of renormalizable field theory]. Uspekhi Fizicheskikh Nauk, 60(11):425–444, November 1956. CODEN UF-NAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1956/11/d/.

Plaga:1989:VPP

[Pla89] R. Plaga. Violations of the Pauli principle and the interior of the Sun. Zeitschrift für Physik, 333(4):397–403, December 1989. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01299694;   http://www.springerlink.com/content/h23n5447m7011464/
fulltext.pdf.

Pauli:1926:DDN

[PM26] Wolfgang Pauli and L. Mensing. Über die Dielektrizitätskonstante von Dipolgasen nach der Quantenmechanik. (German) [On the dielectric constant of dipole gases according to quantum mechanics]. Physikalische Zeitschrift, 27(??):509–512, ???? 1926. CODEN PHZTAO. ISSN 0369-982X.

Pauli:1936:RPE

[PR36] Wolfgang Pauli and M. E. Rose. Remarks on the polarization effects in the positron theory. Physical Review (2), 49(6):462– 465, March 15, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v49/i6/p462_1.

Pauli:1955:NBD

[PRW55] W. Pauli, L. Rosenfeld, and V. Weisskopf, editors. Niels Bohr and the development of physics: essays dedicated to Niels Bohr on the occasion of his 70th birthday. McGraw-Hill, New York, NY, USA, 1955. vii + 195 pp. LCCN QC71 .P3 1955.

Pauli:1932:TUEa

[PS32a] Wolfgang Pauli and J. Solomon. La théorie unitaire d’Einstein et Mayer et les équations de Dirac. (French) [The Einstein–Mayer unitary theory and the Dirac equations]. Journal de Physique et le Radium, 3(10):452–463, ???? 1932. CODEN JPRAAJ. ISSN 0368-3842.

Pauli:1932:TUEb

[PS32b] Wolfgang Pauli and J. Solomon. La théorie unitaire d’Einstein et Mayer et les équations de Dirac. (French) [The Einstein–Mayer unitary theory and the Dirac equations]. Journal de Physique et le Radium, 3(12):582–589, ???? 1932. CODEN JPRAAJ. ISSN 0368-3842.

Pelantova:2006:FGG

[PST06] Edita Pelantová, Milena Svobodová, and Sébastien Tremblay. Fine grading of sl(p^{2}, C) generated by tensor product of generalized Pauli matrices and its symmetries. Journal of Mathematical Physics, 47(1):013512, January 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/
i1/p013512_s1.

Pauli:1959:RCF

[PT59] W. Pauli and B. Touschek. Report and comment on F. Gürsey’s “Group structure of elementary particles”. Il Nuovo Cimento (10), 14(supplemento):205–211, 1959. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Pauli:1949:IRR

[PV49] W. Pauli and F. Villars. On the invariant regularization in relativistic quantum theory. Reviews of Modern Physics, 21(3): 434–444, July 1, 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.21.434;   http://rmp.
aps.org/abstract/RMP/v21/i3/p434_1. Special issue in honor

of Albert Einstein on his 70th birthday.

Pauli:1934:QSR

[PW34] W. Pauli and V. Weisskopf. Über die Quantisierung der skalaren relativistischen Wellengleichung. (German) [On the quantization of the scalar relativistic wave equation]. Helvetica Physica Acta, 7(7):709–731, ???? 1934. CODEN HPACAK. ISSN 0018-0238.

Pauling:1935:IQM

[PW35] Linus Pauling and E. Bright Wilson, Jr. Introduction to Quantum Mechanics: With Applications to Chemistry. McGraw-Hill, New York, NY, USA, 1935. xi + 468 pp. LCCN QC174.1 .P34. Reprinted by Dover, 1985.

Patera:1988:PMD

[PZ88] J. Patera and H. Zassenhaus. The Pauli matrices in n dimensions and finest gradings of simple Lie algebras of type A_{n−1}. Journal of Mathematical Physics, 29(3):665–673, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i3/p665_s1.

Qian:1998:AAS

[QS98] Zhixin Qian and Viraht Sahni. Analytical asymptotic structure of the Pauli, Coulomb, and correlation-kinetic components of the Kohn–Sham theory exchange — correlation potential in atoms. International Journal of Quantum Chemistry, 70(4–5):671–680, ???? 1998. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract?ID=75045;
http://www3.interscience.wiley.com/cgi-bin/fulltext?
ID=75045&PLACEBO=IE.pdf.

Raju:1976:GKP

[Raj76] Cvavb Chandra Raju. Generalized Källén–Pauli equation. Journal of Mathematical Physics, 17(1):6–12, January 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i1/p6_s1.

Raju:1977:LMP

[Raj77] Cvavb. Chandra Raju. Lee model with 3V particles. Journal of Mathematical Physics, 18(11):2194–2203, November 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i11/p2194_s1.

Ruark:1925:CFS

[RC25] Arthur Edward Ruark and Roy L. Chenault. CII. fine structures of spectrum lines. Philosophical Magazine, 50(299):937– 956, ???? 1925. CODEN PHMAA4. ISSN 0031-8086.

Reines:1953:DFN

[RC53] Frederick Reines and Clyde L. Cowan. Detection of the free neutrino. Physical Review (2), 92(3):830–831, November 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRev.

92.830. See confirmation [CRH^{+}56].

Rahal:1988:TIV

[RC88] V. Rahal and A. Campa. Thermodynamical implications of a violation of the Pauli Principle. Physical Review A (Atomic, Molecular, and Optical Physics), 38(7):3728–3731, October 1, 1988. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622,

1538-4446, 1538-4519. URL http://link.aps.org/doi/10.
1103/PhysRevA.38.3728.

Reines:1972:CFOa

[Rei72] Frederick Reines, editor. Cosmology, fusion and other matters: George Gamow memorial volume. Colorado Associated University Press, Boulder, CO, USA, 1972. ISBN 0-87081-025-1. xiv + 320 pp. LCCN QC780 .C65.

Richtmyer:1976:PP

[RG76] R. D. Richtmyer and Samuel A. Goudsmit. Prejudice in physics? Physics Today, 29(12):42, December 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See [Gou76, page 42].

Rigden:2002:HEE

[Rig02] John S. Rigden. Hydrogen: the essential element. Harvard University Press, Cambridge, MA, USA, 2002. ISBN 0-674-00738-7. 280 pp. LCCN QD181.H1 R54 2002. URL http://catdir.loc.

gov/catdir/toc/fy022/2001051708.html.

Rai:1977:BTT

[RL77] D. K. Rai and J. Ladik. Breit-type three-electron equation in the Pauli approximation. International Journal of Quantum Chemistry, 12(5):925–935, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roseberg:1991:MW

[Rös91] Ulrich Röseberg, editor. Mathematik und Wirklichkeit. (German) [Mathematics and Reality]. Akad.-Verl., Berlin, Germany, 1991. ISBN 3-05-501349-2. 124 pp. LCCN ????

URL http://www.gbv.de/dms/hbz/toc/ht003858637.PDF;
 http://www.zentralblatt-math.org/zmath/en/search/?
an=0751.00006.

Richter:1979:WPJ

[RP79] Steffen Richter and Wolfgang Pauli. Wolfgang Pauli: die Jahre 1918–1930: Skizzen zu einer wissenschaftlichen Biographie. (German) [Wolfgang Pauli: The years from 1918 to 1930. Sketches of a scientific biography], volume 32 of Veröffentlichungen der Schweizerischen Gesellschaft für Geschichte der Medizin und der Naturwissenschaften. Sauerländer, Aarau, Switzerland, 1979. ISBN 3-7941-1556-2. 112 pp. LCCN ????

Reines:1974:TPE

[RS74] F. Reines and H. W. Sobel. Test of the Pauli Exclusion Principle for atomic electrons. Physical Review Letters, 32(17):954, April 29, 1974. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114

(electronic), 1092-0145. URL http://link.aps.org/doi/10.
1103/PhysRevLett.32.954.

Russell:1924:NFE

[Rus24] Henry Norris Russell. A new form of the Exclusion Principle in optical spectra. Science, 59(1536):512–513, June 6, 1924. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/59/1536/512.2.
full.pdf.

Renneberg:1994:STN

[RW94] Monika Renneberg and Mark Walker, editors. Science, Technology, and National Socialism. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-40374-X (hardcover), 0-521-52860-7 (paperback). xix + 422 pp. LCCN Q127.G3 S36 1994. URL http://www.loc.

gov/catdir/description/cam025/92041633.html;   http:
//www.loc.gov/catdir/samples/cam031/92041633.html;
http://www.loc.gov/catdir/toc/cam028/92041633.html.

Sommerfeld:1933:EMG

[SB33] Arnold Sommerfeld and Hans Bethe. Elektronentheorie der Metalle. (German) [Electron theory of metals]. In Aufbau Der Zusammenhängenden Materie. (German) [Structure of Connected Matter], Handbuch der Physik, pages 333–622. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London,

UK / etc., 1933. URL http://link.springer.com/chapter/
10.1007/978-3-642-91116-3_3. See also English translation

[SB53].

Sommerfeld:1953:ETM

[SB53] Arnold Sommerfeld and Hans A. Bethe. Electron theory of metals. II. Electrons in the periodic potential field. Report AEC-TR-2055, United States Atomic Energy Commission, Technical Information Service, Oak Ridge, TN, USA, October 15, 1953. English translation of [SB33].

Scerri:2016:TSS

[Sce16] Eric R. Scerri. A tale of seven scientists and a new philosophy of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-023299-4 (hardcover), 0-19-023300-1 (e-book). xxxiv + 228 pp. LCCN Q175 .S3044 2016.

Schilpp:1949:AEPa

[Sch49a] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 1 of The Library of Living Philosophers. Cambridge University Press, Cambridge, UK, 1949. ISBN 0-87548-286-4. ISSN 0075-9139. xviii + 781 pp. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Schilpp:1949:AEPb

[Sch49b] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 2 of The Library of Living Philosophers. Cambridge University Press, Cambridge, UK, 1949. ISBN 0-87548-286-4. ISSN 0075-9139. xviii + 781 pp. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Schwinger:1958:SPQ

[Sch58] Julian Schwinger, editor. Selected Papers on Quantum Electrodynamics. Dover books on engineering and engineering physics. Dover, New York, NY, USA, 1958. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35.

Schucking:1996:JPP

[Sch99a] E. L. Schucking. Jordan, Pauli, politics, Brecht … and a variable gravitational constant. In Schucking and Harvey [SH99], pages 1–14. ISBN 0-387-98564-6 (hardcover). LCCN QC173.5

.O6 1999. URL http://www.loc.gov/catdir/enhancements/
fy0816/98020292-d.html;   http://www.loc.gov/catdir/
enhancements/fy0816/98020292-t.html.

Schucking:1999:JPP

[Sch99b] Engelbert L. Schucking. Jordan, Pauli, politics, Brecht, and a variable gravitational constant. Physics Today, 52 (10):26–31, October 1999. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/adsabs.harvard.edu/abs/1999PhT....52j..26S;   http:
//link.aip.org/link/phtoad/v52/i10/p26/s1;   http:/
/www.physicstoday.org/resource/1/phtoad/v52/i10/p26_
s1.

Schirrmacher:2001:WPR

[Sch01] Arne Schirrmacher. Wolfgang Pauli: Relativitätstheorie. Neu herausgegeben und kommentiert von Domenico Giulini [unter Mitwirkung von Karl von Meyenn]. Berlin usw.: Springer Verlag 2000. XV und 300 Seiten, gebunden DM 89. (German) [Wolfgang Pauli, Theory of Relativity. New edited and annotated by Domenico Giulini [with the collaboration of Karl von Meyenn]. Berlin, etc.: Springer Verlag, 2000. xv + 300 pages, bound, DM 89]. Berichte zur Wissenschaftsgeschichte, 24(4):254, 2001. CO-DEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Schucking:2002:BRC

[Sch02] Engelbert L. Schucking. Book review: C. A. Meier, Atom and Archetype: The Pauli–Jung Letters, 1932–1958. Physics Today, 55(9):62–63, September 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v55/i9/p62/s1.

Schwinger:2003:SPQ

[Sch03] Julian Schwinger. Selected Papers on Quantum Electrodynamics. Dover, New York, NY, USA, 2003. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35. Reprint of [Sch58] with ISBN.

Scully:2007:DQP

[Scu07a] Robert J. Scully. The demon and the quantum: from the Pythagorean mystics to Maxwell’s demon and quantum mystery. Wiley-VCH, Weinheim, Germany, 2007. ISBN 3-527-40688-3 (paperback). viii + 271 pp. LCCN QC174.12 .S43 2007. with endnotes by Marlan O. (Marlan Orvil) Scully.

Scully:2007:MMM

[Scu07b] Robert J. Scully. Mathematics, mysticism, and more: From Pythagoras and Plato to Pauli. In The demon and the quantum: from the Pythagorean mystics to Maxwell’s demon and quantum mystery [Scu07a], chapter 1, pages 7–24. ISBN 3-527-40688-3 (paperback). LCCN QC174.12 .S43 2007. with endnotes by Marlan O. (Marlan Orvil) Scully.

Scully:2007:QMB

[Scu07c] Robert J. Scully. On quantum mechanics and the big questions: From quantum controversy to Pauli’s spiritual complementarity. In The demon and the quantum: from the Pythagorean mystics to Maxwell’s demon and quantum mystery [Scu07a], chapter 10, pages 137–164. ISBN 3-527-40688-3 (paperback). LCCN QC174.12 .S43 2007. with endnotes by Marlan O. (Marlan Orvil) Scully.

Serwer:1977:UZP

[Ser77] Daniel Serwer. Unmechanischer Zwang: Pauli, Heisenberg, and the rejection of the mechanical atom, 1923–1925. Historical Studies in the Physical Sciences, 8(??):189–256, ???? 1977. CO-DEN HSPSAS. ISSN 0073-2672. URL http://www.jstor.org/

stable/27757371.

Seth:2009:ZAZ

[Set09] Suman Seth. Zweideutigkeit about “Zweideutigkeit”: Sommerfeld, Pauli, and the methodological origins of quantum mechanics. Studies in History and Philosophy of Modern Physics, 40(4): 303–315, December 2009. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1355219809000276.

Sidhu:1978:DPS

[SG78] S. S. Sidhu and R. H. Good, Jr. Definition of polarization of a spin 1/2 particle in an external electromagnetic field. Journal of Mathematical Physics, 19(5):935–936, May 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i5/p935_s1.

Schucking:1999:EPE

[SH99] E. L. (Engelbert L.) Schucking and Alex Harvey, editors. On Einstein’s path: essays in honor of Engelbert Schucking. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98564-6 (hardcover). LCCN QC173.5 .O6 1999. URL http://www.loc.gov/

catdir/enhancements/fy0816/98020292-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0816/98020292-t.
html.

Spitz:1981:RPP

[SHS81] G. Spitz, K. Hahn, and E. W. Schmid. The role of partly Pauliforbidden states in nucleus-nucleus elastic scattering. Zeitschrift für Physik, 303(3):209–213, September 1981. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01424761;
 http://www.springerlink.com/content/n1008p36772w64q3/
fulltext.pdf.

Spitz:1985:CIP

[SKS85] G. Spitz, H. Klar, and E. W. Schmid. The combined influence of the Pauli principle and saturation on the effective α and α–α interactions. Zeitschrift für Physik, 322(1):49–57, March 1985. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01412017;   http://www.springerlink.com/content/q74733802p748r62/
fulltext.pdf.

Spangenburg:2008:NBA

[SM08] Ray Spangenburg and Diane Kit Moser, editors. Niels Bohr: Atomic Theorist. Makers of modern science. Chelsea House, New York, NY, USA, revised edition, 2008. ISBN 0-8160-6178-

5. xvi + 141 pp. LCCN QC16.B63 S63 2008.

Smoluchowski:1987:PP

[Smo87] R. Smoluchowski. A portrait of Pauli. Physics Today, 40 (2):9, February 1987. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v40/i2/p9/s1.

Smutny:1990:EMW

[Smu90] František Smutný. Ernst Mach and Wolfgang Pauli’s ancestors in Prague. European Journal of Physics, 11(5):257, 1990. CO-DEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic).

URL http://stacks.iop.org/0143-0807/11/i=5/a=001.

Schmid:1982:ITB

[SOCg82] E. W. Schmid, M. Orlowski, and Bao Cheng-guang. Importance of the three-body Pauli potential in three-cluster systems. Zeitschrift für Physik, 308(3):237–241, September 1982. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01418285;   http://www.springerlink.com/content/p1730136r7521821/
fulltext.pdf.

Stachel:1977:CPS

[SP77] John Stachel and Jerzy Plebański. Classical particles with spin. I. The WKBJ approximation. Journal of Mathematical Physics, 18(12):2368–2374, December 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/
i12/p2368_s1.

Schwink:1975:SPE

[SS75] Ch. Schwink and O. Schärpf. Solution of the Pauli-equation for neutrons in varying magnetic fields and its application to reflection and transmission at helical magnetic structures. Zeitschrift für Physik. B, Condensed matter, 21(3):305–311, ???? 1975. CODEN ZPCMDN. ISSN 0722-3277 (print), 1431-584X (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01313312;   http://www.springerlink.com/content/m1n3h0236j4q6u9x/
fulltext.pdf.

Samuelson:1998:NLI

[SS98] Bengt Samuelson and Michael Sohlman. Nobel lectures: including presentation speeches and laureates’ biographies, physics: 1942–1962. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1998. ISBN 981-02-3403-1. ISSN 0169-006X. xi + 619 pp. LCCN ????

Schmid:1982:CPB

[SSF82] E. W. Schmid, S. Saito, and H. Fiedeldey. The concept of a Pauli barrier in nucleus-nucleus scattering. Zeitschrift für Physik, 306 (1):37–42, March 1982. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.

com/article/10.1007/BF01413405;   http://www.springerlink.
com/content/k11562wh56763k16/fulltext.pdf.

Scott:1982:GPC

[ST82] N. S. Scott and K. T. Taylor. A general program to calculate atomic continuum processes incorporating model potentials and the Breit–Pauli Hamiltonian within the R-matrix method. Computer Physics Communications, 25(4):347–387, April 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582900765.

Scott:1984:GPC

[ST84] N. S. Scott and K. T. Taylor. A general program to calculate atomic continuum processes incorporating model potentials and the Breit–Pauli Hamiltonian within the R-matrix method. Computer Physics Communications, 35(1–3):C–794–C–797, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829185.

Stachel:1980:LPL

[Sta80] John Stachel. Light on the Pauli legend. Nature, 285(5765): 515–516, June 12, 1980. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v285/n5765/pdf/285515a0.pdf.

Staley:2007:BRA

[Sta07] Kent W. Staley. Book review: Allan Franklin, No Easy Answers: Science and the Pursuit of Knowledge. Pittsburgh: University of Pittsburgh Press, 2005. Pp. xii + 258. $29.95 (hardback). Michela Massimi, Pauli’s Exclusion Principle: The Origin and Validation of a Scientific Principle. Cambridge: Cambridge University Press, 2005. Pp. xiv+211. $80.00 (hardback). British Journal for the History of Science, 40(3):455–457, September 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4500761.

Stachel:2009:GC

[Sta09] John J. Stachel. Going Critical: Volume One: The Challenge of Practice, volume 201 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-1308-6 (hardcover). 407 (est.) pp. LCCN ????

Sticker:1977:BRW

[Sti77] Bernhard Sticker. Book review: Der Weg der Physik. 2500 Jahre physikalischen Denkens. Texte von Anaximander bis Pauli by Shmuel Sambursky. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 61(3):308–310, ???? 1977. CODEN SUARAH. ISSN 0039-4564. URL http://www.jstor.org/

stable/20776487.

Stoner:1924:DEA

[Sto24] Emund C. Stoner. Distribution of electrons among atomic levels. Philosophical Magazine, 48(286):719–736, ???? 1924. CODEN PHMAA4. ISSN 0031-8086.

Straumann:2000:WPM

[Str00] N. Straumann. On Wolfgang Pauli’s most important contributions to physics. arxiv.org, ??(??):1–21, October 2, 2000.

URL http://arxiv.org/abs/physics/0010003. Opening talk

at the symposium: Wolfgang Pauli and Modern Physics, in honor of the 100th anniversary of Wolfgang Pauli s birthday, ETH (Zurich), May 4–6 2000.

Straumann:2004:BRB

[Str04a] N. Straumann. Book review: No time to be brief: a scientific biography of Wolfgang Pauli: Charles P. Enz; Oxford University Press, New York, 2002, pp. 581, price US$60.00, £35.00, ISBN 0-19-856479-1. Studies in History and Philosophy of Modern Physics, 35(3):544–547, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000322. See [Enz02].

Straumann:2004:REP

[Str04b] N. Straumann. The role of the Exclusion Principle for atoms to stars: a historical account. arxiv.org, ??(??), March 28, 2004.

URL http://www.arxiv.org/abs/quant-ph/0403199.

Straumann:2007:REP

[Str07] N. Straumann. The role of the Exclusion Principle for atoms to stars: a historical account. International Review of Physics, 1 (??):184–196, ???? 2007. ISSN 1971-680X.

Straumann:2008:WPM

[Str08] Norbert Straumann. Wolfgang Pauli and modern physics. arxiv.org, ??(??):1–17, October 14, 2008. URL http://arxiv.

org/pdf/0810.2213.

Strickland:2011:WSC

[Str11] Jeffrey Strickland. Weird scientists — the creators of quantum physics. Lulu.com, ????, 2011. ISBN 1-257-97624-9. LCCN ????

Sullivan:1962:NFT

[Sul62] Walter Sullivan. Neutrino is found to be twins. New York Times, ??(??):??, July 1, 1962. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. News of the discovery of the antineutrino.

Suzuki:1973:CIP

[Suz73] Masuo Suzuki. Correlation inequalities and phase transition in the generalized X −Y model. Journal of Mathematical Physics, 14(7):837–838, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i7/p837_s1.

Schreckenbach:1997:CNS

[SZ97] Georg Schreckenbach and Tom Ziegler. Calculation of NMR shielding tensors based on density functional theory and a scalar relativistic Pauli-type Hamiltonian. the application to transition metal complexes. International Journal of Quantum Chemistry, 61(6):899–918, February 20, 1997. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract?ID=42474;
http://www3.interscience.wiley.com/cgi-bin/fulltext?
ID=42474&PLACEBO=IE.pdf.

Szalek:1997:PVM

[Sza97] Marek A. Szalek. Pauli versus the Maxwell equations and the Biot–Savart law. Physics Essays, 10(1):95–102, ???? 1997. CO-DEN PHESEM. ISSN 0836-1398.

Tchrakian:1972:FFS

[Tch72] D. H. Tchrakian. Form factors for any spin and charge coupling. Journal of Mathematical Physics, 13(10):1477–1484, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v13/i10/p1477_s1.

Telegdi:2003:BRC

[Tel03] Valentine L. Telegdi. Book review: Charles P. Enz, No Time to Be Brief: A Scientific Biography of Wolfgang Pauli. Physics Today, 56(11):67–68, November 2003. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v56/i11/p67/s1.

TerHaar:1965:WP

[Ter65] D. Ter Haar. Wolfgang Pauli 1900–1958. Contemporary Physics, 6(6):465–471, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thomas:1926:MSE

[Tho26a] L. H. (Llewellyn Hilleth) Thomas. The motion of the spinning electron. Nature, 117(2945):514, April 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Thorsen:1926:BSW

[Tho26b] V. Thorsen. Über die Seriendarstellung des Wismutspektrums. (German) [On the series representation of the bismuth spectrum]. Zeitschrift für Physik, 40(8):642–647, January 1926. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01390463;   http://www.springerlink.com/content/k0613580220077v0/

.

Tomonaga:1997:SS

[Tom97] Shinichirō Tomonaga. The Story of Spin. University of Chicago Press, Chicago, IL, USA, 1997. ISBN 0-226-80793-2 (hardcover), 0-226-80794-0 (paperback). xii + 258 pp. LCCN QC793.3.S6 T6513 1997. URL http://catdir.loc.gov/

catdir/description/uchi052/97012189.html;   http://
catdir.loc.gov/catdir/toc/uchi051/97012189.html.

Uhlenbeck:1925:EHU

[UG25] George E. Uhlenbeck and Samuel Goudsmit. Ersetzung der Hypothese vom unmechanischen Zwang durch eine Forderung bezüglich des inneren Verhaltens jedes einzelnen Elektrons. (German) [Replacement of the hypothesis of unmechanical force with a requirement for the inner behavior of each electron]. Naturwissenschaften, 13(47):953–954, November 1925. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01558878;   http://www.springerlink.com/content/u20q85248gp8166q/

. See [?] for a computation in 1927 of the radius of a rotating charged particle with the spin of the electron.

Uhlenbeck:1926:SES

[UG26] G. E. Uhlenbeck and S. Goudsmit. Spinning electrons and the structure of spectra. Nature, 117(2938):264–265, February 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). The paper is followed by a half-column supportive letter from Niels Bohr. See [Tho26a] for a relativistic correction to the Zeeman splitting of spectral lines.

Uhlenbeck:1976:FYS

[Uhl76] George E. Uhlenbeck. Fifty years of spin: Personal reminiscences. Physics Today, 29(6):43–48, June 1976. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/
phtoad/v29/i6/p40_s1.

Unser:2000:WPM

[Uns00] Margit Unser, editor. Wolfgang Pauli and modern physics = Wolfgang Pauli und die moderne Physik. ETH-Bibliothek, Zürich, Switzerland, 2000. ISBN 3-9521386-2-2. 124 pp. LCCN QC16.P37 W635 2000.

vonBaeyer:2010:BRA

[vB10] Hans Christian von Baeyer. Book review: Arthur I. Miller, Deciphering the Cosmic Number: The Strange Friendship of Wolfgang Pauli and Carl Jung. New York: W.W. Norton & Company, 2009, xxv + 336 pages. $27.95 (cloth). Physics in Perspective (PIP), 12(4):497–499, December 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vanderWaerden:1967:SQM

[vdW67] B. L. (Bartel Leendert) van der Waerden, editor. Sources of Quantum Mechanics. Classics of science. North-Holland Publishing Co., Amsterdam, The Netherlands, 1967. xi + 430 pp. LCCN QC174.12 S655.

vanderWaerden:1968:SQM

[vdW68] B. L. (Bartel Leendert) van der Waerden, editor. Sources of quantum mechanics, volume 5 of Classics of science. Dover, New York, NY, USA, 1968. ISBN 0-486-61881-1. xi + 430 pp. LCCN QC174.1 .W3 1968.

vanderWaerden:1973:U

[vdW73] B. L. van der Waerden. [unknown]. In Mehra [Meh73], pages 315–?? ISBN 90-277-0345-0, 90-277-2536-5. LCCN

QC173.96 .S95 1972. URL http://www.springer.com/us/
book/9789027703453.

vanderWaerden:1952:BWS

[vdWPR52] B. L. van der Waerden, W. Pauli, and S. Rosin. Der Begriff der Wahrscheinlichkeit und seine Rolle in den Naturwissenschaften. (German) [The concept of probability and its role in the natural sciences]. Actes de la Société Helvétique des Sciences Naturelles, 132:74–82; Diskussion 82–86 (1953), 1952. CODEN VSNGAY. ISSN 0080-7362.

vanErkelens:1995:SVW

[vE95] Herbert van Erkelens. Het spel van de wijsheid: Pauli, Jung en de menswording van God. (Dutch) [The game of wisdom: Pauli, Jung and the incarnation of God]. Kok Agora, Kampen, The Netherlands, 1995. ISBN 90-391-0668-1. 182 pp. LCCN BL245 .E75 1995.

vanErkelens:1998:WPC

[vE98] Herbert van Erkelens. Wolfgang Pauli and the Chinese animafigure. In Eranos Yearbook, page ?? The Eranos Foundation, Anscona, Switzerland, 1998. Cited in [MEF01, page 234].

vanErkelens:19xx:WPA

[vExx] Herbert van Erkelens. Wolfgang Paulis Alchemistische Queste. (German) [Wolfgang Pauli’s alchemical quest]. Cited without year in [MEF01, page 233]., 19xx.

vanErkelens:2002:WPG

[vEA02] Herbert van Erkelens and Thomas Arzt. Wolfgang Pauli und der Geist der Materie. (German) [Wolfgang Pauli and the spirit of matter], volume 7 of Studien aus der existential-psychologischen Bildungs- und Begegnungsstätte Todtmoos-Rütte. Königshausen and Neumann, Würzburg, Germany, 2002. ISBN 3-8260-2222-X. 278 pp. LCCN QC16.P37 E75 2002.

vonMeyenn:1968:PBW

[vM68] Karl von Meyenn. Paulis Briefe als Wegbereiter wissenschaftlicher Ideen. (German) [Pauli’s letters as scientific pioneering ideas]. In Chandrasekharan [Cha68], pages 20–42.

vonMeyenn:1985:PSC

[vM85] K. von Meyenn. Pauli, Schrödinger and the conflict about the interpretation of quantum mechanics. In Lahti and Mittelstaedt [LM85], pages 289–302. ISBN 9971-5-0004-3. LCCN QC173.96 .S96 1985.

vonMeyenn:1987:PBE

[vM87] K. von Meyenn. Pauli’s belief in exact symmetries. In García Doncel et al. [GHMP87], pages 329–360. ISBN 84-7488-148-X. LCCN QC174.17.S9 I57 1983.

vonMeyenn:1990:BRK

[vM90] Karl von Meyenn. Book review: K. V. Laurikainen, Beyond the Atom: The Philosophical Thought of Wolfgang Pauli. Physics Today, 43(10):99–101, October 1990. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v43/i10/p99/s1.

vonMeyenn:1993:WPW

[vM93] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band III: 1940–1949. Scientific Correspondence with Bohr, Einstein, Heisenberg, a.o. Volume III: 1940–1949, volume 11 of Sources in the history of mathematics and physical sciences. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-54911-0. ISSN 0172-6315. lxiv + 1070

pp. LCCN QC16.P37 A34. URL http://www.springer.com/
physics/book/978-3-540-54911-6.

vonMeyenn:1996:WPW

[vM96] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil I: 1950–1952. Scientific Correspondence with Bohr, Einstein, Heisenberg, a.o. Volume IV, Part I: 1950–1952, volume 14 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-59442-6. ISSN 0172-6315. xxxvii + 968 pp. LCCN QC16.P37 A34. URL http://

www.springerlink.com/content/978-3-540-78803-4.

vonMeyenn:1999:WPW

[vM99] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil II: 1953–1954. Scientific Correspondence with Bohr, Einstein, Heisenberg, a.o. Volume IV, Part II: 1953–1954, volume 15 of Sources in the History of Mathematics and Physical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-64312-5. ISSN 0172-6315. xxxv + 1100 pp. URL http://www.springer.com/

physics/book/978-3-540-64312-8.

vanMeijgaard:2000:WPC

[vM00] Harry van Meijgaard. Wolfgang Pauli centennial 1900–2000: an extension of the dissertation 1998. Emmy Memorial Center, Oldenzaal, The Netherlands, 2000. ISBN 90-805493-1-2. 184 pp. LCCN QC16.P37 M45 2000.

vonMeyenn:2001:WPW

[vM01] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil III: 1955–1956 Scientific Correspondence With Bohr, Einstein, Heisenberg, a.o. Volume IV, Part III: 1955–1956, volume 17 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-67591-4, 3-540-78805-0. ISSN 0172-6315. lxv + 994 pp. LCCN QC16.P37 A34.

vonMeyenn:2005:HZP

[vM05a] Karl von Meyenn. Heisenbergs Zusammenarbeit mit Pauli während der Leipziger Jahre. (German) [Heisenberg’s collaboration with Pauli during the Leipzig years]. In Leipziger Festschrift zu Werner Heisenbergs 100. Geburtstag. (German) [Leipzig Celebrations on Werner Heisenberg’s 100th Birthday], pages 58–81. Berichte der Sächsischen Akademie der Wissenschaften, Leipzig, Germany, 2005. ISBN ???? LCCN ????

vonMeyenn:2005:WPW

[vM05b] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil IV: 1957–1958 Scientific Correspondence With Bohr, Einstein, Heisenberg, a.o. Volume IV, Part IV: A:1957, B:1958, volume 18 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-40296-9. ISSN 0172-6315. xl + 1585 pp. LCCN QC16.P37 A34. URL http:/

/www.springerlink.com/content/978-3-540-26832-1.

vonMeyenn:1985:WPW

[vMHW85] Karl von Meyenn, Armin Hermann, and Victor F. (Victor Frederick) Weisskopf, editors. Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Band II, 1930– 1939. Scientific Correspondence with Bohr, Einstein, Heisenberg a.o. Volume II: 1930–1939, volume 6 of Sources in the History of Mathematics and Physical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-13609-6 (New York), 3-540-13609-6 (Berlin). xxxix + 783 pp. LCCN QC16.P37 A34. URL http://www.

springerlink.com/content/978-3-540-13609-5.

vonMeyenn:2001:WP

[vMS01] Karl von Meyenn and Engelbert Schucking. Wolfgang Pauli. Physics Today, 54(2):43–48, February 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/phtoad/v54/i2/p43/s1.

Varlaki:2008:NAB

[VNB08] Péter Várlaki, László Nadai, and József Bokor. Number archetypes and background control theory concerning the fine structure constant. Acta Polytechnica Hungarica, 5(2): 71–104, ???? 2008. CODEN ???? ISSN 1785-8860.

URL http://web.archive.org/web/20110721103929/http:/
/bmf.hu/journal/Varlaki_Nadai_Bokor_14.pdf.

vonOertzen:1969:TGP

[vO69] Wolfram von Oertzen. On the test of the generalized Pauli Principle for complex nuclei in nuclear reactions. Zeitschrift für Physik, 228(2):182–187, April 1969. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01397539;
 http://www.springerlink.com/content/j12043g28r6rwu88/
fulltext.pdf.

vonWeizsacker:1959:EWP

[vW59] Carl Friedrich von Weizsäcker. Erinnerungen an Wolfgang Pauli. (German) [Memories of Wolfgang Pauli]. Zeitschrift für Naturforschung, 14a(??):439–??, ???? 1959. CODEN ZNTFA2. ISSN 0372-9516.

Weber:1973:RWS

[Web73] Robert L. Weber, editor. A Random Walk In Science. The Institute of Physics, Bristol and London, 1973. ISBN 0-85498-027-X. xvii + 206 pp. LCCN Q167.W42. Compiled by R. L. Weber, edited by E. Mendoza, with a foreword by William Cooper.

Weisskopf:1970:TSS

[Wei70] Victor F. Weisskopf. Three steps in the structure of matter. Physics Today, 23(8):17–24, August 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

www.physicstoday.org/resource/1/phtoad/v23/i8/p17_s1.

Weisskopf:1985:PMP

[Wei85] Victor F. Weisskopf. Personal memories of Pauli. Physics Today, 38(12):36–41, December 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v38/i12/p36/s1.

Weiner:1988:OHS

[Wei88] Charles Weiner. Oral history of science: A mushrooming cloud? The Journal of American History, 75(2):548–559, September 1988. ISSN 0021-8723 (print), 1945-2314 (electronic). URL

http://www.jstor.org/stable/1887871.

Weisskopf:1991:AP

[Wei91a] Victor Frederick Weisskopf. Adventures with Pauli. In The Joy of Insight: Passions of a Physicist [Wei91b], page ?? ISBN 0-465-03678-3, 0-465-03677-5. LCCN QC16.W516 A3 1991.

Weisskopf:1991:JIP

[Wei91b] Victor Frederick Weisskopf. The Joy of Insight: Passions of a Physicist. Alfred P. Sloan Foundation series. Basic Books, New York, NY, USA, 1991. ISBN 0-465-03678-3, 0-465-03677-5. xiv + 336 + 8 pp. LCCN QC16.W516 A3 1991.

Wessels:1985:BRI

[Wes85] Linda Wessels. Book review: Issues in quantum physics: The creation of quantum mechanics and the Bohr–Pauli dialogue, John Hendry, Reidel, Boston, 1984. Science (New Series), 229 (4717):963–964, September 6, 1985. CODEN SCNCAD. ISSN 0036-8075 (print), 1095-9203 (electronic).

Winter:1991:NP

[Win91] Klaus Winter, editor. Neutrino physics. Cambridge monographs on particle physics, nuclear physics, and cosmology. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-36452-3. xiv + 670 pp. LCCN QC793.5.N42 N47 1990.

Woit:2006:EWF

[Woi06] Peter Woit. Not even wrong: the failure of string theory and the search for unity in physical law. Basic Books, New York, NY, USA, 2006. ISBN 0-465-09275-6. xix + 291 pp. LCCN

QC794.6.S85 W65 2006. URL http://www.loc.gov/catdir/
toc/ecip0613/2006013933.html. See critical review [Ber06].

Xue:1991:CGS

[Xue91] She-Sheng Xue. Chiral-gauge symmetry and Pauli–Villars regularization scheme. Zeitschrift für Physik. C, Particles and fields, 50(1):145–147, March 1991. CODEN ZPCFD2. ISSN 0170-9739 (print), 1431-5858 (electronic). URL

http://link.springer.com/article/10.1007/BF01558569;
 http://www.springerlink.com/content/g07567h8q7031725/
fulltext.pdf.

Yorke:1973:BRW

[YK73] Ellen Yorke and Ivan Kramer. Book review: Wolfgang Pauli and C. P. Enz, editors, Pauli Lectures on Physics. Physics Today, 26(12):53–55, December 1973. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v26/i12/p53/s1. See remarks in [LM74].

Yurke:1993:UPE

[YS93] Bernard Yurke and David Stoler. Using the Pauli Exclusion Principle to exhibit local-realism violations in overlapping interferometers. Physical Review A (Atomic, Molecular, and Optical Physics), 47(3):1704–1707, March 1, 1993. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL

http://link.aps.org/doi/10.1103/PhysRevA.47.1704.

Zeeman:1930:HWG

[ZBG30] P. Zeeman, E. Back, and S. Goudsmit. Zur Hyperfeinstruktur des Wismuts. (German) [On the hyperfine structure of bismuth]. Zeitschrift für Physik, 66(1–2):1–12, January 1930. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01397521;
 http://www.springerlink.com/content/n27vx7850217172r/

.

Zatsarinny:2000:GPC

[ZF00] Oleg Zatsarinny and Charlotte Froese Fischer. A general program for computing angular integrals of the Breit–Pauli Hamiltonian with non-orthogonal orbitals. Computer Physics Communications, 124(2–3):247–289, February 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004415.

Zhalij:2002:SPE

[Zha02] Alexander Zhalij. On separable Pauli equations. Journal of Mathematical Physics, 43(3):1365–1389, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ziyang:1982:ADC

[ZSJY82] Sun Ziyang, Wang Shunuan, Zhang Jingshang, and Zhuo Yizhong. Angular distribution calculation based on the exciton model taking account of the influence of the Fermi motion and the Pauli principle. Zeitschrift für Physik, 305(1):61–68, March 1982. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831

(electronic). URL http://link.springer.com/article/10.
1007/BF01415080;   http://www.springerlink.com/content/
k6133j6075w1wh7n/fulltext.pdf.

Zhang:1988:PEE

[ZY88] J. S. Zhang and X. J. Yang. The Pauli exclusion effect in multiparticle and hole state densities. Zeitschrift für Physik, 329(1): 69–76, March 1988. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic). URL http://link.springer.com/
article/10.1007/BF01294817;   http://www.springerlink.
com/content/n215385305621445/fulltext.pdf.